### v2

# A Universal Scheme for Dynamic Partitioned Shortest Path Index

Abstract.
Shortest path (SP) computation is the fundamental operation in various networks such as urban networks, logistic networks, communication networks, social networks, etc. With the development of technology and societal expansions, those networks tend to be massive. This, in turn, causes deteriorated performance of SP computation, and graph partitioning is commonly leveraged to scale up the SP algorithms. However, the partitioned shortest path (PSP) index has never been systematically investigated and theoretically analyzed, and there is a lack of experimental comparison among different PSP indexes. Moreover, few studies have explored PSP index maintenance in dynamic networks. Therefore, in this paper, we systematically analyze the dynamic PSP index by proposing a universal scheme for it. Specifically, we first propose two novel partitioned shortest path strategies (No-boundary and Post-boundary strategies) to improve the performance of PSP indexes and design the corresponding index maintenance approaches to deal with dynamic scenarios. Then we categorize the partition methods from the perspective of partition structure to facilitate the selection of partition methods in the PSP index. Furthermore, we propose a universal scheme for designing the PSP index by coupling its three dimensions (i.e. PSP strategy, partition structure, and SP algorithm). Based on this scheme, we propose five new PSP indexes with prominent performance in either query or update efficiency. Lastly, extensive experiments are implemented to demonstrate the effectiveness of the proposed PSP scheme, with valuable guidance provided on the PSP index design.
PVLDB Reference Format:
PVLDB, (): , .
doi: This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. , No.  ISSN 2150-8097.
doi:
footnotetext: *Both authors contributed equally to this paper. Lei Li is the corresponding author.
1. Introduction
Shortest Path (SP) query in dynamic networks is an essential building block for various applications affecting daily lives. Given an origin-destination pair, it returns the minimum traveling time in a road network (Bast et al., 2016), the fastest connection in the web graph (Boccaletti et al., 2006), or the most intimate relationships in the social network (Kolaczyk et al., 2009; Opsahl et al., 2010; Yao et al., 2013). With the development of urban traffic systems and the evolvement of online interactions platforms, real-life networks tend to be massive, which brings great challenges to the scalability of the state-of-the-art (Li et al., 2020a, 2019b; Zhang et al., 2021a, c, b, 2023a) with either heavy memory or expensive search overheads. For example, 2-Hop Labeling (Cohen et al., 2003) has the fastest query efficiency but requires $O(nm^{1/2})$ space (for a graph with $n$ vertices and $m$ edges) to store the index. Complementary, the search-based methods like Dijkstra’s are inefficient on query processing on large networks. Moreover, because the networks are dynamic in nature with evolving structures or edge weights, extra efforts are required to handle the updates for SP indexes.
Graph partitioning is often used to improve the scalability of SP algorithms, enabling more complicated problems like time-dependent (Li et al., 2019c, 2020b) and constraint path computation (Wang et al., 2016; Liu et al., 2021, 2022a) to be efficiently handled in large networks. It decomposes a large network into several smaller ones such that the construction time can be improved by parallelization, and the index sizes and update workloads can be reduced. As shown in Figure 1, we can compare them briefly in terms of index construction time, storage space, query time, and update time. On one extreme are the direct search algorithms (Dijkstra, 1959; Hart et al., 1968; Li et al., 2020c; Zhang et al., 2020) that require no index and can work in dynamic environments directly but are slow for query answering. Then the partitioned search adds various information to guide and reduce search space (Chondrogiannis and Gamper, 2016; Möhring et al., 2007; Bauer and Delling, 2010) to improve query processing. On the other extreme, Contraction Hierarchy (CH) (Geisberger et al., 2008) and 2-hop Labeling (HL) (Cohen et al., 2003; Akiba et al., 2013; Li et al., 2019b; Ouyang et al., 2018) have larger index sizes, longer construction and maintenance time but faster query performance. Their partitioned versions (Delling et al., 2017; Li et al., 2020a; Li et al., 2019c; Wang et al., 2021; Farhan et al., 2022; Liu et al., 2021, 2022b, 2022a) are faster to construct with smaller sizes but longer query times. As these works have better performance than their unpartitioned versions, they create the following misconceptions regarding the benefits of partitions on pathfinding-related problems: 1) Graph partitioning and shortest path algorithms are irrelevant since they are two different research branches; 2) Different PSP indexes are irrelevant and very different from each other; 3) Applying graph partitioning is always better.
Figure 1. Performance Comparison of SP Algorithms
Motivations. These misconceptions are not always true. Although the partitioned shortest path (PSP) indexes have been widely applied in the past decade, to the best of our knowledge, they have not been studied systematically. Specifically, the existing solutions typically pick one partition method without discriminating their characteristics and figure out a way to make it work, and then claim benefits/superiority. Consequently, there lacks a generalized scheme to organize and compare the PSP indexes insightfully and fairly. Towards that, we postulate that both theoretical and experimental systematic study is needed for PSP indexes, so that a suitable index could be constructed for different scenarios.
Challenges. However, it is non-trivial to design such a scheme. Firstly, the current PSP indexes are delicately designed structures with tightly coupled components and all claim certain superiority. However, since no effort has been put into the abstract PSP index itself, it is unclear what determines their performance, when to use particular indexes, and how to design a suitable new PSP index for a new scenario. Therefore, we study the PSP index systematically in this work and identify three dimensions throughout the rich literature. Secondly, it is unclear how to construct, query, and update a new PSP index as all existing works only claim their unique structures can answer queries correctly and seldom involve index maintenance. Therefore, we dive into the partitioned shortest path (PSP) strategy, which determines how to coordinate multiple partitions for correct query answering, and summarizes the traditional PSP strategy. Additionally, we propose two novel PSP strategies together with their index update methods to improve the index performance. Besides, a pruning-based overlay graph optimization is also proposed to improve the index and update efficiency further. Thirdly, the graph partition methods are normally classified from different perspectives (like vertex-cut / edge-cut, in-memory / streaming, etc.), but these criteria are not path-oriented and thus hardly help to tell which partition method is suitable for certain path-oriented applications. Therefore, we propose a path-oriented partitioning classification from the perspective of partition structure for easier partition method selection. Lastly, after elaborating on each dimension of PSP index, we propose our universal PSP index scheme by entangling these dimensions. Based on this scheme, we propose five new PSP indexes to achieve better efficiency in either query or update operations by selecting the specific strategy or method on each dimension as per system requirements.
Contributions. Our contributions are listed as follows:
  • $\bullet$
    We propose a universal PSP index scheme by decoupling the PSP index into three dimensions: PSP strategy, partition structure, and SP algorithm, such that the PSP index can be analyzed and designed systematically. (Section 2)
  • $\bullet$
    We propose two novel PSP strategies called No-Boundary and Post-Boundary strategies to improve the index construction and maintenance efficiency over the traditional PSP strategy. We also provide non-trivial correctness analysis for their index construction, query processing, and index maintenance. A novel Pruning-based overlay graph optimization is also designed to prune the unnecessary computation. (Section 3)
  • $\bullet$
    We propose a new path-oriented partitioning classification from the perspective of partition structure for easier partition method selection in designing the PSP index. (Section 4)
  • $\bullet$
    Based on the proposed PSP index scheme, we put forward a PSP index generator to create new PSP structures. Among them, we identify and design five new PSP indexes suitable for specific network structures and scenarios. (Section 5)
  • $\bullet$
    We implement comprehensive experimental studies to explore the performance of PSP indexes under our proposed universal scheme and provide guidelines for designing PSP indexes with different application scenarios. (Section 6)
2. Preliminary and PSP Dimensions
In this paper, we focus on the dynamic weighted network $G(V,E,W)$ with $V$ denoting the vertex set, $E$ denoting the edge set and $W:E\rightarrow\mathbf{R}^{+}$ assigning a non-negative weight to each edge – i.e. $w(u,v)\in W,\forall(u,v)\in E$ – which can increase or decrease in an ad-hoc manner. We denote the number of vertices and edges as $n=|V|$ and $m=|E|$. For each $v\in V$, we represent its neighbors as $N(v)=\{u,|(u,v)\in E\}$, u’s commonly called adjacent to $v$. We associate each vertex $v\in V$ with order $r(v)$ indicating its importance in $G$. In this paper, we use boundary-first ordering (Liu et al., 2022a) to generate the order for the PSP indexes. For non-partitioned SP indexes, we leverage degree to decide the order of small-world networks while the minimum degree elimination-based ordering (Berry et al., 2003) for the road networks. A path $p=\left<v_{0},v_{1},\dots,v_{k}\right>((v_{i},v_{i+1})\in E,0\leq i<k)$ is a sequence of adjacent vertices with length of $l(p)=\sum_{i=0}^{k-1}w(v_{i},v_{i+1})$. Given an origin $s$ and destination $t$ vertex (OD pair), the shortest path query $Q(s,t)$ calculates the path with the minimum length between them, i.e. $p_{G}(s,t)$ with the shortest distance $d_{G}(s,t)$. In the following, we introduce the PSP index dimensions and summarize each of them.
2.1. Dimensions of PSP Index
Based on a comprehensive literature review and analysis of the PSP index, we identify three critical dimensions of the PSP index: 1) Graph Partition Method, which divides the graph into multiple subgraphs and is the key difference from the standalone SP indexes. However, the existing partition methods are not designed for path indexes, so how they affect the PSP index is unknown and deserves study; 2) PSP Strategy, which provides a procedure to organize the index construction for multiple subgraphs and coordinate those indexes such that the query correctness is guaranteed. Only straightforward solutions exist currently as will be discussed in Section 2.3. However, this part is crucial because it determines the complexity of both the query and update; 3) SP Algorithm, whose influence was thoroughly studied in our previous work (Zhang et al., 2021a). With these three dimensions, we can form a space to position the existing methods into it. Given a massive dynamic network, these dimensions provide a roadmap to construct a PSP index by selecting a specific graph partition method, SP algorithm and PSP strategy such that the application requirement (query/update efficiency) can be satisfied. Next, we discuss and review these three dimensions.
2.2. Graph Partitioning
Definition 0 (Graph Partitioning).
A graph $G$ is decomposed into multiple disjoint subgraphs $\{G_{i}\}$ $(i=1,2,\dots,k)$ with $V_{i}\cap V_{j}=\emptyset$ and $\bigcup V_{i}=V$. denoted as $E_{inter}$ and $E_{intra}$.
We now discuss and categorize the vertex and edge in graph partitioning. $\forall v\in G_{i}$, we say $v$ is a boundary vertex if there exists a neighbor of $v$ in the another subgraph, that is $\exists u\in N(v),u\in G_{j}(i\neq j)$. Otherwise, $v$ is an inner vertex. We represent the boundary vertex set of $G_{i}$ as $B_{i}$ and that of $G$ as $B=\bigcup B_{i}$. For $(u,v)\in E$, we say $(u,v)$ is an inter-edge if both its two endpoints $u$ and $v$ are boundary vertices from different subgraphs, i.e., $u\in B_{i},v\in B_{j},i\neq j$. Otherwise, it is an intra-edge. The corresponding edge sets are denoted as $E_{inter}$ and $E_{intra}$, respectively. For example in Figure 2, the example graph is partitioned into four subgraphs $G_{1}$ to $G_{4}$, with the red vertices denoting the boundaries ($B_{1}=\{v_{3},v_{10}\}$) and the white vertices representing the inner vertices. Edges outside partitions like $(v_{3},v_{5})$ are inter-edges, and $(v_{3},v_{10})$ is an intra-edge. Besides, the boundary vertex plays an important role in the shortest path computation with its cut property: given a shortest path $p(s,t)$ with its endpoints in two different subgraphs $s\in G_{i},t\notin G_{i}$, there is at least one vertex $v$ which is in the boundary vertex set of $G_{i}$ on the path, that is $\exists v\in p(s,t),v\in B_{i}$. Therefore, boundary vertices would determine the correctness and complexity of partitioned shortest path computation (discussed in the following sections). Various partition methods are involved in the existing PSP methods such as METIS (Karypis and Kumar, 1998) in G-Tree (Zhong et al., 2013) and SHARC (Bauer and Delling, 2010) ; PUNCH (Delling et al., 2011) in COLA (Wang et al., 2016), FHL (Liu et al., 2021, 2022b), FHL-Cube (Liu et al., 2022a) and THop (Li et al., 2019c, 2020b) ; Tree Decomposition-based Partition (Ouyang et al., 2018) in Core-Tree (Li et al., 2020a, 2019b) ; multi-level partitioning (Jung and Pramanik, 1996) in G-tree (Zhong et al., 2013, 2015), SPAH (Jung and Pramanik, 1996), CPR (Delling et al., 2017) and (Holzer et al., 2009; Chondrogiannis and Gamper, 2016, 2014). But, their selections are performed in an ad-hoc manner or consider the boundary number, partition balance, and spatial awareness, rather than the SP computation itself.
2.3. Traditional PSP Strategy
The PSP index $L$ typically consists of two components: the partition indexes $\{L_{i}\}$ for each subgraph (partition) $G_{i}$, and the overlay index $\tilde{L}$ for the overlay graph which is composed of the boundary vertices of all partitions, i.e., $L=\{L_{i}\}\cup\tilde{L}$. Regardless of the specific SP algorithm, almost all the existing PSP indexes (Jung and Pramanik, 2002; Delling et al., 2017; Holzer et al., 2009; Chondrogiannis and Gamper, 2014, 2016; Zhong et al., 2013; Li et al., 2019a; Zhong et al., 2015; Ta et al., 2017; Wang et al., 2019, 2016; Liu et al., 2021, 2022b, 2022a; Li et al., 2019c, 2020b; Li et al., 2020a; Zheng et al., 2022; Wang et al., 2021; Farhan et al., 2022; Gubichev et al., 2010; Qiao et al., 2012) construct the PSP index and process the queries under the following procedure.
Figure 2. Example Graph $G$ and Overlay Graph $\tilde{G}$
Index Construction. It first leverages a partition method to divide the graph $G$ into multiple subgraphs $\{G_{i}\}$ and then builds the PSP index $L$ by the following four steps. As illustrated in Figure 3, Step 1 precomputes the global distance between all boundary vertex pairs $(b_{i1},b_{i2}),(b_{i1},b_{i2}\in B_{i})$ for each partition $G_{i}$ and insert shortcuts $e(b_{i1},b_{i2})=d_{G}(b_{i1},b_{i2})$ into $G_{i}$ to get $G^{\prime}_{i}$. For instance, the global distance between boundary vertex pairs $(v_{3},v_{10})\in G_{1}$, $(v_{8},v_{9})\in G_{2}$, $(v_{5},v_{12})\in G_{3}$, $(v_{6},v_{11})\in G_{4}$ are calculated and inserted into their corresponding subgraphs to form new subgraphs $G^{\prime}_{1},G^{\prime}_{2},G^{\prime}_{3},G^{\prime}_{4}$, respectively; Step 2 constructs the SP index $L_{i}$ based on $G^{\prime}_{i}$. The index can be any type like CH (Geisberger et al., 2008), PLL (Akiba et al., 2013) or H2H (Ouyang et al., 2018), and the index construction procedure among multiple partitions can be parallelized as it takes each partition as input independently; Step 3 constructs the overlay graph $\tilde{G}$ based on the precomputed shortcuts in Step 1. Specifically, $\tilde{G}$ is composed of shortcuts between boundary vertex pairs $\{B_{i}\times B_{i}\}$ ($(v_{3},v_{10}),(v_{8},v_{9}),(v_{5},v_{12}),(v_{6},v_{11})$) and the inter-edge set $E_{inter}$ ($(v_{3},v_{5})$, $(v_{10},v_{12})$, $(v_{9},v_{10})$,$(v_{9},v_{12})$, $(v_{9},v_{11})$, $(v_{8},v_{11})$, $(v_{12},v_{11})$, $(v_{12},v_{6})$), that is $V_{\tilde{G}}=\{B_{i}\},E_{\tilde{G}}=\{B_{i}\times B_{i}\}\cup E_{inter}$; Step 4 constructs the SP index $\tilde{L}$ on $\tilde{G}$. Similar to $L_{i}$, $\tilde{L}$ can be any index type.
Figure 3. Traditional PSP Strategy Example
Note that the construction of $\tilde{L}$ (Step 3, Step 4) can be parallelized with $L_{i}$ (Step 2) since they are independent and both rely on Step 1. All pairs of boundaries are independent of each other so we only need $|B|$ times of Dijkstra’s with time complexity $O(nlogn+m)$. Then each partition’s label $L_{i}$ can be constructed in parallel, and $\tilde{L}$’s construction is also independent to them, so its complexity is the worst case of them: $max\{O_{c}(G_{i}),O_{c}(\tilde{G})\}$, where $O_{c}$ is the complexity of underlying index’s construction time as our discussion is not fixed to any specific index type.
Query Processing. By referring to the cut property of boundary vertex, the shortest path queries are divided into two categories and processed by utilizing $L$:
Case 1: Same-Partition i.e., $\forall s,t\in G_{i}$, $Q(s,t)=d_{L_{i}}(s,t)$;
Case 2: Cross-Partition, i.e., $\forall s\in G_{i},t\in G_{j}(i\neq j),Q(s,t)=$
$\begin{cases}d_{\tilde{L}}(s,t)&s,t\in B\\ \min\limits_{b_{q}\in B_{j}}\{d_{\tilde{L}}(s,b_{q})+d_{L_{j}}(b_{q},t)\}&s\in B,t\notin B\\ \min\limits_{b_{p}\in B_{i}}\{d_{L_{i}}(s,b_{p})+d_{\tilde{L}}(b_{p},t)\}&s\notin B,t\in B\\ \min\limits_{b_{p}\in B_{i},b_{q}\in B_{j}}\{d_{L_{i}}(s,b_{p})+d_{\tilde{L}}(b_{p},b_{q})+d_{L_{j}}(b_{q},t)\}&s\notin B,t\notin B\\ \end{cases}$
In summary, when $s$ and $t$ are in the same partition, we can use $L_{i}$ to answer $d_{G}(s,t)$; otherwise, we have to use $L_{i},L_{j}$ and $\tilde{L}$. The same-partition query complexity is $O_{q}(L_{i})$, where $O_{q}$ is the index’s query complexity. The cross-partition query is made up of three parallel query sets, and the complexity is the worst of them: $max\{B_{i}\times O_{q}(L_{i}),B_{i}\times B_{j}\times O_{q}(\tilde{L}),B_{j}\times O_{q}(L_{j})\}$. The proof of the query processing can be found in the Appendix of our extended version (Zhang et al., 2023b).
2.4. Shortest Path Algorithms
We summarize the shortest path algorithms as follows: 1) Direct Search such as Dijkstra’s (Dijkstra, 1959) and $A^{*}$ (Hart et al., 1968) searches the graph when the index cannot be constructed due to the large graph size (QbS (Wang et al., 2021), ParDist (Chondrogiannis and Gamper, 2016)) or complicated problem (COLA (Wang et al., 2016)). It takes no time in index update, but queries are slowest; 2) Contraction Hierarchy (CH) (Geisberger et al., 2008) is a widely used lightweight index that contracts the vertices in a pre-defined order and preserves the shortest path information by adding shortcuts among the contracted vertex’s neighbors. The search-based CH (Geisberger et al., 2008) has a small index size but takes longer to build and maintain (Zhang et al., 2021a; Geisberger et al., 2012) while concatenation-based CH (Ouyang et al., 2020; Wei et al., 2020) is much faster to build and maintain if the tree-width is small. CH’s query performance is generally around 10$\times$ faster than the direct search but much slower than the hub labelings. Surprisingly, no PSP index has used CH as underlying index; 3) Pruned Landmark Labeling (PLL). Although many hub labeling methods have been proposed in the past decade, only two are widely used. Built by either pruned search (Akiba et al., 2013) or propagation (Li et al., 2019b; Jin et al., 2020; Zhang et al., 2021b, 2023a), PLL is the only index that can work on the graph with large treewidth. Therefore, graphs with this property use it as the underlying index (e.g., QsB (Wang et al., 2021; Farhan et al., 2022) for its landmarks, Core-Tree (CT) (Li et al., 2020a; Zheng et al., 2022) for its core); 4) Tree Decomposition (TD) (Ouyang et al., 2018; Chang et al., 2012; Wei, 2012; Chen et al., 2021; Zhang et al., 2021c). As another widely used hop labeling, it is much faster than PLL for graphs with smaller treewidth but cannot scale to large ones. For instance, CT (Li et al., 2020a; Zheng et al., 2022) uses it for its periphery, and FHL (Liu et al., 2021, 2022a) use it as a forest; 5) All Pair Pre-computation pre-computes all-pair distance with the fastest query, but incurs long index construction and large index storage. G-Tree (Zhong et al., 2013, 2015) and ROAD (Lee et al., 2009, 2010) store the distance between boundary vertices in each partition.
3. Novel PSP Strategies
As stated in Section 2, the traditional PSP strategy can be time-consuming for large networks since it requires a laborious computation of all-pair boundary distances using Dijkstra’s search. To tackle this problem, we propose two novel PSP strategies and analyze how they could be leveraged to correctly answer the shortest path queries. Besides, we design the index update approaches systematically to support the dynamic application of both the existing and novel PSP strategies. Lastly, we proposed a pruning technique to slim the overlay graph to achieve better index performance. It should be noted that the PSP strategy is a general procedure that any PSP index needs to follow and can work with any type of graph partition method or SP algorithm. So we introduce the PSP strategies without mentioning the specific partitioning and SP approach.
3.1. No-Boundary and Post-Boundary Strategy
Since the traditional partitioned index starts by precomputing the all-pair distance among boundary vertices, we call this approach Pre-Boundary Strategy. It seems that this strategy, with the theoretical guarantee for correct partitioned shortest-path processing, could handle all the partitioned queries well. Nevertheless, the first step of the Pre-Boundary could be very time-consuming because only the index-free SP algorithms like Dijkstra’s (Dijkstra, 1959) or A (Hart et al., 1968) could be utilized. As a result, the index construction and maintenance efficiency suffer when the graph has numerous boundary vertex pairs. However, it appears that pre-computing the all-pair boundary distance is an essential procedure for constructing the “correct” PSP index, otherwise the correctness of the subgraph index $L_{i}$ cannot be guaranteed because the shortest distance between boundary vertices within one partition could pass through another partition. For instance, as shown in Figure 3, $p(v_{6},v_{11})=\left<v_{6},v_{12},v_{11}\right>$ goes outside $G_{4}$ and passes through $v_{12}\in G_{3}$. As a result, $d(v_{6},v_{11})$ cannot be answered correctly only with $G_{3}$ and the global shortest distance $d(v_{6},v_{11})$ calculation is essential before constructing $L_{3}$. Then a question naturally arises: do we really require precomputing the boundary all-pair distance first? In other words, does there exist a chance to keep index correctness by dropping this time-consuming step? Based on this brainstorm, we break the traditional misconception and propose a novel No-Boundary Strategy as follows, which significantly reduces the index construction/maintenance time by skipping the time-consuming pre-computation step.
Index Construction. It contains three steps as shown in Figure 4. With a graph partitioned into subgraphs $\{G_{i}\}$, we start by constructing the shortest path index $L_{i}$ for each subgraph $G_{i}$ parallelly in Step 1; Then Step 2 construct the overlay graph $\tilde{G}$ based on $\{L_{i}\}$. For instance, $\tilde{G}$ is made up of boundary all-pair edges $(v_{3},v_{10})$, $(v_{8},v_{9})$, $(v_{5},v_{12})$, $(v_{6},v_{11})$ (with their weight calculated based on $L_{1},L_{2},L_{3}$, and $L_{4}$) and inter-edges. It should be noted that although the overlay graph $\tilde{G}$ has the same structure as that in the Pre-boundary Strategy, their edge weights differ as $\{L_{i}\}$ of them are constructed from different subgraphs; Step 3 constructs the shortest path index $\tilde{L}$ for $\tilde{G}$. Since the partition indexes $\{L_{i}\}$ are constructed in parallel first followed by the construction of $\tilde{L}$, the No-Boundary takes $max\{O_{c}(G_{i})\}+O_{c}(\tilde{G})$ time in index construction.
Figure 4. No-Boundary and Post-Boundary Strategies
Query Processing. Now that $L_{i}$ cannot answer $G_{i}$’s query correctly as the global distance between boundary all-pair is not covered, then how can No-Boundary answer query correctly? Before revealing it, we first prove that the correctness of $\tilde{L}$ still holds even though it is built upon the incorrect $\{L_{i}\}$, that is the queries between boundary vertices can be correctly processed with $\tilde{L}$, using:
Theorem 1.
$\forall s,t\in B,d_{G}(s,t)=d_{\tilde{L}}(s,t)$.
Proof.
We divide all the scenarios into three cases as shown in Figure 5-(a): 1) $s,t\in G_{i}$ and $p_{G}(s,t)$ only passes through the interior of $G_{i}$, then it is obvious that $d_{G}(s,t)=d_{G_{i}}(s,t)$ holds. Since $\tilde{w}(s,t)=d_{G_{i}}(s,t)$ is leveraged for $\tilde{G}$’s construction, we can obtain that $d_{G}(s,t)=d_{\tilde{G}}(s,t)=d_{\tilde{L}}(s,t)$; 2) $s,t\in G_{i}$ with $p_{G}(s,t)$ going outside of $G_{i}$; 3) $s\in G_{i}$ and $t\in G_{j}(i\neq j)$. In the latter two cases, we take the concise form of $p_{G}(s,t)$ by extracting only the boundary vertices as $p_{c}=\left<s,b_{0},b_{1},\dots,b_{n},t\right>$ ($b_{i}\in B,0\leq i\leq n$). For two adjacent vertices $b_{i},b_{j}\in p_{c}$, if $b_{i}$ and $b_{j}$ are in the same partition, then $\tilde{w}(b_{i},b_{j})$ can be correctly obtained in $\tilde{G}$ via case 1); otherwise, $(b_{i},b_{j})$ is an inter-edge with $\tilde{w}(b_{i},b_{j})=$$w(b_{i},b_{j})$ naturally correct. Thus, the shortest distance among boundary vertices can be correctly calculated by accumulating the edge weights on $\tilde{G}$. ∎
Figure 5. Different Categories of OD Distribution
Based on the above Theorem, we continue to answer the shortest distance queries by discussing all scenarios in two cases:
Case 1: Same Partition $s,t\in G_{i}$, we report $Q(s,t)$ as follows:
Theorem 2.
[…]
Proof.
We denote $d_{L_{i}}(s,t)$ as $d_{2}$, $\min\{d_{L_{i}}(s,b_{i1})+d_{\tilde{L}}(b_{i1},b_{i2})+d_{L_{i}}(b_{i2},t)\}$ as $d_{4}$, and divide all the scenarios into two subcases:
Subcase 1: $p_{G}(s,t)$ does not go outside of $G_{i}$, as shown in the first row of Figure 5. No matter $s$ and $t$ are borders or not, $d_{L_{i}}(s,t)$ (i.e., $d_{2}$) is enough to answer $d_{G}(s,t)$ as $L_{i}$ is built based on $G_{i}$ which contains all necessary information for finding the shortest path.
Subcase 2: $p_{G}(s,t)$ passes outside of $G_{i}$, as shown in the second row of Figure 5. If $s$ and $t$ are both boundary vertices, $d_{G}(s,t)=d_{\tilde{L}}(s,t)$ holds by referring to Theorem 1. If $s$ and $t$ are both non-boundary vertices, we take the concise form of $p_{G}(s,t)$ by extracting only the boundary vertices as $p_{c}=\left<s,b_{0},b_{1},\dots,b_{n},t\right>$ ($b_{i}\in B,0\leq i\leq n$). Therefore, $d_{4}$ can correctly handle this case as the $d_{G}(b_{0},b_{n})$ can be answered by $d_{\tilde{L}}(b_{0},b_{n})$ as per Theorem 1, while $d_{G}(s,b_{0})$ and $d_{G}(b_{n},t)$ can be answered by $d_{L_{i}}(s,b_{0})$ and $d_{L_{i}}(b_{n},t)$ by referring to Case 1. If either $s$ or $t$ is a non-boundary vertex, its distance is the special case of $d_{4}$ and can be easily proved. ∎
Case 2: Cross-Partitions, i.e., $s\in G_{i},t\in G_{j},i\neq j$, we process queries the same manner as that of Pre-boundary Strategy.
Lemma 3.
The cross-partition queries can be correctly processed in the No-boundary Strategy.
Note that the proof of Lemma 3 and all the other Lemmas of this paper are available in the Appendix of our extended version (Zhang et al., 2023b). We can use $\tilde{L}$ to answer $d_{G}(s,t)$ when $s$ and $t$ are both boundary vertices; otherwise, we have to use $L_{i},L_{j}$ and $\tilde{L}$. The intra-query complexity is $max\{O_{q}(L_{i}),B_{i}\times B_{j}\times O_{q}(\tilde{L}),B_{j}\times O_{q}(L_{j})\}$, while the inter-query takes $O_{q}(\tilde{L})$ time when $s$ and $t$ are both boundary vertices, and the complexity of other case is the worst of them: $max\{B_{i}\times O_{q}(L_{i}),B_{i}\times B_{j}\times O_{q}(\tilde{L}),B_{j}\times O_{q}(L_{j})\}$. Therefore, we make a bold attempt to enhance the performance of PSP index by skipping the heavy pre-computation, such that the index construction time is largely reduced so as the index maintenance (as will introduced in Section 3.2). Surprisingly, with incorrect distance value in the No-boundary index, we prove theoretically that it can support the query answering with correctness guarantee.
By comparing the query processing procedure of the traditional Pre-boundary strategy and our proposed No-boundary strategy, we find that it is more complex for No-boundary to process those queries with endpoints in the same partition because of the incorrectness of $\{L_{i}\}$. We repair this weakness by fixing the incorrect boundary all-pair distance in each partition $G_{i}$ and then transforming $\{L_{i}\}$ to their correct version. Specifically, we calculate the boundary all-pair distance $d(b_{i1},b_{i2})$ within each partition $G_{i}$ by leveraging $\tilde{L}$ and insert its corresponding edge $e(b_{i1},b_{i2})$ into $G_{i}$ to get $G^{\prime}_{i}$. For instance, $d(v_{3},v_{10})$, $d(v_{8},v_{9})$, $d(v_{5},v_{12})$ and $d(v_{6},v_{11})$ are calculated based on $\tilde{L}$ and these edges (as colored red in Figure 4) are inserted into their corresponding partitions with new subgraphs $\{G^{\prime}_{i}\}$ formed. Followed by, we treat those newly inserted edges as graph updates and refresh $\{L_{i}\}$ to $\{L^{\prime}_{i}\}$ by invoking the index maintenance algorithms (Ouyang et al., 2020; Zhang et al., 2021c; Zhang and Yu, 2022; Farhan et al., 2022) in each partition parallelly. As these procedures are implemented after the No-boundary, we call this strategy as Post-boundary. By referring to Theorem 1, the correct global distance information is contained in $\{G^{\prime}_{i}\}$, so $\{L^{\prime}_{i}\}$ can support the correct query processing of $Q(s,t)$ with $s,t\in G_{i}$. With the consistent query procedure as that of Pre-boundary, the query efficiency of Post-boundary reaches the state-of-the-art level.
3.2. Index Update for Different PSP Strategies
The partitioned shortest path has been widely used, however, how to maintain the index to support its application in dynamic environments has never been discussed. In this section, we propose the partitioned index update algorithms to guarantee the correctness of the partitioned index in dynamic scenarios.
Index Update for Pre-boundary Strategy. As shown in Figure 6-(a), when the weight of edge $e\in G$ changes, we first recalculate the boundary-all pair distance (as the step 1 in index construction) and identify the changed weight $e(b_{i1},b_{i2})$ between boundary vertices in each partition $G_{i}$. Then we need to update the corresponding partition index $L_{i}$ and overlay index $\tilde{L}$.
Lemma 4.
The indexes of Pre-Boundary Strategy can be correctly updated with the above strategy.
Step 1 takes $O_{u}(nlogn+m)$ time, while the partition and $\tilde{G}$ index takes $max\{O_{u}(G_{i}),O_{u}(\tilde{G})\}$ time to update in parallel, where $O_{u}$ is the update complexity for different indexes.
Figure 6. Index Update of Different PSP Strategies
Index Update for No-boundary Strategy. Since the weight change of inter-edges does not affect the index for each subgraph, we divide index updates into two scenarios as shown in Figure 6-(b). Scenario 1: Inter-edge weight change. When $e\in E_{inter}$ changes, only $\tilde{L}$ needs an update; scenario 2: Intra-edge weight change. When $e\in E_{intra}$ ($e\in E_{j}$) changes, we first update $L_{j}$ and compare the old and new weights of $e(b_{j1},b_{j2})$ between boundary in $G_{j}$. If there is an edge weight update, we need to further update $\tilde{L}$. The update complexity is $max\{O_{u}(G_{i})\}+O_{u}(\tilde{G})$.
Lemma 5.
The indexes of No-Boundary can be correctly maintained with the above update strategy.
Index Update for Post-boundary Strategy. It is similar to No-Boundary, with an additional judgment and processing as shown in Figure 6-(b). Scenario 1: Intra-edge weight change. Suppose $e\in E_{i}$ changes, we update $G_{i},L_{i}$ and then update $\tilde{G},\tilde{L}$ if any $d_{L_{j}}(b_{j1},b_{j2})$ changes. After that we may need to further update $\{G^{\prime}_{i}\},\{L^{\prime}_{i}\}$ if $d_{G_{i}^{\prime}}(d_{i1},d_{i2})$ and $d_{\tilde{L}}(d_{i1},d_{i2})$ are different; scenario 2: Inter-edge weight change. Suppose $e\in E_{inter}$ changes, we update $\tilde{G},\tilde{L}$ and then update $\{G_{i}^{\prime}\},\{L_{i}^{\prime}\}$. Therefore, the update time complexity is the linear combination of the above procedures: $max\{O_{u}(G_{i})\}+O_{u}(\tilde{G})+O_{q}(\tilde{G})+max\{O_{u}(G_{i})\}$.
Lemma 6.
The indexes of Post-Boundary Strategy can be correctly maintained with the above update strategy.
Remark 1. To construct the PSP index, Pre-Boundary precomputes the all-pair shortest distance among boundary vertices of each partition, and thus the newly added edges of $G_{i}$ and $\tilde{G}$ (red edges) are of correct edge weight. For example, the edge weight of $e(v_{5},v_{12})$ is 6 which is the path length of the shortest path $p_{G}(v_{5},v_{12})=\left<v_{5},v_{3},v_{10},v_{12}\right>$. While No-Boundary skips the heavy precomputation and only leverage $L_{i}$ to calculate boundary all-pair distance for overlay graph $\tilde{G}$ construction. As a result, the edge weight $w_{\tilde{G}}(v_{5},v_{12})=7$ is incorrect since it only preserves the local distance with corresponding lcoal shortest path $p(v_{5},v_{12})=\left<v_{5},v_{4},v_{12}\right>$ in $G_{3}$. This innovation largely saves the index construction time while drags down the query answering speed. Then Post-Boundary makes up the query efficiency loss by inserting the correct shortest distance value between boundary pairs into each partition (e.g., insert $w(v_{5},v_{12})=6$ to $G^{\prime}_{3}$) and refresh $\{L_{i}\}$ to its correct version $\{L^{\prime}_{i}\}$. In summary, there exists a balance of index performance among the three boundary strategies. In terms of the query processing, the No-Boundary is slower than Pre-Boundary and Post-Boundary, since the boundary vertices and path concatenation should be considered for OD within one partition. In terms of the index construction and update, the No-Boundary and Post-Boundary enjoy faster speed as the boundary all-pair distance computation can be neglected. In terms of the index size, the Post-Boundary needs twice the storage space as that of No-Boundary and Pre-Boundary, even though it has advantages in both query processing and index update. Therefore, no one strategy is better than the others in all aspects and could be selected based on the application scenario. Besides, it is worth noting that we illustrate PSP strategies in the form of edge-cut for clear presentation since the vertex-cut partitions can be easily converted to edge-cut partitions. Due to limited space, the conversion method is presented in extended version (Zhang et al., 2023b).
3.3. Pruning-based Overlay Graph Optimization
Though it is flexible to select the suitable boundary strategy according to application scenarios, we observe that all boundary strategies are hindered by the extremely dense overlay graph. The density of overlay graph $\tilde{G}$ increases dramatically as the all-pair boundary vertices in each partition are connected during its construction, which badly affects the index performance by slowing down the index construction, query processing, and index update (Zhang et al., 2021a). Then is it possible to improve index performance and decrease the density of $\tilde{G}$ by deleting some unnecessary edges in $\{B_{i}\times B_{i}\}$? As proved in Theorem 1, the shortest distance between boundary vertices is well-preserved in $\tilde{G}$, then could the distance still be preserved in $\tilde{G}$ if some edges are removed? We start exploring this question by categorizing the boundary vertices as followed:
Definition 0 (Half/Full-Connected Boundary Vertex).
$b\in B_{i}$ is a half-connected boundary vertex if $\exists u\in N(b),u\in G_{i},u\notin B_{i}$ and we denote them as $B_{i}^{H}$ with $B^{H}$=$\{B_{i}^{H}\}$; otherwise, $b$ is a full-connected boundary vertex and denoted as $B_{i}^{F}$ with $B^{F}$=$\{B_{i}^{F}\}$.
Figure 7. Example of Overlay Graph Simplification
As illustrated in Figure 7-(a), $b_{i3}$ is a half-connected boundary vertex since there exists $v_{i1}\in N(b_{i3}),v_{i1}\notin B_{i}$; and $b_{i1},b_{i2},b_{i4}$ are full-connected boundary vertices. Initially, $\tilde{G}$ can be viewed as composed of multiple complete graphs $\{B_{i}\times B_{i}\}$, which are connected by the inter-edges. Within this dense structure, the global distance between any two boundary vertices $s,t\in B$ is well-preserved. Let us consider a shortest path $p$ with starting vertex $b_{s}\in B_{i}^{F}$ and target vertex $b_{t}\in B$, then the second vertex $v_{0}$ along $p$ is a boundary vertex with $v_{0}\in B,v_{0}\in N(b_{s})$ according to the definition above. Therefore, the global distance information of $p$ will still be kept in $\tilde{G}$ if we delete edges $\{(b_{s},b_{i})\}$ with $b_{i}\in B_{i},b_{i}\notin N(b_{s})$, since it is impossible for $p$ to pass through these edges. The same principle holds for a shortest path ending with a fully connected boundary vertex $b_{t}\in B_{i}^{F}$ that its distance will also be preserved in $\tilde{G}$ after deleting edges $\{(b_{i},b_{t})\}$ with $b_{i}\in B_{i},b_{i}\notin N(b_{t})$. Then for the shortest path starting and ending with the half-connected boundary vertices $b_{s},b_{t}\in B_{i}^{H}$, it is necessary to insert an edge $(b_{s},b_{t})$ into $\tilde{G}$ to preserve its distance information, as the second vertex on this path could be a non-boundary vertex which is not contained in the overlay graph. As analyzed above, we can shrink $\tilde{G}$ ($\{B_{i}\times B_{i}\}\cup E_{inter}$) to its slimmer version $\tilde{G}^{\prime}$ (composed of $\{B_{i}^{H}\times B_{i}^{H}\}$, $E_{inter}$ and $B_{i}^{F}$’s adjacent edges). Suppose $b_{i1},b_{i2}\in N(b_{i4})$ and let us only take $G_{i}$ as an example for the clear presentation, the original overlay graph as shown in Figure 7-(b) will be shrunk to a sparser one in Figure 7-(c) after the pruning technique. The following lemma demonstrates that the global distance can still be preserved in the overlay graph with the pruning technique.
Lemma 8.
$d_{G}(b_{s},b_{t})=d_{\tilde{G^{\prime}}}(b_{s},b_{t})$, $\forall(b_{s},b_{t})\in B\times B$.
Therefore, we can enhance the index performance by pruning some edges in $\tilde{G}$ and utilizing its sparser version $\tilde{G}^{\prime}$, while preserving the global distance by referring to Lemma 8.
4. SP-Oriented Partition Method Classification
Having figured out how to achieve efficient query and update in partitioned graphs while guaranteeing the index correctness with the novel PSP strategies, we move on to the remaining two dimensions. As the SP index is well studied and analyzed in Section 2.4. Then, only the mist on the partition method remains: how to choose a partition method to achieve our preferred performance? This is not trivial because: 1) dozens of partition methods with different characteristics were proposed and applied in the past decades; 2) they are classified under different criteria (Buluç et al., 2016), for example, from the perspectives of partition manners (spectral, flow, graph glowing, contraction and multi-level), partition objectives (balance and minimal cut), computation manner (in-memory, distributed and streaming), and cut category (edge-cut and vertex-cut), but it is not clear which criterion is beneficial to SP index; 3) their relationship with (or the influence on) SP index is unknown and has never been studied. Based on these doubts, we propose a novel SP-oriented classification of partition methods as follows:
1) Planar Partition treats partitions equally on one level. It decomposes a graph $G$ into multiple equally-important subgraphs $\{G_{i}|1\leq i\leq k\}(k\in[2,\infty])$ with 1) $\bigcup_{i\in[1,k]}V(G_{i})=V$, $V(G_{i})\cap V(G_{j})=\emptyset$, or 2) $\bigcup_{i\in[1,k]}E(G_{i})=E$, $E(G_{i})\cap E(G_{j})=\emptyset,\forall i,j\in[1,k],i\neq j$. Representative methods include spectral partitioning (Pothen et al., 1990; Barnard and Simon, 1994), graph growing (Diekmann et al., 2000), flow-based partitioning (Sanders and Schulz, 2011; Delling et al., 2011), node-swapping (Kernighan and Lin, 1970; Fiduccia and Mattheyses, 1982), multilevel graph partition (MGP) (Karypis and Kumar, 1998; Barat et al., 2018), etc..
2) Core-Periphery Partition ((Borgatti and Everett, 2000; Rombach et al., 2014; Elliott et al., 2020; Maehara et al., 2014a)) treats the partitions discriminately by taking some important vertices as “Core” and the remaining ones as “Peripheries”. It decomposes a graph $G$ into two distinct parts: a core subgraph $G_{c}$ and $k-1$ peripheral subgraphs $\{G_{i}|1\leq i\leq k-1\}(k\in[2,\infty])$, satisfying $\bigcup_{i\in[1,k-1]}V(G_{i})\cup V(G_{c})=V,V(G_{c})\cap V(G_{i})=\emptyset(\forall 1\leq i\leq k-1)$ and $V(G_{i})\cap V(G_{j})=\emptyset(\forall i,j\in[1,k-1],i\neq j)$. There are two big streams depending on how the core is formed: Core-Tree Decomposition (Maehara et al., 2014b; Li et al., 2020a; Zheng et al., 2022) forms the core through tree decomposition (Robertson and Seymour, 1984), and the resulting periphery part is a set of small-width trees. Specifically, it leverages minimum degree elimination (Xu et al., 2005) to contract vertices of lower degree first such that the graph is contracted from the edge towards the center, generating a set of growing trees (“Peripheries”) around a shrinking while denser graph (“Core”) formed by those non-contracted vertices. The contraction terminates when the width of one tree reaches the previously set threshold; Sketch (Potamias et al., 2009; Das Sarma et al., 2010; Gubichev et al., 2010; Tretyakov et al., 2011; Qiao et al., 2012; Qi et al., 2013; Farhan et al., 2018; Wang et al., 2021; Farhan et al., 2022) selects a set of vertices as landmarks, which can be regarded as the core and treats the remaining parts as periphery. It generally works on extremely large graphs where the above tree decomposition is impossible.
3) Hierarchical Partition organizes the network partitions hierarchically and each level is equivalent to planar partitioning. It organizes a graph $G$ hierarchically within $H$ levels where each level $h$ is a planar partitioning $\{G^{h}_{i}\}$ of $G$ and $\exists G^{h-1}_{j}\supset G^{h}_{i},\forall G^{h}_{i}(1\leq h<H)$. Classic methods include HiTi (Jung and Pramanik, 1996), SHARC (Bauer and Delling, 2010), G-Tree (Zhong et al., 2015), G*-Tree (Li et al., 2019a), SMOG (Holzer et al., 2009), and so on.A typical approach to create hierarchical partitions is by utilizing MGP, such as METIS (Karypis and Kumar, 1998) in G-Tree and G*-Tree.
Figure 8. Different SP-Oriented Partitioning
Table 1. SP-Oriented Partition Methods Classification.
Partition Structure Partition Method Cut Category Partition Objectives Spatial-Aware
Category Representative Methods Edge Cut Vertex Cut Balance Minimal Cut
Planar Spectral Partitioning SP (Pothen et al., 1990), RSB (Barnard and Simon, 1994) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
Graph Growing Bubble (Diekmann et al., 2000) $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
Flow-Based Partitioning PUNCH (Delling et al., 2011) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Geometric Partitioning RCB (Simon, 1991; Huang et al., 1996) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Node-Swapping Heuristic KL (Kernighan and Lin, 1970), FM (Fiduccia and Mattheyses, 1982) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
Minimum $k$-cut $k$-cut (Goldschmidt and Hochbaum, 1994) $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
Multilevel Graph Partitioning SCOTCH (Pellegrini and Roman, 1996), METIS (Karypis and Kumar, 1998), KaHyPar (Akhremtsev et al., 2017) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
Streaming Edge Partitioning HDRF (Petroni et al., 2015; Mayer and Jacobsen, 2021), CLUGP (Kong et al., 2022) $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
In-Memory Edge Partitioning NE (Zhang et al., 2017), HEP (Mayer and Jacobsen, 2021) $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
Core-Periphery Core-Tree Decomposition Core-Tree (Maehara et al., 2014b; Li et al., 2020a; Zheng et al., 2022; Elliott et al., 2020) $\checkmark$ $\times$ $\times$ $\times$ $\times$
Sketch Sketch (Potamias et al., 2009; Das Sarma et al., 2010; Gubichev et al., 2010; Tretyakov et al., 2011; Qiao et al., 2012; Qi et al., 2013; Farhan et al., 2018; Wang et al., 2021; Farhan et al., 2022) $\checkmark$ $\times$ $\times$ $\times$ $\times$
Hierarchical Edge-Cut Based HP HiTi (Jung and Pramanik, 1996) $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$
SHARC (Bauer and Delling, 2010), G-Tree (Zhong et al., 2015), G*-Tree (Li et al., 2019a) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
Vertex-Separator Based HP SMOG (Holzer et al., 2009), ROAD  (Lee et al., 2009, 2010) $\times$ $\checkmark$ $\times$ $\times$ $\times$
Figure 8 shows the partition results of an example graph $G$ with different partitioning methods, where the red vertices represent the corresponding boundary vertices. Specifically, planar partitioning generates four partitions with equivalent vertex size while core-periphery partitioning produces both “Core” and “Periphery” (the periphery of Core-Tree decomposition is a set of small-width trees while Sketch treats the non-core part as periphery. The yellow vertex in each periphery is the “root vertex”). Hierarchical partitioning (HP) organizes partitions hierarchically, and the leaf nodes (the lowest level partition result) may have the same partition result as planar partitioning if the same partitioning method is used. In addition, we summarize and categorize existing partition methods under our SP-oriented classification, as shown in Table 1.
Remark 2. As analyzed in Section 3, the boundary number of each partition will affect the partition method section as it is crucial to the computation complexity of PSP index. Specifically, the planar and hierarchical partitions have no limit to the boundary number, even though reducing the cut size is one of their optimization goals as the path index may suffer from large boundary numbers. Complementary, the core-tree decomposition limits the boundary number to be no larger than the pre-defined bandwidth such that their performance would not be deteriorated by the boundary if we set the bandwidth wisely. Therefore, the planar and hierarchical partitions are better used for small treewidth networks such that balanced partitions and fewer boundaries can be achieved together; while core-periphery is suitable for large treewidth networks as it deliberately limits the boundary vertex number by bandwidth.
5. Universal PSP Index Scheme and Newly Generated PSP Indexes
In this section, we propose a universal PSP index scheme and generator and put forward five representative PSP indexes for query-oriented or update-oriented applications based on this scheme.
Figure 9. PSP Index Generator with PSP Index Scheme and Representative Generated PSP Indexes
5.1. PSP Index Scheme and Generator
After elaborating on each component of the PSP index, we now have our PSP index scheme ready as shown in Figure 9. The main part is the PSP index generator, which is formed by coupling those three components. For easy selection of the PSP strategy and SP algorithm, we sort all of their choices roughly in terms of query and update efficiency. As shown in Figure 9, the lighter color signifies faster efficiency for PSP strategy and SP algorithm. Meanwhile, planar and hierarchical partitions are more suitable for low-treewidth networks, while core-periphery can deal with high-treewidth graphs. Given the specific application requirements in terms of computation efficiency, space consumption, or partition type, our scheme could provide one PSP index with a specified PSP strategy, SP algorithm, and partition structure through the PSP index generator. For starters, the existing PSP indexes can find their positions in our scheme as will be discussed in Section 7, and we can compare them theoretically and fairly. Apart from this, we can further utilize it as a PSP Index Generator to create new PSP indexes that are suitable to user requirements, especially when all the existing ones are query-oriented, while there leaves a vast space of combinations for new ones. Specifically, the generated PSP index needs to choose the components from each of the three dimensions, and its name also has three parts corresponding to the three dimensions. Although we can enumerate them all, we choose to introduce the representative ones from the perspectives of query- and update-oriented under different partition structures, as shown in Figure 9.
5.2. Representative Generated New PSP Indexes
1. Query-Oriented Planar PSP Index: P-TD-P. Given a planar partition, the PSP strategies can apply directly to it. Next, we introduce how to construct the PSP indexes that are efficient in query processing under planar partition. Firstly in terms of PSP strategy, since both Pre-Boundary and Post-Boundary are efficient in query answering and our proposed Post-Boundary is faster than Pre-Boundary in index construction, we use it as the PSP strategy. Secondly, in terms of the SP algorithm, we could choose TD as the index for both partitions and overlay graph. Although all-pair is faster, its space consumption is intolerable. Its structure is shown in Figure 9-(a). Specifically, the index construction takes $O(V_{max}\cdot(\log V_{max}+h_{max}\cdot w_{max}))$ time for the partition TD, $O(V_{\tilde{G}}(\log V_{\tilde{G}}+h_{\tilde{G}}\cdot w_{\tilde{G}}))$ for the overlay TD, $O(B^{2}_{max}w_{\tilde{G}})$ for boundary correction, and $O(w_{\tilde{G}}+w^{2}_{max}\cdot\delta)$ for the partition TD refresh, where $\delta$ is the affected shortcut number and $max$ are the corresponding max values in the partitions. For the query processing, the intra-query takes $O(w_{max})$ as partition index is correct, and the inter-query takes $O(max\{B_{max}w_{max},B^{2}{max}w_{\tilde{G}}\})$ for the partition and overlay query and $O(B^{2}_{max})$ for the combinations. For the index update, it takes $O(w^{2}_{max}\delta)$ to update partition TD, $O(w^{2}_{\tilde{G}}\delta)$ to update overlay TD, and $O(B^{2}_{max}w_{\tilde{G}})$ to check boundaries.
2. Update-Oriented Planar PSP Index: N-CH-P. In terms of the PSP strategy, we use No-Boundary as it requires the least effort to update. In terms of SP algorithm, we can choose CH as the underlying index as it is fast to update while the query processing is better than direct search (Figure 9-(b)). Its index construction is faster with $O(V_{max}\cdot w^{2}_{max}\cdot\log V_{max})$ for partition CH and $O(V_{\tilde{G}}\cdot w^{2}_{\tilde{G}}\cdot\log V_{\tilde{G}})$ for overlay CH; the query time is longer with $O(max\{B_{max}\cdot w_{max}\log V_{max},$ $B^{2}_{max}\cdot w_{\tilde{G}}\log V_{\tilde{G}}\})$ for the intra- and overlay searching, and $O(B^{2}_{max})$ for the combinations; the index update is faster with $O(\delta w_{max})$ for partition CH maintenance and $O(\delta w_{\tilde{G}})$ for overlay CH maintenance.
3. Query-Oriented Core-Tree PSP Index: P-PT-CP. The core-periphery partition index comprises the core index $L_{c}$ and the periphery index $\{L_{i}\}$. It seems that $L_{c}$ and $L_{i}$ can be constructed by their corresponding subgraphs $G_{c}$ and $G_{i}$, respectively. Although the core does not belong to any partition, they are connected to the partitions, and we treat the core as the overlay graph. Different from the previous planar partition, the core here usually has a large degree so its index is limited to PLL. As for sketch, we omit it here because its PLL core + pruned direct search seems to be the only solution for huge networks. Next, we discuss the remaining parts. For the PSP strategy, Post-Boundary is utilized as the queries within the periphery can be handled without the core index. The periphery index uses TD because the periphery usually has a small degree. This structure is shown in Figure 9-(c). In terms of index construction, because the periphery is constructed through contraction, we regard it as a by-product of the partition phase and do not construct their labels in the first step. Then it takes $O(w_{c}E_{c}\log V_{c}+w^{2}_{c}V_{c}\log^{3}V_{c})$ for the core PLL, $O(B^{2}_{max}w_{c}\log V_{c})$ for boundary correction, and $O(V_{max}\cdot w^{2}_{max}\cdot\log V_{max})$ for periphery label. In terms of query processing, the intra-query takes $O(w_{max})$ time. The inter-query takes $O(max\{B_{max}w_{max},B^{2}_{max}w_{c}\log V_{c}\})$ for the periphery and core, and $O(B^{2}_{max})$ for the combinations. In terms of index update, it takes $O(w^{2}_{max}\cdot\delta)$ for the periphery TD, $O(w_{c}E_{c}\log V_{c})$ for the core PLL, and $O(B^{2}_{max}w_{c}\log V_{c})$ for boundary correction.
4. Update-Oriented Core-Tree PSP Index: N-PC-CP. As shown in Figure 9-(d), N-PC-CP uses CH for faster periphery update while PLL is for the core. It adopts No-Boundary strategy for faster update. In index construction, only core needs $O(w_{c}E_{c}\log V_{c}+w^{2}_{c}V_{c}\log^{3}V_{c})$ time. The query time is longer with $O(max\{B_{max}\allowbreak w_{max}\log V_{max},$ $B^{2}_{max}w_{c}\log V_{c}\})$ for intra and core, and $O(B^{2}_{max})$ for combinations. The index update is fast with $O(\delta w_{max})$ for periphery shortcuts and $O(w_{c}E_{c}\log V_{c})$ for the core.
5. Update / Query-Oriented Hierarchical PSP Index: N-TS-HP. This category organizes $L$ levels of partitions hierarchically with several lower partitions forming a larger partition on the higher level. We use $L^{l}_{i}$ to denote the index of partition $G^{l}_{i}$ on level $l$. Different from the SOTA G-Tree stream of indexes which uses all-pair in their hierarchical overlay graph, we replace it with the hierarchical labels for better query and update performance. Specifically, for vertices in each layer, we store their distance to vertices in their upper layers. As this is essentially 2-hop labeling, we use TD to implement it with orderings corresponding to the boundary vertex hierarchy. Such a replacement in the overlay index could answer queries and update much faster than the original dynamic programming-based layer all-pair index. As for the partitions, this structure tends to generate small partitions so we inherit the original search for fast query processing. Consequently, no partition index leads to inevitable boundary all-pair searches. Fortunately, our No-Boundary restricts the search space to the small partition compared with G-Tree’s whole graph Pre-Boundary (Figure 9-(e) and (f)). To construct the index, the partition boundary all-pair takes $O(B_{max}\cdot V_{max}(logV_{max}+E_{max}))$, and its by-product boundary-to-partition can be cached for faster inter-query. Then the overlay TD takes $O(V_{\tilde{G}}(logV_{\tilde{G}}+h_{\tilde{G}}w_{\tilde{G}}))$ time. In query processing, the intra-query takes $O(B_{max}V_{max}(logV_{max}+E_{max})$ for the direct search (very rare as the partitions are small), while the inter-query The query takes $O(B_{max}w_{\tilde{G}}\})$ for intra- (constant time with cache) and hierarchical query, and $O(B^{2}_{max})$ for combination. As for index update, it takes $O(B_{max}\cdot V_{max}(logV_{max}+E_{max}))$ to update the partition all-pairs and $O(w^{2}_{\tilde{G}}\cdot\delta)$ to update the overlay graph.
Remark 3. We note that these five PSP indexes are novel structures, each with unique technical challenges, especially in updating. Nevertheless, we believe the PSP scheme holds greater significance than these new indexes, as it can guide new PSP index designs. We omit their details due to limited space, but their source codes are provided (GitHub, [n.d.]). Besides, we take the N-CH-P index as an example and elaborate on its details in the Appendix (FullVersion).
Table 2. Dataset Description
[] Type Name Dataset $|V|$ $|E|$ Road Network NY 1 New York City 264,346 730,100 FL 1 Florida 1,070,376 2,687,902 W 1 Western USA 6,262,104 15,119,284 US 1 Full USA 23,947,347 57,708,624 Complex Network GO 2 Google 855,802 8,582,704 SK 3 Skitter 1,689,805 21,987,076 WI 3 Wiki-pedia 3,333,272 200,923,676 FR 2 com-Friendster 65,608,366 3,612,134,270 [1] http://www.dis.uniroma1.it/challenge9/download.shtml [2] http://snap.stanford.edu/data [3] http://konect.cc/
Table 3. Average Vertex Degree of Overlay Graph
Dataset NY FL GO SK
Method PUNCH HEP PUNCH HEP KaHyPar HEP KaHyPar HEP
w/o Opt 48.68 325.98 36.17 719.25 947.36 9754.52 1420.18 34889.5
w/ Opt 48.68 320.61 36.17 714.30 912.24 7678.05 1409.98 27326.5
6. Experimental Evaluation
In this section, we evaluate the effectiveness of the proposed methods and provide a systematic evaluation of PSP indexes. All the algorithms are implemented in C++ with full optimization on a server with 4 Xeon Gold 6248 2.6GHz CPUs (total 80 cores / 160 threads) and 1.5TB memory. The thread number is set to 150.
6.1. Experimental Settings
Datasets and Queries. We test on eight real-life datasets, including four weighted road networks (NY, FL, W, and US) and four complex networks (Web graph: GO and WI; Social network: SK and FR), as shown in Table 2. The edge weight of complex networks is generated by following the manner (Zhang et al., 2021b) that is inversely proportional to the highest degree of the endpoints. We randomly generate 10,000 queries and 1,000 update instances for each dataset to assess the query processing and index maintenance efficiency, respectively.
Partition Methods and Their Performance Metrics. To select the suitable partition method for PSP index, we implement 8 representative partition methods: Bubble (Diekmann et al., 2000), RCB (Simon, 1991), KaHyPar (Akhremtsev et al., 2017), PUNCH (Delling et al., 2011), SCOTCH (Barat et al., 2018), METIS (Karypis and Kumar, 1998), HEP (Mayer and Jacobsen, 2021), CLUGP (Kong et al., 2022), all of which are planar partition methods with PUNCH and RCB only working on road networks. Since core-periphery partition (including Tree-Decomposition-based partition (Ouyang et al., 2018) and landmark selection in Sketch (Farhan et al., 2022)) and hierarchical partition (mostly using METIS (Karypis and Kumar, 1998)) methods are finite, they are selected by default in the corresponding PSP index. We evaluate the performance of a partition method from three aspects: overall border number $|B|=\sum_{i\in[1,k]}|B_{i}|$, average partition border number $\overline{|B_{i}|}=\frac{\sum_{i\in[1,k]}|B_{i}|}{k}$, and partition connectivity $R_{C}=\frac{\sum_{i\in[1,k]}|C(G_{i})|}{k}$, where $C(G_{i})$ is the connected component number of subgraph $G_{i}$.
PSP Indexes and Their Performance Metrics In terms of PSP indexes comparison, we evaluate 5 existing representative PSP indexes spreading over our proposed three partition structures: a) FHL (Liu et al., 2021, 2022a) and THop (Li et al., 2019c) belongs to planar partition; b) Core-Tree (CT) (Li et al., 2020a; Zheng et al., 2022) and Sketch (Farhan et al., 2022) belong to the core-periphery partition; c) G-Tree (Zhong et al., 2013) belongs to the hierarchical partition. Besides, we implement our 5 newly proposed PSP indexes (see Section 5): P-TD-P, N-CH-P, P-PT-CP, N-PC-CP, and N-TS-HP. For the state-of-the-art shortest path algorithms CH (Geisberger et al., 2008; Ouyang et al., 2020), TD (Ouyang et al., 2018; Zhang et al., 2021c), and PLL (Akiba et al., 2013; Zhang et al., 2021b), we implement their partitioned version (named PCH, PTD, and PPLL respectively) to test the performance of different PSP strategies and select the suitable partition number. We measure the performance of the PSP index from four aspects: index construction time $t_{c}$, query time $t_{q}$, update efficiency $t_{u}$, and index size $s$. Note that we do not report the result (or set it to $\infty$) if the running time of a partition method or PSP index exceeds 24 hours.
Parameter Setting. We set the default partition number as $32$ for all planar partitions and the bandwidth of core-tree indexes as $40$ as per experimental results. The landmark number of Sketch is set to $20$ by referring to (Farhan et al., 2022). As for the hierarchical partition, we follow G-tree (Zhong et al., 2013; Li et al., 2019a) and set the fan-out $f$ as $4$ and the maximum leaf node size $\tau$ as 128 in NY, 256 in FL, 512 in W and US, respectively.
Figure 10. Partition Performance When Varying $k$.
Figure 11. Effect of Partition Methods. Bar: Left, Ball: Right
Figure 12. Effect of Partition Number. Bar: Left, Ball: Right
Figure 13. Comparison with Non-partitioned SP Indexes. Bar: Left, Ball: Right
Figure 14. Effectiveness of Partitioned Shortest Path Strategy. (Bar: Left Axis; Ball: Right Axis)
6.2. Experiment Result Analysis
Exp 1: Performance of Partition Methods When Varying $k$. We first report the performance of different partition methods with varying partition number $k$ (we only report results on FL, US, and GO due to limited space; the full results are in the full version (Zhang et al., 2023b))). As shown in Figure 10 (a)-(f), the overall border number $|B|$ generally increases, while the average border number $\overline{|B_{i}|}$ decreases as $k$ grows from 4 to 256. This indicates that having a larger partition number can potentially speed up the query processing of the PSP index due to a smaller average partition border number. However, it also leads to a larger overlay graph, which can slow down the query processing and construction of the overlay index. Therefore, there should be a trade-off between them, and the best partition number typically varies for different PSP indexes. Furthermore, as per Figure 10 (g)-(i), only PUNCH and Bubble have a partition connectivity ratio $R_{C}$ of 1, which implies that at least one partition generated by other partition methods is unconnected (with multiple separate components). Since PTD requires the vertices within one partition to be connected, only PUNCH and Bubble suit it.
Exp 2: Effect of Partition Method. Next, we further evaluate the effect of the partition method on the PSP indexes (PCH, PTD, and PPLL). In particular, we test PCH and PPLL with Bubble, PUNCH, KaHyPar, and METIS on road networks (NY and FL) while only testing PTD with PUNCH and Bubble by referring to Exp 1. Figure 11 shows that PUNCH and KaHyPar perform better than other partition methods in most cases due to smaller overall border number $|B|$ and average partition border number $\overline{B_{i}}$, which indicates that $|B|$ and $\overline{B_{i}}$ are important factors in the partition method selection. We select PUNCH as the default planar partition method due to its superior PSP index performance. It is worth noting that we also test PPLL with KaHyPar, METIS, and SCOTCH on complex networks. However, it fails to complete the index construction on the smallest graph GO within 48 hours due to the large and condensed overlay graph, which reveals that the PSP index with planar partition is generally unsuitable for complex networks.
Exp 3: Effect of Partition Number. As shown in Figure 12 (a)-(b), the indexing time of PCH and PTD increased as $k$ increased from 4 to 256. This is because a larger $k$ could lead to a larger overlay graph that cannot leverage thread parallelization, thus resulting in higher indexing time. By contrast, PPLL’s indexing time first decreases and then increases in NY while tending to decrease in FL because a larger $k$ could also lead to a smaller $\overline{|B_{i}|}$, thus reducing the overlay graph construction time. In Figure 12 (c)-(d), we observe that a smaller $k$ generally improves the query time but leads to worse index update efficiency. Therefore, it is crucial to balance the query and update efficiency by selecting the appropriate partition number. We set the default value of $k$ as 32 since it provided satisfactory performance in most cases.
Exp 4: Comparison with Non-partitioned SP Algorithms. We conduct a comparison between PSP indexes (including PCH, PTD, PPLL, and CT) and their non-partitioned counterparts (namely CH (Geisberger et al., 2008), TD (Ouyang et al., 2018), PLL (Akiba et al., 2013)). As depicted in Figure 13, PSP indexes generally provide faster indexing time, smaller index size, and better scalability than their non-partitioned counterparts. For instance, PCH and PTD exhibit up to 2.76×2.76$\times$2.76 × and 2.25×2.25$\times$2.25 × faster index construction compared to CH and TD, respectively. Moreover, CT demonstrates 72.3×72.3$\times$72.3 × and 3.58×3.58$\times$3.58 × improvement over PLL in terms of indexing time and index size, respectively. However, PSP indexes may compromise query processing efficiency and index update time. It is worth noting that it also shows that planar structure is more suited for road networks, while core-periphery structure is better suited for complex networks, as PPLL does not outperform PLL in terms of indexing time and scalability.
Exp 5: Effectiveness of PSP Strategy. As shown in Figure 14 (a)-(d), the no-boundary strategy has the minimum index construction (about $14-35\%$ speed-up) and index update time (up to two orders of magnitude speed-up compared with the pre-boundary strategy for PTD and PCH), which indicates the superiority of our no-boundary strategy. Suffering from the highest index construction time, the post-boundary strategy has the same query time as the pre-boundary strategy but a significant update efficiency improvement (14.636.7×14.6-36.7$\times$14.6 - 36.7 × speed up for PTD). To further demonstrate the effectiveness of our post-boundary strategy, we test the same-partition query efficiency as presented in Figure 14 (e)-(f). The post-boundary archives about 3.76.3×3.7-6.3$\times$3.7 - 6.3 × and 4.830.5×4.8-30.5$\times$4.8 - 30.5 × speed-up compared with the no-boundary strategy for PTD and PPLL, respectively. Such a significant performance gain comes from the post-boundary’s correct local partition index. Finally, we evaluate the effectiveness of pruning-based overlay graph optimization by measuring the average vertex degree of the overlay graph, as shown in Table 3. We test PUNCH and HEP on road networks while the KaHyPar and HEP on complex networks. The degree decreases with the pruning-based optimization in most cases. It cannot improve PUNCH since PUNCH already has a great partition result.
Figure 15. Comparison of All Shortest Path Indexes. Bar: Left, Ball: Right
Exp 6: Comparison of All PSP Algorithms. We compare our proposed PSP indexes with the existing PSP indexes in Figure 15. We first analyze the performance of our proposed indexes against their counterparts: 1) P-TD-P has the same query performance and index size as FHL, but it is much faster to update with $1$ 202.7 × speed-up. 2) N-CH-P has the fastest index construction and maintenance efficiency and the smallest index size among all planar PSP indexes, achieving up to 3.4×3.4$\times$3.4 ×, 1004×1004$\times$1004 ×, and 7.6×7.6$\times$7.6 × speed-up over FHL. 3) N-PC-CP has faster index construction and update than CT since it utilizes faster CH for tree index. P-PT-CP outperforms CT in terms of query processing as it is faster in same-partition queries. 4) N-TS-HP has faster index construction, smaller index size, faster query and update than G-Tree by leveraging TD for overlay index construction. 5) Sketch can scale up to very large graphs (the only algorithm that can be applied on graph FR), but its query efficiency is much slower than N-PC-CP and P-PT-CP due to the light index.
We next discuss the advantages of different partition structures. Figure 15-(b) shows that the state-of-the-art planar PSP indexes (N-CH-P, P-TD-P) and hierarchical PSP indexes (N-TS-HP) generally provide better query and index update efficiency for road networks than core-periphery PSP indexes. However, only core-periphery PSP indexes can be used on complex networks, which typically have large treewidth. Therefore, planar and hierarchical partitions are more suitable for road networks, while complex networks work better with core-periphery partitions.
6.3. PSP Application Guidance
Based on the experimental analysis in Section 6.2, we clarify the misconceptions mentioned in Section 1 and conclude the following observations: 1) all PSP indexes are essentially interrelated because they all consist of three dimensions in our PSP index scheme: graph partition method, PSP strategy, and shortest path algorithm. Moreover, we could combine the elements of these three dimensions to create a new PSP index; 2) it is important to choose the right partition structure for different types of graphs. Generally, for road networks, planar and hierarchical PSP indexes are more suitable than core-periphery PSP due to their efficient query and index update efficiency. On the other hand, core-periphery PSP indexes are more suitable for complex networks due to their excellent scalability; 3) although PSP indexes typically have faster indexing efficiency, smaller index size, and better scalability than their non-partitioned counterparts, they sacrifice query and index update efficiency.
According to our theoretical analysis and experimental results, we further provide our decision tree for PSP index design as illustrated in Figure 16. The first step in creating a new PSP index is to select a partition structure that is suitable for the specific application, such as graph type and size. For graphs with low treewidth, such as road networks, we recommend using planar or hierarchical partition structures. For larger graphs, we suggest using the core-periphery structure, with the core-tree structure for million-level graphs and sketch for billion-level graphs. The second step is to select the PSP strategy based on the required query processing and index update efficiency. For query-oriented applications, we recommend using the post-boundary strategy, while for update-oriented applications, the no-boundary strategy is more suitable. The third step is to choose the SP algorithm for the overlay and partitions. For a small number of vertices, we recommend using the most efficient all-pair table. For larger graphs, if tree decomposition is feasible, we suggest using TD or CH to achieve better query/update efficiency. If tree decomposition is infeasible, we recommend selecting PLL or even the direct search for SP algorithm selection.
Figure 16. Our Decision Tree of PSP Index Design
7. Related Work
Recently, multiple surveys about the SP computation have emerged, focusing on either specific graph types (like transportation networks (Bast et al., 2016) or complex networks (Zhang et al., 2022)) or static graphs (Madkour et al., 2017; Li et al., 2017; Wu et al., 2012). (Zhang et al., 2021a) investigates the computation in dynamic graphs, but it does not consider the partitioned counterpart. The partitioned pathfinding framework is proposed in (Chondrogiannis and Gamper, 2014, 2016), however, the dynamic version is not discussed and they can be covered by the traditional PSP strategy. In this paper, we study the dynamic PSP problem systematically and explore its performance in both road networks and complex networks by involving the state-of-the-art discovery of SP algorithms and partition methods. Next, we discuss and categorize the existing PSP methods based on our scheme:
1) Pre-Boundary + Search/All-Pair + Hierarchy/Planar: These methods (HiTi (Jung and Pramanik, 2002), Graph Separators, Customizable Route Planning (Delling et al., 2017) (Holzer et al., 2009), ParDiSP (Chondrogiannis and Gamper, 2014, 2016)) pre-compute the shortest distance between boundaries (some hierarchically) to guide the search. In a broad sense, Arc-flag (Möhring et al., 2007) and SHARC (Bauer and Delling, 2010) also belong to this category. G-tree (Zhong et al., 2013; Li et al., 2019a) uses dynamic programming to compute the distance between layer’s all-pair information to replace searching, and it is widely used in kNN (Zhong et al., 2015), ride-sharing (Ta et al., 2017), time-dependent routing (Wang et al., 2019), and machine learning-based path finding (Huang et al., 2021). They are slow to construct due to Pre-B and slower to query to direct search;
2) Pre-Boundary+Search/PLL+Planar: COLA (Wang et al., 2016) builds the labels for the skyline shortest path on the overlay graph to answer the constrained shortest path query. This structure’s construction suffers from Pre-B and query suffers from searching;
3) Pre-Boundary+TD/TD+Planar: FHL (Liu et al., 2021, 2022b, 2022a) partition builds the TD both within and between partitions for multi-dimensional skyline paths. As validated in Exp 4, our Post-B can speedup Pre-B’s construction dramatically;
4) Pre-Boundary+PLL/ PLL+Planar: T2Hop (Li et al., 2019c, 2020b) utilizes two layers of PLL to reduce the complexity of long range time-dependent paths, and this structure’s performance is limited by PLL;
5) No-Boundary+ PLL/TD+Core: Core-Tree (Li et al., 2020a; Zheng et al., 2022) is the baseline that we already discussed;
6) Pre-Boundary+Search/All-Pair/PLL+Sketch: This category works on huge graphs where index is nearly impossible so only a small number of landmarks are selected to either help prune the search (Das Sarma et al., 2010; Wang et al., 2021; Farhan et al., 2022) or approximate result (Gubichev et al., 2010; Qiao et al., 2012).
In sum, all the existing PSP methods have their limitation, warranting the comprehensiveinvestigation embarked on in this paper.
8. Conclusions
In this work, we systematically study the partitioned shortest path index and propose a universal PSP scheme by concomitantly considering three dimensions: PSP strategy, SP algorithm, and partition structure. For the PSP strategy, we propose two new strategies and pruned-boundary optimization for better index performance. We also provide index maintenance solutions for both traditional and novel PSP strategies. For partition structure, we propose a new path-oriented partition method classification. Furthermore, we construct our PSP index scheme by seamlessly coupling those three dimensions. We also propose a PSP index generator and five new PSP indexes that are either more efficient in query or update than the current state-of-the-art. Systematic and extensive experimental evaluations fully validate our findings and prove the effectiveness of our new indexes guided by the insightful PSP index scheme. Finally, PSP index design guidelines are provided to help future researchers and practitioners design their indexes. Our future work will address recommendation aspects for such designs.

### v3

@@
-A Universal Scheme for Dynamic Partitioned Shortest Path Index
+A Universal Scheme for Dynamic Partitioned Shortest Path Index: Survey, Improvement, and Experiments
 Abstract.
-Shortest path (SP) computation is the fundamental operation in various networks such as urban networks, logistic networks, communication networks, social networks, etc.
-With the development of technology and societal expansions, those networks tend to be massive.
-This, in turn, causes deteriorated performance of SP computation, and graph partitioning is commonly leveraged to scale up the SP algorithms.
-However, the partitioned shortest path (PSP) index has never been systematically investigated and theoretically analyzed, and there is a lack of experimental comparison among different PSP indexes.
-Moreover, few studies have explored PSP index maintenance in dynamic networks.
-Therefore, in this paper, we systematically analyze the dynamic PSP index by proposing a universal scheme for it.
-Specifically, we first propose two novel partitioned shortest path strategies (No-boundary and Post-boundary strategies) to improve the performance of PSP indexes and design the corresponding index maintenance approaches to deal with dynamic scenarios.
-Then we categorize the partition methods from the perspective of partition structure to facilitate the selection of partition methods in the PSP index.
-Furthermore, we propose a universal scheme for designing the PSP index by coupling its three dimensions (i.e.
-PSP strategy, partition structure, and SP algorithm).
-Based on this scheme, we propose five new PSP indexes with prominent performance in either query or update efficiency.
-Lastly, extensive experiments are implemented to demonstrate the effectiveness of the proposed PSP scheme, with valuable guidance provided on the PSP index design.
-PVLDB Reference Format:
-PVLDB, (): , .
-doi: This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
-Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
-For any use beyond those covered by this license, obtain permission by emailing info@vldb.org.
-Copyright is held by the owner/author(s).
-Publication rights licensed to the VLDB Endowment.
-Proceedings of the VLDB Endowment, Vol. , No.  ISSN 2150-8097.
-doi:
-footnotetext: *Both authors contributed equally to this paper. Lei Li is the corresponding author.
+Shortest Path (SP) computation is a fundamental operation in many real-life applications such as navigation on road networks, link analysis on social networks, etc.
+These networks tend to be massive, and graph partitioning is commonly leveraged to scale up the SP algorithms.
+However, the Partitioned Shortest Path (PSP) index has never been systematically investigated.
+Moreover, few studies have explored its index maintenance in dynamic networks.
+In this paper, we survey the dynamic PSP index and propose a universal scheme for its design and analysis.
+Specifically, we first review the SP algorithms and put forward a novel structure-based partition method classification to facilitate the selection of partition methods.
+Furthermore, we summarize the existing Pre-boundary PSP strategy and propose two novel strategies (No-boundary and Post-boundary) to improve its index performance.
+Lastly, we propose a universal scheme with three dimensions (SP algorithm, partition method, and PSP strategy) to facilitate the analysis and design of the PSP index.
+Based on this scheme, we put forward five new PSP indexes with a prominent query or update efficiency performance.
+Extensive experiments are conducted to evaluate the performance of the PSP index and the effectiveness of the proposed techniques, with valuable guidance on the PSP index design.
+Shortest Path Index, Dynamic Graph, Graph Partition
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
 1. Introduction
-Shortest Path (SP) query in dynamic networks is an essential building block for various applications affecting daily lives.
-Given an origin-destination pair, it returns the minimum traveling time in a road network (Bast et al., 2016), the fastest connection in the web graph (Boccaletti et al., 2006), or the most intimate relationships in the social network (Kolaczyk et al., 2009; Opsahl et al., 2010; Yao et al., 2013).
-With the development of urban traffic systems and the evolvement of online interactions platforms, real-life networks tend to be massive, which brings great challenges to the scalability of the state-of-the-art (Li et al., 2020a, 2019b; Zhang et al., 2021a, c, b, 2023a) with either heavy memory or expensive search overheads.
-For example, 2-Hop Labeling (Cohen et al., 2003) has the fastest query efficiency but requires $O(nm^{1/2})$ space (for a graph with $n$ vertices and $m$ edges) to store the index.
-Complementary, the search-based methods like Dijkstra’s are inefficient on query processing on large networks.
-Moreover, because the networks are dynamic in nature with evolving structures or edge weights, extra efforts are required to handle the updates for SP indexes.
-Graph partitioning is often used to improve the scalability of SP algorithms, enabling more complicated problems like time-dependent (Li et al., 2019c, 2020b) and constraint path computation (Wang et al., 2016; Liu et al., 2021, 2022a) to be efficiently handled in large networks.
-It decomposes a large network into several smaller ones such that the construction time can be improved by parallelization, and the index sizes and update workloads can be reduced.
-As shown in Figure 1, we can compare them briefly in terms of index construction time, storage space, query time, and update time.
+Shortest Path (SP) query in networks is an essential building block for various daily life applications.
+Given an origin-destination pair such as $(s,t)$, the point-to-point shortest path query aims to find the path with the minimum length from vertex $s$ to vertex $t$ on network $G$.
+In real-life applications, this shortest path could be the path with minimum traveling time in a road network (Ding et al., 2008; Li et al., 2017, 2018, 2019c, 2020b; Bast et al., 2016; Zhou et al., 2025), the fastest connection in the web graph or Internet (Boccaletti et al., 2006; Scheideler, 2002), or the most intimate relationships in the social network (Gong et al., 2016; Kolaczyk et al., 2009; Opsahl et al., 2010).
+With the development of urban traffic systems and the evolvement of online interaction platforms, real-life networks tend to be massive, which brings great challenges to the scalability of the existing shortest-path algorithms with either heavy memory or expensive search overheads.
+For example, index-free algorithms such as Dijkstra’s (Dijkstra, 1959) and A* (Hart et al., 1968) directly search on the graphs in a best-first manner for query processing, thus are time-consuming on large networks due to large search space.
+To this end, index-based algorithms such as Contraction Hierarchy (CH) (Geisberger et al., 2008) and 2-Hop Labeling (HL) (Cohen et al., 2003) are proposed to expedite the query efficiency.
+In particular, 2-hop labeling (Cohen et al., 2003) such as Pruned Landmark Labeling (PLL) (Akiba et al., 2013) and Hierarchical 2-Hop labeling (H2H) (Ouyang et al., 2018) has the state-of-the-art query efficiency but requires $O(nm^{1/2})$ space to store the index for a graph with $n$ vertices and $m$ edges.
+Moreover, since the networks are dynamic in nature with evolving edge weights or structures, extra efforts such as space and maintenance overheads are required to handle the updates for shortest path indexes, further deteriorating their scalability.
+Graph partitioning is often used to improve the scalability of shortest path algorithms and enable more complicated path problems like time-dependent (Li et al., 2019c, 2020b), constraint (Wang et al., 2016; Liu et al., 2021, 2022), and top-$k$ path enumeration (Yu et al., 2020, 2024).
+We refer to the shortest path algorithms (indexes) adopting graph partitioning as the Partitioned Shortest Path (PSP) algorithms (indexes) [^1] The term “index” also refers to “index-based algorithm”.
+We use the terms “algorithm” and “index” interchangeably when the context is clear..
+They generally decompose a large network into several smaller ones such that the query processing could be enhanced for index-free algorithms, or the construction time and index size can be reduced for index-based algorithms.
+As shown in Figure 1, we classify the existing SP algorithms into six categories and compare them briefly in terms of index construction time, storage space, query time, and update time.
 On one extreme are the direct search algorithms (Dijkstra, 1959; Hart et al., 1968; Li et al., 2020c; Zhang et al., 2020) that require no index and can work in dynamic environments directly but are slow for query answering.
 Then the partitioned search adds various information to guide and reduce search space (Chondrogiannis and Gamper, 2016; Möhring et al., 2007; Bauer and Delling, 2010) to improve query processing.
-On the other extreme, Contraction Hierarchy (CH) (Geisberger et al., 2008) and 2-hop Labeling (HL) (Cohen et al., 2003; Akiba et al., 2013; Li et al., 2019b; Ouyang et al., 2018) have larger index sizes, longer construction and maintenance time but faster query performance.
-Their partitioned versions (Delling et al., 2017; Li et al., 2020a; Li et al., 2019c; Wang et al., 2021; Farhan et al., 2022; Liu et al., 2021, 2022b, 2022a) are faster to construct with smaller sizes but longer query times.
-As these works have better performance than their unpartitioned versions, they create the following misconceptions regarding the benefits of partitions on pathfinding-related problems: 1) Graph partitioning and shortest path algorithms are irrelevant since they are two different research branches; 2) Different PSP indexes are irrelevant and very different from each other; 3) Applying graph partitioning is always better.
+On the other extreme, index-based solutions such as CH (Geisberger et al., 2008) and HL (Cohen et al., 2003; Akiba et al., 2013; Li et al., 2019b; Ouyang et al., 2018) have larger index sizes, longer construction and maintenance time but faster query performance.
+Their partitioned versions (Li et al., 2020a, 2019c; Wang et al., 2021; Farhan et al., 2022; Liu et al., 2021, 2023, 2022) are faster to construct with smaller sizes but longer query times.
+As these works have better performance than their unpartitioned versions, they may create the following misconceptions regarding the benefits of partitions on pathfinding-related problems: 1) Graph partitioning and shortest path algorithms are irrelevant since they are two different research branches.
+This is often phrased as “graph partitioning is a well-studied problem and their methods could use any of the existing partition solutions” such that only one partition method is used without justification experimentally (Wang et al., 2016; Li et al., 2019c; Liu et al., 2021; Abraham et al., 2011; Chondrogiannis and Gamper, 2014; Yu et al., 2020); 2) Applying graph partitioning is always better, which is the reason why these PSP indexes were proposed in the first place; 3) Different PSP indexes are irrelevant and very different from each other, as evidenced by the fact that they only compared with their non-partitioned counterparts (Chondrogiannis and Gamper, 2016; Li et al., 2020a; Zheng et al., 2022; Liu et al., 2022; Dan et al., 2022) but not with each other.
+However, these misconceptions are not always true.
+Motivation.
+Although the PSP indexes have been widely applied in the past few decades, they have not been studied systematically.
+Specifically, the existing solutions typically glue one partition method with one SP algorithm without discriminating their characteristics, figure out a way to query correctly, and then claim superiority.
+For instance, THop (Li et al., 2019c) utilizes PUNCH (Delling et al., 2011b) for partitions and PLL (Akiba et al., 2013) for boundary and inner indexes, and processed the queries by concatenating inner and boundary indexes.
+While G-tree (Zhong et al., 2013) uses METIS (Karypis and Kumar, 1998) for hierarchical partitions and builds the distance matrices for vertices on different levels, answering the queries by dynamic programming.
+These two PSPs seem very different, with different partition methods and path indexes.
+However, they essentially share the same strategy for organizing index construction and query processing procedures.
+Besides, most PSP indexes only compare with the non-partitioned counterparts and claim their superiority without comparing with other PSP indexes.
+Consequently, there lacks a generalized scheme to organize, analyze, and compare the PSP indexes insightfully and fairly.
+Moreover, how to extend the PSP indexes to dynamic networks is also unknown.
+To this end, we aim to theoretically and systematically study the dynamic PSP index by conducting a comprehensive survey and proposing a universal scheme to guide its design for different scenarios.
 Figure 1. Performance Comparison of SP Algorithms
-Motivations.
-These misconceptions are not always true.
-Although the partitioned shortest path (PSP) indexes have been widely applied in the past decade, to the best of our knowledge, they have not been studied systematically.
-Specifically, the existing solutions typically pick one partition method without discriminating their characteristics and figure out a way to make it work, and then claim benefits/superiority.
-Consequently, there lacks a generalized scheme to organize and compare the PSP indexes insightfully and fairly.
-Towards that, we postulate that both theoretical and experimental systematic study is needed for PSP indexes, so that a suitable index could be constructed for different scenarios.
-Challenges.
-However, it is non-trivial to design such a scheme.
-Firstly, the current PSP indexes are delicately designed structures with tightly coupled components and all claim certain superiority.
+Challenges. However, achieving this goal presents several significant challenges.
+Firstly, graph partition methods are normally classified from different perspectives (like vertex-cut / edge-cut, in-memory / streaming, and etc..) with various optimization objectives (such as cut size minimization, balancing, and etc..), but these criteria are not path-oriented and thus hardly help to tell which partition method is suitable for certain path applications.
+Therefore, we propose a new partitioning classification from the perspective of partition structure for easier partition method selection.
+We also briefly review the shortest path algorithms used in the PSP index from the perspective of query and update efficiency to guide their selection.
+Secondly, it is unclear how to construct, query, and update a new PSP index as all existing works only claim their unique structures can answer queries correctly and seldom involve index maintenance.
+Therefore, we dive into the PSP strategy, which determines how to build the index, coordinate multiple partitions for correct query answering, and deal with index updates.
+In particular, we summarize classic pre-boundary PSP strategy and then propose two novel PSP strategies (no-boundary and post-boundary strategies) and index update methods to improve its index performance.
+Besides, a pruning-based overlay graph optimization is also proposed to improve the index and update efficiency further.
+Thirdly, the current PSP indexes are delicately designed structures with tightly coupled components and all claim certain superiority.
 However, since no effort has been put into the abstract PSP index itself, it is unclear what determines their performance, when to use particular indexes, and how to design a suitable new PSP index for a new scenario.
-Therefore, we study the PSP index systematically in this work and identify three dimensions throughout the rich literature.
-Secondly, it is unclear how to construct, query, and update a new PSP index as all existing works only claim their unique structures can answer queries correctly and seldom involve index maintenance.
-Therefore, we dive into the partitioned shortest path (PSP) strategy, which determines how to coordinate multiple partitions for correct query answering, and summarizes the traditional PSP strategy.
-Additionally, we propose two novel PSP strategies together with their index update methods to improve the index performance.
-Besides, a pruning-based overlay graph optimization is also proposed to improve the index and update efficiency further.
-Thirdly, the graph partition methods are normally classified from different perspectives (like vertex-cut / edge-cut, in-memory / streaming, etc.), but these criteria are not path-oriented and thus hardly help to tell which partition method is suitable for certain path-oriented applications.
-Therefore, we propose a path-oriented partitioning classification from the perspective of partition structure for easier partition method selection.
-Lastly, after elaborating on each dimension of PSP index, we propose our universal PSP index scheme by entangling these dimensions.
+Therefore, based on the rich literature, we propose an insightful and universal PSP index scheme with three critical dimensions, i.e., shortest path algorithm, partition method, and PSP strategy.
 Based on this scheme, we propose five new PSP indexes to achieve better efficiency in either query or update operations by selecting the specific strategy or method on each dimension as per system requirements.
-Contributions. Our contributions are listed as follows:
+Contributions. Our contributions and survey structure are listed as follows:
   • $\bullet$
-    We propose a universal PSP index scheme by decoupling the PSP index into three dimensions: PSP strategy, partition structure, and SP algorithm, such that the PSP index can be analyzed and designed systematically.
+    [Literature Review] We propose a novel structure-based partitioning classification to facilitate the partition method selection for the PSP index, and provide comprehensive literature reviews of corresponding partitioning methods and PSP indexes.
+Besides, we also briefly review the shortest path algorithms used in PSP index.
+(Sections 3-4)
+  • $\bullet$
+    [Improvements] We summarize the traditional PSP strategy as Pre-boundary strategy and propose two novel PSP strategies called No-Boundary and Post-Boundary strategies to improve its index construction and update efficiency.
+Furthermore, we investigate the index maintenance of these PSP strategies and design a novel pruning-based overlay graph optimization to prune the unnecessary computation.
 (Section 5)
   • $\bullet$
-    We propose two novel PSP strategies called No-Boundary and Post-Boundary strategies to improve the index construction and maintenance efficiency over the traditional PSP strategy.
-We also provide non-trivial correctness analysis for their index construction, query processing, and index maintenance.
-A novel Pruning-based overlay graph optimization is also designed to prune the unnecessary computation.
+    [PSP Index Scheme and New PSP Indexes] We propose a universal PSP index scheme by decoupling the PSP index into three dimensions: shortest path algorithm, partition method, and PSP strategy, such that the PSP index can be analyzed and designed systematically.
+Based on the scheme, we identify and design five new PSP indexes suitable for specific network structures and scenarios to validate the effectiveness of our scheme.
 (Section 6)
   • $\bullet$
-    We propose a new path-oriented partitioning classification from the perspective of partition structure for easier partition method selection in designing the PSP index. (Section 4)
-  • $\bullet$
-    Based on the proposed PSP index scheme, we put forward a PSP index generator to create new PSP structures.
-Among them, we identify and design five new PSP indexes suitable for specific network structures and scenarios.
+    [Experimental Evaluation] We implement comprehensive experimental studies to explore the performance of various PSP indexes under our proposed universal scheme and provide guidelines for designing PSP indexes with different application scenarios.
 (Section 7)
-  • $\bullet$
-    We implement comprehensive experimental studies to explore the performance of PSP indexes under our proposed universal scheme and provide guidelines for designing PSP indexes with different application scenarios.
-(Section 6)
-2. Preliminary and PSP Dimensions
-In this paper, we focus on the dynamic weighted network $G(V,E,W)$ with $V$ denoting the vertex set, $E$ denoting the edge set and $W:E\rightarrow\mathbf{R}^{+}$ assigning a non-negative weight to each edge – i.e.
-$w(u,v)\in W,\forall(u,v)\in E$ – which can increase or decrease in an ad-hoc manner.
+2. Preliminary
+In this paper, we focus on the dynamic weighted network $G(V,E)$ with $V$ denoting the vertex set and $E$ denoting the edge set.
 We denote the number of vertices and edges as $n=|V|$ and $m=|E|$.
-For each $v\in V$, we represent its neighbors as $N(v)=\{u,|(u,v)\in E\}$, u’s commonly called adjacent to $v$.
+Each edge $e\in E$ is assigned a non-negative weight $w=|e|$, which can increase or decrease in ad-hoc.
+Note that we focus on edge weight change updates since the structural update (i.e., edge/vertex deletion/insertion) can be converted to edge weight updates and leverage their update mechanisms for structural updates (Ouyang et al., 2020; Zhang et al., 2021c, b, a), as will be discussed in Section 5.3.
+For each $v\in V$, we represent its neighbors as $N(v)=\{u,|(u,v)\in E\}$.
 We associate each vertex $v\in V$ with order $r(v)$ indicating its importance in $G$.
-In this paper, we use boundary-first ordering (Liu et al., 2022a) to generate the order for the PSP indexes.
+In this paper, we use boundary-first ordering (Liu et al., 2022) to generate the order for the PSP indexes.
 For non-partitioned SP indexes, we leverage degree to decide the order of small-world networks while the minimum degree elimination-based ordering (Berry et al., 2003) for the road networks.
-A path $p=\left<v_{0},v_{1},\dots,v_{k}\right>((v_{i},v_{i+1})\in E,0\leq i<k)$ is a sequence of adjacent vertices with length of $l(p)=\sum_{i=0}^{k-1}w(v_{i},v_{i+1})$.
-Given an origin $s$ and destination $t$ vertex (OD pair), the shortest path query $Q(s,t)$ calculates the path with the minimum length between them, i.e.
-$p_{G}(s,t)$ with the shortest distance $d_{G}(s,t)$.
-In the following, we introduce the PSP index dimensions and summarize each of them.
-2.1. Dimensions of PSP Index
-Based on a comprehensive literature review and analysis of the PSP index, we identify three critical dimensions of the PSP index: 1) Graph Partition Method, which divides the graph into multiple subgraphs and is the key difference from the standalone SP indexes.
-However, the existing partition methods are not designed for path indexes, so how they affect the PSP index is unknown and deserves study; 2) PSP Strategy, which provides a procedure to organize the index construction for multiple subgraphs and coordinate those indexes such that the query correctness is guaranteed.
-Only straightforward solutions exist currently as will be discussed in Section 2.3.
-However, this part is crucial because it determines the complexity of both the query and update; 3) SP Algorithm, whose influence was thoroughly studied in our previous work (Zhang et al., 2021a).
-With these three dimensions, we can form a space to position the existing methods into it.
-Given a massive dynamic network, these dimensions provide a roadmap to construct a PSP index by selecting a specific graph partition method, SP algorithm and PSP strategy such that the application requirement (query/update efficiency) can be satisfied.
-Next, we discuss and review these three dimensions.
-2.2. Graph Partitioning
-Definition 0 (Graph Partitioning).
-A graph $G$ is decomposed into multiple disjoint subgraphs $\{G_{i}\}$ $(i=1,2,\dots,k)$ with $V_{i}\cap V_{j}=\emptyset$ and $\bigcup V_{i}=V$. denoted as $E_{inter}$ and $E_{intra}$.
-We now discuss and categorize the vertex and edge in graph partitioning.
+A path $p=\left<v_{0},v_{1},\dots,v_{k}\right>((v_{i},v_{i+1})\in E,0\leq i<k)$ is a sequence of adjacent vertices with length of $l(p)=\sum_{i=0}^{k-1}w(v_{i},v_{i+1})$, where $w(v_{i},v_{i+1})$ is the edge length from $v_{i}$ to $v_{i+1}$.
+Given an origin $s$ and destination $t$ vertex (OD pair), the shortest path query $Q(s,t)$ calculates the path with the minimum length between them, i.e., $p_{G}(s,t)$ with the shortest distance $d_{G}(s,t)$.
+We omit the notation $G$ when the context is clear.
+Note that we consider $G$ as an undirected graph, and our techniques can be easily extended to directed graphs.
+Graph partitioning generally divides a graph $G$ into multiple disjoint subgraphs $\{G_{i}\}$ $(i=1,2,\dots,k)$ with [^2] In this paper, we generally refer to graph partitioning as edge-cut partitioning for easier illustration.
+$V_{i}\cap V_{j}=\emptyset$ and $\bigcup V_{i}=V$.
 $\forall v\in G_{i}$, we say $v$ is a boundary vertex if there exists a neighbor of $v$ in the another subgraph, that is $\exists u\in N(v),u\in G_{j}(i\neq j)$.
@@
 We represent the boundary vertex set of $G_{i}$ as $B_{i}$ and that of $G$ as $B=\bigcup B_{i}$.
-For $(u,v)\in E$, we say $(u,v)$ is an inter-edge if both its two endpoints $u$ and $v$ are boundary vertices from different subgraphs, i.e., $u\in B_{i},v\in B_{j},i\neq j$.
+The PSP indexes then build an overlay graph $\tilde{G}$ on all boundary vertices to connect different partitions and preserve the global distances for all cross-partition queries (i.e., $s\in G_{i},t\notin G_{j},i\neq j$).
+For $(u,v)\in E$, we say $(u,v)$ is an inter-edge if both its two endpoints $u$ and $v$ are boundary vertices from different subgraphs (i.e., $u\in B_{i},v\in B_{j},i\neq j$).
 Otherwise, it is an intra-edge.
 The corresponding edge sets are denoted as $E_{inter}$ and $E_{intra}$, respectively.
-For example in Figure 2, the example graph is partitioned into four subgraphs $G_{1}$ to $G_{4}$, with the red vertices denoting the boundaries ($B_{1}=\{v_{3},v_{10}\}$) and the white vertices representing the inner vertices.
-Edges outside partitions like $(v_{3},v_{5})$ are inter-edges, and $(v_{3},v_{10})$ is an intra-edge.
-Besides, the boundary vertex plays an important role in the shortest path computation with its cut property: given a shortest path $p(s,t)$ with its endpoints in two different subgraphs $s\in G_{i},t\notin G_{i}$, there is at least one vertex $v$ which is in the boundary vertex set of $G_{i}$ on the path, that is $\exists v\in p(s,t),v\in B_{i}$.
-Therefore, boundary vertices would determine the correctness and complexity of partitioned shortest path computation (discussed in the following sections).
-Various partition methods are involved in the existing PSP methods such as METIS (Karypis and Kumar, 1998) in G-Tree (Zhong et al., 2013) and SHARC (Bauer and Delling, 2010) ; PUNCH (Delling et al., 2011) in COLA (Wang et al., 2016), FHL (Liu et al., 2021, 2022b), FHL-Cube (Liu et al., 2022a) and THop (Li et al., 2019c, 2020b) ; Tree Decomposition-based Partition (Ouyang et al., 2018) in Core-Tree (Li et al., 2020a, 2019b) ; multi-level partitioning (Jung and Pramanik, 1996) in G-tree (Zhong et al., 2013, 2015), SPAH (Jung and Pramanik, 1996), CPR (Delling et al., 2017) and (Holzer et al., 2009; Chondrogiannis and Gamper, 2016, 2014).
-But, their selections are performed in an ad-hoc manner or consider the boundary number, partition balance, and spatial awareness, rather than the SP computation itself.
-2.3. Traditional PSP Strategy
+Example 0.
+Figure 2 presents an example graph $G$ which is partitioned into four subgraphs $G_{1}$ to $G_{4}$, with the red vertices denoting the boundary vertices (e.g., $B_{1}=\{v_{3},v_{10}\}$) and the white vertices representing the inner vertices (e.g., $v_{1}$ and $v_{2}$).
+Edges outside partitions like $e(v_{3},v_{5})$ are inter-edges, and $e(v_{3},v_{10})$ is an intra-edge.
+Figure 2-(c) illustrates an overlay graph $\tilde{G}$ built on boundary vertices to connect all subgraphs while preserving the global distances (e.g., $|e_{\tilde{G}}(v_{8},v_{9})|=2$).
+Figure 2. Example Graph $G$ and Overlay Graph $\tilde{G}$
+It is worth noting that boundary vertex plays an important role in the shortest path computation with its cut property: given a shortest path $p(s,t)$ with its endpoints in two different subgraphs $s\in G_{i},t\notin G_{i}$, there is at least one vertex $v$ which is in the boundary vertex set of $G_{i}$ on the path, that is $\exists v\in p(s,t),v\in B_{i}$.
+Therefore, boundary vertices determines the correctness and complexity of partitioned shortest path computation, as will be discussed in Section 5.
+Various partition methods are involved in the existing PSP methods such as METIS (Karypis and Kumar, 1998) in CRP (Delling et al., 2017) and (Holzer et al., 2009; Chondrogiannis and Gamper, 2016, 2014); PUNCH (Delling et al., 2011b) in COLA (Wang et al., 2016), FHL (Liu et al., 2021, 2023), FHL-Cube (Liu et al., 2022) and THop (Li et al., 2019c, 2020b); core-tree decomposition (Li et al., 2020a) in Core-Tree (Li et al., 2020a; Zheng et al., 2022); multi-level partitioning (Karypis and Kumar, 1998) in G-tree (Zhong et al., 2013, 2015), HiTi (Jung and Pramanik, 1996), and CRP (Delling et al., 2017).
+However, their selections are performed in an ad-hoc manner without comparing each other and discerning the SP computation requirements.
+3. Shortest Path Algorithms in PSP Index
+In this Section, we briefly review the shortest path algorithms used in the PSP indexes and roughly compare their query efficiency and update efficiency to facilitate the selection of shortest path algorithms.
+It is worth noting that there are several surveys regarding the shortest path algorithms, such as SP algorithms in static networks (Sommer, 2014; Zhang et al., 2022) and SP algorithms in dynamic networks (Zhang et al., 2021a).
+The interested reader could refer to (Sommer, 2014; Zhang et al., 2022, 2021a) for more details.
+Direct Search.
+Direct search algorithms such as Dijkstra’s (Dijkstra, 1959), $A^{*}$ (Hart et al., 1968), or ALT (Goldberg and Harrelson, 2005) search the graph in a best-first manner.
+They are generally used when the PSP index cannot be constructed due to the large graph size (QbS (Wang et al., 2021)) or complicated problem (Wang et al., 2016; Zhou et al., 2024a).
+Direct search algorithms take no time in index maintenance on dynamic networks, but their query processing is the slowest due to a large search space.
+The PSP indexes adopting direct search include G-tree (Zhong et al., 2013), CRP (Delling et al., 2017), PbS (Chondrogiannis and Gamper, 2014), BMHPS (Zhou et al., 2024a), and so on.
+Contraction Hierarchy (CH).
+CH (Geisberger et al., 2008) is a widely used lightweight index that contracts the vertices in a pre-defined order and preserves the shortest path information by adding shortcuts among the contracted vertex’s neighbors.
+The search-based CH (Geisberger et al., 2008) has a small index size but takes longer to build and maintain (Zhang et al., 2021a; Geisberger et al., 2012) while concatenation-based CH (Ouyang et al., 2020; Wei et al., 2020) is much faster to build and maintain and is more suitable for the graph with small treewidth (Zhang et al., 2021a; Ouyang et al., 2018; Li et al., 2020a; Zhou et al., 2024b).
+CH’s query performance is generally around 10$\times$ faster than Dijkstra’s but much slower than the 2-hop labeling (Cohen et al., 2003).
+The Dynamic CH (DCH) (Ouyang et al., 2020) extends the CH index to dynamic networks and achieves almost the fastest index update efficiency among index-based solutions.
+Surprisingly, no PSP index has used CH as the underlying index so far.
+Pruned Landmark Labeling (PLL).
+Although many 2-hop labeling methods have been proposed in the past decade, only two are widely used: PLL (Akiba et al., 2013; Li et al., 2019b) and Tree Decomposition (TD) (Ouyang et al., 2018).
+Built by either pruned search (Akiba et al., 2013) or propagation (Li et al., 2019b; Jin et al., 2020; Zhang et al., 2021b, 2023), PLL (Akiba et al., 2013) and its parallel indexing version PSL (Li et al., 2019b) is the only index that can work on graphs with large treewidth, such as large social networks, web graphs, and etc.
+The query efficiency of PLL is much higher than CH but lower than TD.
+The index maintenance of PLL can be divided into search-based (Akiba et al., 2014; D’angelo et al., 2019; Qin et al., 2017) or propagation-based (Zhang et al., 2021b; Zhou et al., 2024b).
+The search-based solutions involve time-consuming Dijktra’s searches during the index update.
+By contrast, propagation-based solutions rely on propagation to rectify the incorrect labels, thus achieving much faster efficiency.
+Nevertheless, the update efficiency of PLL is still slower than both CH and TD.
+The PSP indexes adopting PLL include THop (Li et al., 2019c) that uses it to build the overlay and partition indexes, QbS (Wang et al., 2021; Farhan et al., 2022) that leverages it for building labels among landmarks for billion-scale small-world networks, and Core-tree (Li et al., 2020a; Zheng et al., 2022; Zhou et al., 2024b) that uses it to build the core index.
+Tree Decomposition (TD).
+Combining the 2-hop labeling with hierarchy, the query processing and index update of TD (a.k.a.
+Hierarchical 2-Hop labeling (H2H)(Ouyang et al., 2018; Chen et al., 2021) are much faster than PLL for graphs with smaller treewidth (Zhang et al., 2021a), such as road networks.
+However, due to the vertex contraction mechanism, it cannot scale to large-treewidth networks.
+The PSP indexes adopting TD includes Core-tree (Li et al., 2020a; Zheng et al., 2022; Zhou et al., 2024b) that leverage it to build the tree index, and FHL (Liu et al., 2021, 2022) that uses it to build the forest index, and etc.
+All-pair Table.
+The pre-computation of all-pair distances among different vertices enables the fastest query but incurs time-consuming index construction and large index storage, which is infeasible for large graphs.
+Nevertheless, PSP indexes such as G-Tree (Zhong et al., 2013, 2015) and ROAD (Lee et al., 2009, 2010) store the distance between boundary vertices in each partition to expedite the query processing.
+4. Structure-based Partition Method Classification
+An important step for designing a PSP index is to choose a suitable graph partition method.
+However, it is non-trivial to choose a partition method fulfilling our preferred performance because 1) Numerous partition methods with different characteristics were proposed in the past decades; 2) They are classified under different criteria (Buluç et al., 2016), for example, from the perspectives of partition manners (spectral, flow, graph glowing, contraction and multi-level, etc.), partition objectives (balance and minimal cut), computation manner (in-memory, distributed and streaming), and cut type (edge-cut and vertex-cut), but it is not clear which criterion is beneficial to PSP index; 3) Their relationship with (or the influence on) PSP index is unknown and has never been studied.
+Based on these doubts, in this Section, we propose a novel structure-based classification of partition methods to facilitate the partition method selection.
+In particular, we categorize existing partition methods into three types: Planar Partitioning, Core-Periphery Partitioning, and Hierarchical Partitioning, and provide comprehensive reviews about corresponding partition methods and PSP indexes in Section 4.1-4.3.
+An insightful discussion about these partition structures are also provide in Section 4.4.
+4.1. Planar Partitioning
+Planar Partitioning (PP) treats all partitions equally on one level. We provide its formal definition as follows.
+Definition 0.
+(Planar Partitioning).
+Given a graph $G$, planar partitioning decomposes $G$ into multiple equally-important subgraphs $\{G_{i}|1\leq i\leq k\}(k\in[2,\infty])$ with 1) $\bigcup_{i\in[1,k]}V(G_{i})=V$, $V(G_{i})\cap V(G_{j})=\emptyset,\forall i,j\in[1,k],i\neq j$ , or  [^3] We also consider the vertex-cut scenario of graph partition to generalize the definition of planar partitioning.
+2) $\bigcup_{i\in[1,k]}E(G_{i})=E$, $E(G_{i})\cap E(G_{j})=\emptyset,\forall i,j\in[1,k],i\neq j$.
+4.1.1. Planar Partitioning Algorithms
+Sharing almost the same constraint with edge-cut (or vertex-cut) partitioning, the planar partitioning highlights that all partitions are treated equally.
+Representative methods include spectral partitioning (Pothen et al., 1990; Barnard and Simon, 1994), growing-based partitioning (Diekmann et al., 2000), flow-based partitioning (Sanders and Schulz, 2011; Delling et al., 2011b), geometric partitioning (Berger and Bokhari, 1987; Simon, 1991), node-swapping (Kernighan and Lin, 1970; Fiduccia and Mattheyses, 1982), multilevel graph partitioning (Karypis and Kumar, 1998; Barat et al., 2018), etc.
+We next briefly introduce these partition methods and corresponding PSP indexes.
+Spectral Partitioning.
+Spectral Bisection (SB) (Pothen et al., 1990) is a powerful but expensive algorithm that can be applied in various applications.
+Its basic idea is to calculate the eigenvector $z_{2}$ corresponding to the second smallest eigenvalue (i.e., Fiedler vector) of the graph Laplacian matrix $L$ and utilizes the median value $\overline{m}$ of it to bisect the graph, for the purpose of equal size.
+It assigns the vertices with an entry larger than $\overline{m}$ to one partition and all other vertices to the other.
+The computation of $z_{2}$ is usually done by Lanczos algorithm (Lanczos, 1950) which is quite time-consuming.
+Barnard and Simon (Barnard and Simon, 1994) propose a multi-level Recursive Spectral Bisection (RSB) method to improve the efficiency of SB and achieved about an order-of-magnitude improvement without any loss in the quality of the edge-cut.
+Leete et al.
+extend RSB method to handle an arbitrary number of partitions and propose a parallel implementation of it to improve efficiency.
+Hendrickson and Leland (Hendrickson and Leland, 1995a) extend the bisection problem to $k$-way partitioning by using multiple eigenvalues, which can produce better results than recursive bisection when $k$ is 4 or 8.
+Overall, although spectral partitioning is a powerful method that is applied in many applications, it is very time-consuming and cannot scale to large-scale graphs.
+Growing-based Partitioning.
+The bisection version of graph growing-based partitioning bisects a graph by picking a vertex and growing the region centered at it in a breath-first fashion until half of the vertices are included (Buluç et al., 2016).
+Similarly, the extension to $k$ partitions conducts the $k$ breadth-first searches from $k$ seed vertices until all vertices are traversed (Goehring and Saad, 1994).
+The partition quality of graph-growing algorithms is sensitive to the initial seed vertices.
+Therefore, some methods such as using nodes from a two-way partitioning, and finding the “farthest” vertex from current seeds are explored to find the initial seed vertices (Goehring and Saad, 1994).
+Diekmann et al.
+(Diekmann et al., 2000) extend the graph growing algorithm to its iterative version Bubble, which tries to adjust the seed vertices to the centers of their regions and restart the growing process in each iteration.
+Such a procedure continues until the seed vertices are close to the centers for all partitions, thus producing higher partition quality.
+However, Bubble cannot guarantee balanced partitions.
+Distance measures-based improvements for Bubble are proposed by (Schamberger, 2004; Meyerhenke et al., 2006) to improve the partition quality.
+Due to the complexity caused by iterations, the graph growing partitioning is time-consuming when the graph is large.
+Flow-based Partitioning.
+Flow-based Partitioning is based on the well-known max-flow min-cut theorem (Ford and Fulkerson, 1956), which tries to find a minimum cut between two parts by computing the maximum flow.
+It is obvious that such an approach cannot ensure the balance requirement of partitioning.
+Therefore, flow-based partitioning is often used as a building block for local improvement in many advanced partitioning algorithms.
+For example, as a famous partitioning algorithm for road networks, PUNCH (Delling et al., 2011b) leverages flow-based partitioning to improve the partition quality.
+It utilizes the max-flow computation to find natural cuts (i.e., a sparse cut separating a local region from the rest of the graph) which are geographically formed by road networks.
+To be specific, it first randomly picks a vertex $v$ as a center from a certain area and finds the ring and core of $v$ by growing a BFS from $v$.
+Then, we contract the ring and core into two vertices $s$ and $t$, and compute the $s$-$t$ cut algorithm such as the push-relabel method (Goldberg and Tarjan, 1988) to get the min-cut (i.e., natural cut) of this area.
+Apart from natural cut detection, other techniques such as contraction and re-balancing are also leveraged to get the final partition result.
+PUNCH achieves very high partition quality on road networks because of the preservation of the natural cuts and the contraction of dense regions.
+However, it is slower than many multilevel graph partitioning algorithms such as METIS (Karypis and Kumar, 1998).
+Besides, it is spatial-aware and thus is only applicable on road networks.
+Geometric Partitioning.
+Geometric partitioning solely leverages the coordinate information of the graph to generate the partition, targeting at minimizing some related concepts such as the size of subdomain boundary (Schloegel et al., 2000).
+Recursive Coordinate Bisection (RCB) (Berger and Bokhari, 1987; Simon, 1991) is one of the basic algorithms to solve load-balancing problems in multi-dimensional space by recursively collapsing each dimension.
+In each recursion, RCB divides the computational domain into two partitions by a cutting plane perpendicularly to the coordinate direction of the longest expansion of the domain.
+Many theoretical works (Teng and Points, 1991; Miller et al., 1993, 1998; Gilbert et al., 1998) provide a performance guarantee of partitioning on geometrically defined graphs.
+Space-filling curve (Pilkington and Baden, 1994; Sagan, 2012) is another representative geometric partitioning method that maps a $d$-dimensional space to a $1$-dimensional line to preserve the locality of vertices.
+Overall, geometric partitioning is only suitable for road networks and is conducive to applications that are often defined geometrically.
+Node-Swapping Heuristic.
+Kernighan–Lin (KL) (Kernighan and Lin, 1970) is one of the most famous heuristic algorithms that iteratively swaps nodes of different partitions to improve the partition quality.
+It first defines the edge-cut gain of exchanging two vertices from two partitions.
+Based on the edge-cut gain function, it greedily chooses a set of the pairs to maximize the improvement of the partitions in each iteration until no further gain can be achieved.
+The complexity of KL algorithm is $O(n^{2}\log n)$.
+Fiduccia-Mattheyses (FM) (Fiduccia and Mattheyses, 1982) algorithm improves the expensive complexity of KL algorithm to $O(m)$ by several optimizations.
+Specifically, instead of swapping pairs of vertices, the FM algorithm moves one vertex with the greatest gain from one partition to another at each time and locks it for no further relocation.
+Moreover, the elaborately-designed data structures are utilized to improve efficiency.
+In general, the node-swapping heuristic usually serves as a local improvement for multilevel graph partitioning.
+For example, METIS (Karypis and Kumar, 1998) adopts the improved KL/FM algorithm by confining the scope of gain re-computation to the boundary vertices to improve efficiency.
+Multilevel Graph Partitioning.
+Multilevel Graph Partitioning (MGP) scheme (Barnard and Simon, 1994; Hendrickson and Leland, 1995b) is one of the most powerful approaches that can produce high-quality partitions efficiently.
+It generally includes three phases: coarsening phase, partitioning phase, and uncoarsening phase.
+The coarsening phase sequentially collapses the graph into a coarser one with fewer vertices (normally ends with hundreds of nodes).
+The second phase partitions the coarser graph by utilizing the algorithms mentioned above while the last phase projects the partition of the coarser graph back to the original graph with refinement during each projection step.
+Classic multilevel algorithms include Scotch (Pellegrini and Roman, 1996), METIS (Karypis and Kumar, 1998), KaPPa (Holtgrewe et al., 2010), PaToH (Çatalyürek and Aykanat, 2011) and so on.
+For example, METIS (Karypis and Kumar, 1998) adopts heavy edge matching in the coarsening phase, improved KL/FM algorithm in the partitioning phase, and boundary KL refinement in the uncoarsening phase.
+In particular, the heavy edge matching matches the vertex $v$ for $u$ such that $|e(u,v)|$ is maximum over all the adjacent edges of $u$.
+The intuition behind heavy edge matching is that incorporating adjacent vertices with heavy edge weight results in lightweight edges being the potential edge cuts to decrease the cut size.
+In the partitioning phase, METIS improves the efficiency of KL algorithm by early termination condition, i.e., we stop exchanging vertices between partitions when the edge-cut does not decrease for several iterations.
+Due to the high efficiency and reasonable partition quality, METIS is widely used as a building block for many downstream applications.
+Overall, MGP can not only generate partitions with high quality but also have significant efficiency, thus they are widely used in many applications.
+In-Memory and Streaming Edge Partitioning.
+Over the past decade, the vertex-cut partitioning (a.k.a.
+edge partitioning) that cuts the vertex and targets minimizing the replication factor is drawing much attention as it is more effective than edge-cut partitioning in reducing the communication volume of distributed query processing on skewed power-law graphs (Mayer and Jacobsen, 2021; Bourse et al., 2014).
+We divide edge partitioning methods into two categories: in-memory and online streaming algorithms.
+Intuitively, in-memory algorithms need to load the whole graph into the main memory and hence can produce higher partition quality than online streaming methods as global information is available (Zhang et al., 2017; Mayer and Jacobsen, 2021).
+In contrast, online streaming algorithms ingest edge stream and perform on-the-fly partitioning based on partial information of the graph and thus generate relatively low-quality partitions (Mayer and Jacobsen, 2021; Kong et al., 2022).
+One of the representatives of the in-memory algorithms is the Neighbor Expansion (NE) algorithm proposed by Zhang et al.
+(Zhang et al., 2017), which aims to decrease the replication factor by greedily maximizing edge locality.
+Hybrid Edge Partitioner (HEP) (Mayer and Jacobsen, 2021) improves the efficiency and memory overhead of NE by pruned graph representation and lazy edge removal strategy.
+In addition, it combines NE with streaming algorithm HDRF (Petroni et al., 2015) in a framework to take advantage of them.
+Streaming algorithms aim to improve partitioning efficiency and reduce the memory overhead.
+The simplest streaming algorithm is Random (Gonzalez et al., 2012) which randomly assigns edges to a partition based on hashing, thus having limited partition quality.
+Therefore, heuristic-based methods such as HDRF (Petroni et al., 2015) and CLUGP (Kong et al., 2022) are exploited to improve the partition quality.
+4.1.2. PSP Indexes with Planar Partitioning
+Typical PSP indexes adopting planar partitioning include Arc-Flag (ich Lauther, 2006; Hilger et al., 2009), TNR (Bast et al., 2007), PbS (Chondrogiannis and Gamper, 2014), ParDiSP (Chondrogiannis and Gamper, 2016), and BMHPS (Zhou et al., 2024a), etc..
+We next briefly review some of them.
+Arc-Flag.
+Arc-flag (ich Lauther, 2006) proposed by Lauther is a partition-based arc labeling that utilizes a goal-directed heuristic to speed up Dijkstra’s algorithm.
+It divides a road network into several regions $P=\{p_{j}|j=1,...,k\}$ by rectangular geographic partition and precompute the arc flags whether an edge $e\in E$ is on any shortest path into the region $p_{j}\in P,j=1,...,k$.
+During the query phase, if the target vertex $t$ locates in partition $p_{t}$, then only the edge with the true arc flag of $p_{t}$ will be searched, thereby narrowing down the search space of Dijkstra’s algorithm.
+The query efficiency can be further improved by leveraging the bidirectional technique.
+(Hilger et al., 2009) explores the effect of graph partitioning methods including grid, quadtree, $kd$ tree, and METIS in Arc-Flag.
+It turns out that $kd$ tree and METIS are better than the others.
+TNR.
+Transit Node Routing (TNR) (Bast et al., 2007) is a famous framework for processing long-distance shortest path queries on road networks.
+It is based on the observation that when traveling to a distant location, you will depart from your current location through one of the limited and important traffic junctions.
+In other words, for every vertex $v$, there are a small set of vertices that $v$ has to pass through one of them to access a distant vertex.
+We call such a small set of vertices as the access node of $v$, denoted as $A(v)$.
+The union of all access nodes forms the transit node $\mathcal{T}$ of graph $G$.
+The idea of TNR is to leverage grid-based geometric partitioning to obtain subnetworks and then precompute distances between $v$ and $u\in A(v)$, and the distances between all transit nodes.
+Given a query $Q(s,t)$, a locality filter is used to decide whether such a query is a local query or not.
+If it is a long-distance query then we can answer it by $Q(s,t)=\min_{u\in A(s),w\in A(t)}\{d(s,u)+d(u,w)+d(w,t)\}$.
+Otherwise, $Q(s,t)$ is answered by another shortest path algorithm.
+TNR demonstrates a good performance on long-distance queries on large road networks.
+Nonetheless, its preprocessing is very expensive.
+PbS and ParDiSP.
+PbS (Chondrogiannis and Gamper, 2014) partitions the road network into several parts by METIS (Karypis and Kumar, 1998) and constructs an overlay graph among boundary vertices for query processing.
+In particular, it PbS builds in-component shortcuts for each vertex $v\in G_{i}$ (i.e., the shortcuts from $v$ to its boundary vertices in $B_{i}$).
+Based on the in-component shortcuts, an overlay graph $H$ comprising a complete clique for each partition is constructed.
+To respond to a query about an OD pair $(s,t)$, the modified bidirectional search is performed from $s$ to $t$ if they belong to the same partition.
+In cases where they do not, $Q(s,t)$ is calculated as $\min_{b_{i}\in B_{s},b_{j}\in B_{t}}\{d_{G_{s}}(s,b_{i})+d_{H}(b_{i},b_{j})+d_{G_{t}}(t,b_{j})\}$, where $B_{s}$ and $B_{t}$ represent the boundary vertex set of $G_{s}$ and $G_{t}$, respectively.
+ParDiSP (Chondrogiannis and Gamper, 2016) improves the query efficiency of PbS by modifying the overlay graph and the query processing method.
+Specifically, after partitioning the graph $G$ by METIS, it builds an extended component for each partition $G_{i}$.
+These components denoted as $G^{*}_{i}$, include all the vertices and edges in $G_{i}$ as well as those located on the shortest path between two boundary vertices of $G_{i}$.
+The overlay graph of ParDiSP consists of all shortest paths between all boundary vertices of each partition $G_{i}$, along with the inter-edges.
+To answer the query $Q(s,t)$, the ALT algorithm (Goldberg and Harrelson, 2005) is employed on the extended component $G^{*}_{i}$ if $s,t\in G_{i}$.
+Additionally, the boundary vertices of $G_{i}$ are utilized as landmarks for ALT.
+For cross-partition queries, ParDiSP adopts the same query strategy as PbS.
+In summary, planar PSP indexes leverage graph partitioning for shortest path computation in small-treewidth graphs such as road networks, which can achieve good query performance.
+Additionally, graph partitioning is the only solution that enables the more complicated shortest path problems like time-dependent (Li et al., 2019c, 2020b), constraint paths (Liu et al., 2021, 2022, 2023), and multi-criteria shortest paths (Zhou et al., 2024a) to construct their indexes in large networks.
+Figure 3. Different Partition Structures
+4.2. Core-Periphery Partitioning
+Core-periphery Partition (CP) (Borgatti and Everett, 2000; Rombach et al., 2014; Elliott et al., 2020; Maehara et al., 2014a) treats the partitions discriminately by taking some important vertices as “Core” and the remaining ones as “Peripheries” with the formal definition as follows.
+Definition 0.
+(Core-Periphery Partitioning).
+Given a graph $G$, core-periphery partitioning decomposes $G$ into two distinct parts: a core subgraph $G_{c}$ that contains the most important vertices, and $k-1$ peripheral subgraphs $\{G_{i}|1\leq i\leq k-1\}(k\in[2,\infty])$, satisfying $\bigcup_{i\in[1,k-1]}V(G_{i})\cup V(G_{c})=V,V(G_{c})\cap V(G_{i})=\emptyset(\forall 1\leq i\leq k-1)$ and $V(G_{i})\cap V(G_{j})=\emptyset(\forall i,j\in[1,k-1],i\neq j)$.
+There are two big streams depending on how the core is formed: Core-Tree Decomposition (Maehara et al., 2014a, b; Li et al., 2020a; Zheng et al., 2022; Zhou et al., 2024b) and Sketch (Potamias et al., 2009; Das Sarma et al., 2010; Gubichev et al., 2010; Tretyakov et al., 2011; Qiao et al., 2012; Qi et al., 2013; Farhan et al., 2018; Wang et al., 2021; Farhan et al., 2022).
+It is worth noting that these two core-periphery partitioning is tightly integrated with corresponding PSP indexes.
+As such, it is hard for the core-periphery PSP index to choose other partition methods (e.g., planar partitioning) for graph partitioning.
+Core-Tree Decomposition (Maehara et al., 2014a) generates the core through tree decomposition (Robertson and Seymour, 1984), and the resulting periphery part is a set of small-width trees.
+Specifically, it leverages minimum degree elimination (Xu et al., 2005; Berry et al., 2003) to contract vertices of lower degree first such that the graph is contracted from the periphery towards the center, generating a set of growing trees (“Peripheries”) around a shrinking while denser graph (“Core”) formed by those non-contracted vertices.
+The contraction terminates when the width of one tree reaches the previously-set bandwidth threshold.
+A classic PSP index adopting core-tree decomposition is Core-tree (CT) index (Li et al., 2020a; Zhou et al., 2024b), which aims to scale up the distance labeling on graphs with core-periphery properties such as small-world networks.
+It leverage core-tree decomposition to divide the graph $G$ into a “core” (i.e., a large partition of high-degree vertices) and a set of “trees” (i.e., a number of small partitions with bounded treewidth).
+To achieve efficient shortest path computation, a 2-hop labeling index (i.e., PLL (Akiba et al., 2013)) is built on the core, while a tree decomposition-based index (i.e., H2H (Ouyang et al., 2018)) is built on the “trees”.
+There are several reasons behind such a differentiation strategy.
+Firstly, lower-degree vertices tend to have larger label sizes compared to higher-degree ones since the high-degree vertices are more likely to be passed through by the shortest path of a random source-target query.
+Therefore, by removing them from the core, the index size of 2-hop labeling can be significantly reduced.
+Secondly, H2H is unsuitable for large-treewidth graphs (the core) as the tree decomposition method cannot handle these graphs.
+However, it is very efficient on small-treewidth networks, which makes it an ideal SP index for the peripheries.
+DCT (Zhou et al., 2024b) extends CT to dynamic networks by efficient index maintenance algorithms.
+Sketch (Potamias et al., 2009; Das Sarma et al., 2010; Gubichev et al., 2010; Tretyakov et al., 2011; Qiao et al., 2012; Qi et al., 2013; Farhan et al., 2018; Wang et al., 2021; Farhan et al., 2022) selects a set of vertices as landmarks, which can be regarded as the core and treats the remaining parts as a periphery.
+A classic PSP index adopting Sketch is QbS (Wang et al., 2021) which picks a set of landmarks as meta-graph (i.e., the core) while the other part as sparsified graph.
+The 2-hop labeling is computed for the meta-graph and a sketch is obtained based on the meta-graph to summarize the important structure of shortest paths for the query pair.
+The shortest path computation on the sparsified graph is based on a bidirectional search and guided by the sketch.
+Similarly, BatchHL (Farhan et al., 2018, 2022) also selects a set of landmarks and then builds a highway index on these landmarks.
+Besides, a highway cover labeling is built to connect the periphery vertices to at least one landmark to prune the search space during query processing.
+(Farhan et al., 2022) investigates the batch updates for BatchHL in unweighted networks.
+In summary, the identification of a small yet crucial set of vertices as the “core” of the network, with all other vertices being the “periphery”, allows core-periphery PSP indexes to achieve exceptional scalability in computing shortest paths on small-world networks.
+Core-Tree (Li et al., 2020a) index improves the scalability of state-of-the-art non-partitioned PSL (Li et al., 2019b) with a negligible performance drop in the query time, while Sketch-based methods (Farhan et al., 2018; Wang et al., 2021) have the best scalability for shortest path computation but sacrificing the query efficiency due to the search on periphery.
+4.3. Hierarchical Partitioning
+Hierarchical Partitioning (HP) organizes the network partitions hierarchically and each level is equivalent to a planar partitioning. We formally define it as follows.
+Definition 0.
+(Hierarchical Partitioning).
+Given a graph $G$, hierarchical partitioning method organizes $G$ hierarchically within $H$ levels where each level $h$ is a planar partitioning $\{G^{h}_{i}\}$ of $G$ and $\exists G^{h-1}_{j}\supset G^{h}_{i},\forall G^{h}_{i}(1\leq h\leq H)$.
+HP aims to organize the partition results in multiple layers so that each layer only needs to focus on the search/index of this level.
+Such a level-by-level approach is effective in pruning the search space for faster query processing.
+Since each level of hierarchical partitioning is equivalent to planar partitioning, we further divide it into two categories according to different cut types: edge-cut based HP that cuts edges and vertex-separator based HP that cuts vertices.
+A general way to obtain hierarchical partitioning is to leverage multi-level graph partition methods such as METIS (Karypis and Kumar, 1998) since they naturally produce multiple levels of partitions during the partitioning process.
+Classic PSP indexes adopting edge-cut based HP include G-Tree (Zhong et al., 2015), G*-Tree (Li et al., 2019a), CRP (Delling et al., 2017), HiTi (Jung and Pramanik, 1996), and SHARC (Bauer and Delling, 2010), etc.
+While PSP indexes adopting vertex-separator based HP include SMOG (Holzer et al., 2009) and ROAD (Lee et al., 2010), etc.
+We next briefly introduce some classic ones.
+G-Tree.
+G-Tree (Zhong et al., 2013) is a balanced search tree for location-based queries on road networks (e.g., point-to-point shortest path query, $k$ nearest neighbor ($k$NN) query, and keyword-based $k$NN query) by hierarchically organizing the partitions produced by METIS (Karypis and Kumar, 1998).
+Specifically, they first partition graph $G$ into $f$ subgraphs with roughly equal size and take them as the root’s children nodes, and then recursively partition them into another $f$ children nodes until each leaf node’s subgraph has no more than $\tau$ vertices.
+By precomputing the shortest path distance of the border-border pairs in the non-leaf nodes and border-vertex pairs in the leaf nodes, an assembly-based method is proposed for efficiently retrieving the shortest distance between two vertices from G-tree.
+In particular, when $s$ and $t$ are in the same leaf node (i.e., $s,t\in G_{i}$), there are two subcases for $sp(s,t)$: (1) $sp(s,t)$ does not contain a vertex outside the leaf node $G_{i}$.
+In this case, we can obtain $d(s,t)$ by performing Dijkstra’s algorithm in $G_{i}$.
+(2) $sp(s,t)$ contains a vertex outside the leaf node $G_{i}$.
+In this case, $d(s,t)$ can be computed by $\min_{b_{1},b_{2}\in B_{i}}\{d(s,b_{1})+d(b_{1},b_{2})+d(b_{2},t)\}$.
+For the case that $s$ and $t$ belong to different leaf nodes (i.e., $s\in G_{s}$ and $t\in G_{t}$), we implement a dynamic programming algorithm that searches from $G_{s}$ to $G_{LCA}$ and then to $G_{t}$, where $G_{LCA}$ is the lowest common ancestor (LCA) of $G_{s}$ and $G_{t}$.
+G*-Tree (Li et al., 2019a) further improve the query efficiency of G-Tree by adding shortcuts among the lowest level partitions.
+CRP.
+To answer the shortest path queries on the road networks with arbitrary metrics, Delling et al.
+propose a novel partition-based method named Customizable Route Planning (CRP) (Delling et al., 2011a, 2017) to support real-time traffic updates and personalized optimization functions.
+The general idea of CRP is to decouple topological and metric properties of the road network by graph partitioning.
+In particular, CRP has three main stages: metric-independent preprocessing, metric customization, and query stage.
+Metric-independent preprocessing decomposes graph $G$ into multi-level partitions by PUNCH (Delling et al., 2011b) in top-down fashion.
+We first generate top-level partitions, while the lower-level partitions are obtained by running PUNCH on each subgraphs of the higher level.
+Suppose the top-level partitions are $\{G_{i}|i=1,...,k\}$, then in metric customization stage, a shortest-distance clique between the boundary vertices $B_{i}$, denoted as $C_{i}$, is precomputed for each partition $G_{i}$.
+These boundary distances are computed in a bottom-up fashion for all levels.
+The overlay graph on all boundary vertices is also constructed.
+It is worth mentioning that the overlay graph will be further optimized by pruning techniques such as edge reduction and skeletons, resulting in a much-sparsified graph called skeleton graph.
+Finally, bidirectional search is used for query processing in the query stage.
+In summary, hierarchical PSP indexes organize the partitions in a tree style hierarchically, allowing for the exploration of pathfinding in a layer-by-layer manner.
+As the state-of-the-art method in this line of research, G-Tree (Zhong et al., 2013; Li et al., 2019a) achieves good shortest path query performance on road networks.
+Nonetheless, it cannot outperform tree decomposition-based methods such as H2H (Ouyang et al., 2018) and is unsuitable for small-world networks.
+Table 1. Structure-Based Partition Method Classification.
+Structure Partition Method Objectives Spatial-Aware Cut Type
+Category Representative Methods Balance Minimal Cut
+PP Spectral partition SP (Pothen et al., 1990), RSB (Barnard and Simon, 1994) $\checkmark$ $\checkmark$ $\times$ Edge
+Growing-based partition Bubble (Diekmann et al., 2000) $\times$ $\checkmark$ $\times$ Edge
+Flow-based partition PUNCH(Delling et al., 2011b) $\checkmark$ $\checkmark$ $\checkmark$ Edge
+Geometric partition RCB (Simon, 1991; Huang et al., 1996) $\checkmark$ $\checkmark$ $\checkmark$ Edge
+Node-swapping heuristic KL (Kernighan and Lin, 1970), FM (Fiduccia and Mattheyses, 1982) $\checkmark$ $\checkmark$ $\times$ Edge
+Minimum $k$-cut $k$-cut (Goldschmidt and Hochbaum, 1994) $\times$ $\checkmark$ $\times$ Edge
+Multilevel graph partition SCOTCH (Pellegrini and Roman, 1996), METIS (Karypis and Kumar, 1998), KaHyPar (Akhremtsev et al., 2017) $\checkmark$ $\checkmark$ $\times$ Edge
+Streaming edge partition HDRF (Petroni et al., 2015; Mayer and Jacobsen, 2021), CLUGP (Kong et al., 2022) $\checkmark$ $\checkmark$ $\times$ Vertex
+In-Memory edge partition NE(Zhang et al., 2017), HEP(Mayer and Jacobsen, 2021) $\checkmark$ $\checkmark$ $\times$ Vertex
+CP Core-tree decomposition Core-Tree(Maehara et al., 2014b; Li et al., 2020a; Zheng et al., 2022; Elliott et al., 2020; Zhou et al., 2024b) $\times$ $\times$ $\times$ Edge
+Sketch QbS (Wang et al., 2021), BatchHL (Farhan et al., 2018, 2022) $\times$ $\times$ $\times$ Edge
+HP Edge-cut based HP HiTi (Jung and Pramanik, 1996) $\times$ $\times$ $\checkmark$ Edge
+SHARC (Bauer and Delling, 2010), G-Tree (Zhong et al., 2015), CRP (Delling et al., 2017) $\checkmark$ $\checkmark$ $\times$ Edge
+Vertex-Separator based HP SMOG (Holzer et al., 2009), ROAD  (Lee et al., 2009, 2010) $\times$ $\times$ $\times$ Vertex
+PP is Planar Partitioning, CP is Core-Periphery Partitioning, and HP is Hierarchical Partitioning.
+Balance denotes if it targets at balanced partitioning, Minimum Cut denotes if it targets at minimize the cut size, Spatial-Aware denotes if it needs geographic coordinates, and Cut Type is the cut type with Edge denotes edge-cut and Vertex denotes vertex-cut.
+4.4. Summary and Discussion
+We present the partition results of different partition structures in Figure 3, where the red vertices represent the corresponding boundary vertices.
+Specifically, planar partitioning generates four partitions with equivalent vertex size while core-periphery partitioning produces both “Core” and “Periphery” (the periphery of Core-Tree decomposition is a set of small-width trees while Sketch treats the non-core part as a periphery).
+Hierarchical partitioning organizes partitions hierarchically, and the leaf nodes (the lowest level partition result) may have the same partition result as planar partitioning if the same MGP method is used.
+We also summarize and categorize existing partition methods under our structure-based classification, as shown in Table 1.
+Note that for CP and HP, we use the corresponding PSP indexes as representative methods, given the absence of specific partitioning methods associated with them.
+Almost all planar partition methods adopt minimal cut size as the optimization objective because a smaller cut size could facilitate cross-partition communication.
+Some partition methods also consider the size balance of different partitions for a more balanced workload distribution.
+Some of them also require spatial coordinates, confining their applications to road networks.
+We next analyze the requirement of the PSP index and discuss the pros and cons of different partition structures.
+For the PSP index, the boundary vertex number is a crucial factor in determining the index construction time, query time, and update time.
+However, the planar and hierarchical partitions have no limit to the boundary number, even though reducing the cut size is one of their optimization goals.
+By contrast, the core-periphery partition limits the boundary number by pre-defined bandwidth or limited landmark number such that their performance would not deteriorate by the large border number.
+Therefore, the planar and hierarchical partitions are better for small-treewidth networks (e.g., road networks) such that balanced partitions and fewer boundaries can be achieved together; while core-periphery is better for large-treewidth networks (e.g., social networks) as it limits the boundary vertex number by bandwidth or landmark number.
+5. Novel Partitioned Shortest Path Strategies
+Apart from the shortest path algorithms and partition methods, another key problem for the PSP index is how to organize and coordinate the shortest path indexes of different partitions for correct index construction, query processing, and index maintenance.
+We call the strategy for solving the above problem as PSP Strategy.
+In this section, we first summarize the classic pre-boundary PSP strategy and then propose two novel PSP strategies (no-boundary and post-boundary strategy) to improve the index construction and maintenance efficiency of existing solution with non-trivial correctness guarantee.
+In addition, we also propose a pruning-based overlay optimization to further enhance these strategies.
+Note that we use planar PSP index for illustrating the PSP strategy in this section for simplicity.
+5.1. Traditional Pre-Boundary Strategy
 The PSP index $L$ typically consists of two components: the partition indexes $\{L_{i}\}$ for each subgraph (partition) $G_{i}$, and the overlay index $\tilde{L}$ for the overlay graph which is composed of the boundary vertices of all partitions, i.e., $L=\{L_{i}\}\cup\tilde{L}$.
-Regardless of the specific SP algorithm, almost all the existing PSP indexes (Jung and Pramanik, 2002; Delling et al., 2017; Holzer et al., 2009; Chondrogiannis and Gamper, 2014, 2016; Zhong et al., 2013; Li et al., 2019a; Zhong et al., 2015; Ta et al., 2017; Wang et al., 2019, 2016; Liu et al., 2021, 2022b, 2022a; Li et al., 2019c, 2020b; Li et al., 2020a; Zheng et al., 2022; Wang et al., 2021; Farhan et al., 2022; Gubichev et al., 2010; Qiao et al., 2012) construct the PSP index and process the queries under the following procedure.
-Figure 2. Example Graph $G$ and Overlay Graph $\tilde{G}$
+Regardless of the specific SP algorithm, almost all the existing PSP indexes (Jung and Pramanik, 2002; Delling et al., 2017; Holzer et al., 2009; Chondrogiannis and Gamper, 2014, 2016; Zhong et al., 2013; Ta et al., 2017; Wang et al., 2019, 2016; Liu et al., 2021, 2022; Li et al., 2019c, 2020b; Wang et al., 2021; Farhan et al., 2022; Gubichev et al., 2010; Qiao et al., 2012) construct the PSP index and process the queries under the following procedure.
 Index Construction.
-It first leverages a partition method to divide the graph $G$ into multiple subgraphs $\{G_{i}\}$ and then builds the PSP index $L$ by the following four steps.
-As illustrated in Figure 3, Step 1 precomputes the global distance between all boundary vertex pairs $(b_{i1},b_{i2}),(b_{i1},b_{i2}\in B_{i})$ for each partition $G_{i}$ and insert shortcuts $e(b_{i1},b_{i2})=d_{G}(b_{i1},b_{i2})$ into $G_{i}$ to get $G^{\prime}_{i}$.
-For instance, the global distance between boundary vertex pairs $(v_{3},v_{10})\in G_{1}$, $(v_{8},v_{9})\in G_{2}$, $(v_{5},v_{12})\in G_{3}$, $(v_{6},v_{11})\in G_{4}$ are calculated and inserted into their corresponding subgraphs to form new subgraphs $G^{\prime}_{1},G^{\prime}_{2},G^{\prime}_{3},G^{\prime}_{4}$, respectively; Step 2 constructs the SP index $L_{i}$ based on $G^{\prime}_{i}$.
-The index can be any type like CH (Geisberger et al., 2008), PLL (Akiba et al., 2013) or H2H (Ouyang et al., 2018), and the index construction procedure among multiple partitions can be parallelized as it takes each partition as input independently; Step 3 constructs the overlay graph $\tilde{G}$ based on the precomputed shortcuts in Step 1.
-Specifically, $\tilde{G}$ is composed of shortcuts between boundary vertex pairs $\{B_{i}\times B_{i}\}$ ($(v_{3},v_{10}),(v_{8},v_{9}),(v_{5},v_{12}),(v_{6},v_{11})$) and the inter-edge set $E_{inter}$ ($(v_{3},v_{5})$, $(v_{10},v_{12})$, $(v_{9},v_{10})$,$(v_{9},v_{12})$, $(v_{9},v_{11})$, $(v_{8},v_{11})$, $(v_{12},v_{11})$, $(v_{12},v_{6})$), that is $V_{\tilde{G}}=\{B_{i}\},E_{\tilde{G}}=\{B_{i}\times B_{i}\}\cup E_{inter}$; Step 4 constructs the SP index $\tilde{L}$ on $\tilde{G}$.
-Similar to $L_{i}$, $\tilde{L}$ can be any index type.
-Figure 3. Traditional PSP Strategy Example
+It first leverages a partition method to divide the graph $G$ into multiple subgraphs $\{G_{i}\}$ and then builds the PSP index $L$ by the following four steps, as illustrated in Figure 4.
+Step 1 precomputes the global distance between all boundary vertex pairs $(b_{i1},b_{i2}),(b_{i1},b_{i2}\in B_{i})$ for each partition $G_{i}$ and insert shortcuts $e(b_{i1},b_{i2})=d_{G}(b_{i1},b_{i2})$ into $G_{i}$ to get $G^{\prime}_{i}$.
+For instance, the global distance between boundary vertex pairs $(v_{3},v_{10})\in G_{1}$, $(v_{8},v_{9})\in G_{2}$, $(v_{5},v_{12})\in G_{3}$, $(v_{6},v_{11})\in G_{4}$ are calculated and inserted into their corresponding subgraphs to form new subgraphs $G^{\prime}_{1},G^{\prime}_{2},G^{\prime}_{3},G^{\prime}_{4}$, respectively;
+Step 2 constructs the partition index $L_{i}$ based on $G^{\prime}_{i}$.
+The index can be CH (Geisberger et al., 2008), PLL (Akiba et al., 2013) or H2H (Ouyang et al., 2018), and the index construction procedure among multiple partitions can be parallelized as it takes each partition as input independently;
+Step 3 constructs the overlay graph $\tilde{G}$ based on the precomputed shortcuts in Step 1.
+Specifically, $\tilde{G}$ is composed of shortcuts between boundary vertex pairs $\{B_{i}\times B_{i}\}$ and the inter-edges, that is $V_{\tilde{G}}=\{B_{i}\},E_{\tilde{G}}=\{B_{i}\times B_{i}\}\cup E_{inter}$;
+Step 4 constructs the overlay index $\tilde{L}$ on $\tilde{G}$. Similar to $L_{i}$, $\tilde{L}$ can be any index type.
+Figure 4. Traditional PSP Strategy (Pre-boundary Strategy)
 Note that the construction of $\tilde{L}$ (Step 3, Step 4) can be parallelized with $L_{i}$ (Step 2) since they are independent and both rely on Step 1.
-All pairs of boundaries are independent of each other so we only need $|B|$ times of Dijkstra’s with time complexity $O(nlogn+m)$.
-Then each partition’s label $L_{i}$ can be constructed in parallel, and $\tilde{L}$’s construction is also independent to them, so its complexity is the worst case of them: $max\{O_{c}(G_{i}),O_{c}(\tilde{G})\}$, where $O_{c}$ is the complexity of underlying index’s construction time as our discussion is not fixed to any specific index type.
-Query Processing. By referring to the cut property of boundary vertex, the shortest path queries are divided into two categories and processed by utilizing $L$:
-Case 1: Same-Partition i.e., $\forall s,t\in G_{i}$, $Q(s,t)=d_{L_{i}}(s,t)$;
-Case 2: Cross-Partition, i.e., $\forall s\in G_{i},t\in G_{j}(i\neq j),Q(s,t)=$
-$\begin{cases}d_{\tilde{L}}(s,t)&s,t\in B\\ \min\limits_{b_{q}\in B_{j}}\{d_{\tilde{L}}(s,b_{q})+d_{L_{j}}(b_{q},t)\}&s\in B,t\notin B\\ \min\limits_{b_{p}\in B_{i}}\{d_{L_{i}}(s,b_{p})+d_{\tilde{L}}(b_{p},t)\}&s\notin B,t\in B\\ \min\limits_{b_{p}\in B_{i},b_{q}\in B_{j}}\{d_{L_{i}}(s,b_{p})+d_{\tilde{L}}(b_{p},b_{q})+d_{L_{j}}(b_{q},t)\}&s\notin B,t\notin B\\ \end{cases}$
+For Step 1, we need $|B|$ times of Dijkstra’s with time complexity $O(|B|\cdot(n\log n+m))$.
+Then each partition’s label $L_{i}$ can be constructed in parallel, and $\tilde{L}$’s construction is also independent of them, so its complexity is the worst case of them: $max\{O_{c}(G_{i}),O_{c}(\tilde{G})\}$, where $O_{c}$ is the complexity of underlying index’s construction time as our discussion is not fixed to any specific index type.
+Query Processing. The shortest path queries can be divided into two categories and processed by utilizing $L$:
+[…]
+[…]
 In summary, when $s$ and $t$ are in the same partition, we can use $L_{i}$ to answer $d_{G}(s,t)$; otherwise, we have to use $L_{i},L_{j}$ and $\tilde{L}$.
 The same-partition query complexity is $O_{q}(L_{i})$, where $O_{q}$ is the index’s query complexity.
-The cross-partition query is made up of three parallel query sets, and the complexity is the worst of them: $max\{B_{i}\times O_{q}(L_{i}),B_{i}\times B_{j}\times O_{q}(\tilde{L}),B_{j}\times O_{q}(L_{j})\}$.
-The proof of the query processing can be found in the Appendix of our extended version (Zhang et al., 2023b).
-2.4. Shortest Path Algorithms
-We summarize the shortest path algorithms as follows: 1) Direct Search such as Dijkstra’s (Dijkstra, 1959) and $A^{*}$ (Hart et al., 1968) searches the graph when the index cannot be constructed due to the large graph size (QbS (Wang et al., 2021), ParDist (Chondrogiannis and Gamper, 2016)) or complicated problem (COLA (Wang et al., 2016)).
-It takes no time in index update, but queries are slowest; 2) Contraction Hierarchy (CH) (Geisberger et al., 2008) is a widely used lightweight index that contracts the vertices in a pre-defined order and preserves the shortest path information by adding shortcuts among the contracted vertex’s neighbors.
-The search-based CH (Geisberger et al., 2008) has a small index size but takes longer to build and maintain (Zhang et al., 2021a; Geisberger et al., 2012) while concatenation-based CH (Ouyang et al., 2020; Wei et al., 2020) is much faster to build and maintain if the tree-width is small.
-CH’s query performance is generally around 10$\times$ faster than the direct search but much slower than the hub labelings.
-Surprisingly, no PSP index has used CH as underlying index; 3) Pruned Landmark Labeling (PLL).
-Although many hub labeling methods have been proposed in the past decade, only two are widely used.
-Built by either pruned search (Akiba et al., 2013) or propagation (Li et al., 2019b; Jin et al., 2020; Zhang et al., 2021b, 2023a), PLL is the only index that can work on the graph with large treewidth.
-Therefore, graphs with this property use it as the underlying index (e.g., QsB (Wang et al., 2021; Farhan et al., 2022) for its landmarks, Core-Tree (CT) (Li et al., 2020a; Zheng et al., 2022) for its core); 4) Tree Decomposition (TD) (Ouyang et al., 2018; Chang et al., 2012; Wei, 2012; Chen et al., 2021; Zhang et al., 2021c).
-As another widely used hop labeling, it is much faster than PLL for graphs with smaller treewidth but cannot scale to large ones.
-For instance, CT (Li et al., 2020a; Zheng et al., 2022) uses it for its periphery, and FHL (Liu et al., 2021, 2022a) use it as a forest; 5) All Pair Pre-computation pre-computes all-pair distance with the fastest query, but incurs long index construction and large index storage.
-G-Tree (Zhong et al., 2013, 2015) and ROAD (Lee et al., 2009, 2010) store the distance between boundary vertices in each partition.
-3. Novel PSP Strategies
-As stated in Section 2, the traditional PSP strategy can be time-consuming for large networks since it requires a laborious computation of all-pair boundary distances using Dijkstra’s search.
-To tackle this problem, we propose two novel PSP strategies and analyze how they could be leveraged to correctly answer the shortest path queries.
-Besides, we design the index update approaches systematically to support the dynamic application of both the existing and novel PSP strategies.
-Lastly, we proposed a pruning technique to slim the overlay graph to achieve better index performance.
-It should be noted that the PSP strategy is a general procedure that any PSP index needs to follow and can work with any type of graph partition method or SP algorithm.
-So we introduce the PSP strategies without mentioning the specific partitioning and SP approach.
-3.1. No-Boundary and Post-Boundary Strategy
+The cross-partition query is made up of three parallel query sets, and the complexity is the worst of them: $max\{|B_{i}|\cdot O_{q}(L_{i}),|B_{i}|\cdot|B_{j}|\cdot O_{q}(\tilde{L}),|B_{j}|\cdot O_{q}(L_{j})\}$.
+The proof of the query processing is in the Appendix of our extended version (Zhang et al., [n.
+d.]).
 Since the traditional partitioned index starts by precomputing the all-pair distance among boundary vertices, we call this approach Pre-Boundary Strategy.
-It seems that this strategy, with the theoretical guarantee for correct partitioned shortest-path processing, could handle all the partitioned queries well.
-Nevertheless, the first step of the Pre-Boundary could be very time-consuming because only the index-free SP algorithms like Dijkstra’s (Dijkstra, 1959) or A (Hart et al., 1968) could be utilized.
+5.2. Novel No-Boundary and Post-Boundary Strategies
+The pre-boundary strategy with the theoretical guarantee for correct partitioned shortest-path processing could handle all the queries well.
+However, the Step 1 of the Pre-Boundary could be very time-consuming because only the index-free SP algorithms like Dijkstra’s (Dijkstra, 1959) or A (Hart et al., 1968) could be utilized.
 As a result, the index construction and maintenance efficiency suffer when the graph has numerous boundary vertex pairs.
-However, it appears that pre-computing the all-pair boundary distance is an essential procedure for constructing the “correct” PSP index, otherwise the correctness of the subgraph index $L_{i}$ cannot be guaranteed because the shortest distance between boundary vertices within one partition could pass through another partition.
-For instance, as shown in Figure 3, $p(v_{6},v_{11})=\left<v_{6},v_{12},v_{11}\right>$ goes outside $G_{4}$ and passes through $v_{12}\in G_{3}$.
+However, it appears that pre-computing the all-pair boundary distance is an essential procedure for constructing the “correct” PSP index.
+Otherwise, the correctness of the subgraph index $L_{i}$ cannot be guaranteed because the shortest distance between boundary vertices within one partition could pass through another partition.
+For instance, as shown in Figure 4, $sp(v_{6},v_{11})=\left<v_{6},v_{12},v_{11}\right>$ goes outside $G_{4}$ and passes through $v_{12}\in G_{3}$.
 As a result, $d(v_{6},v_{11})$ cannot be answered correctly only with $G_{3}$ and the global shortest distance $d(v_{6},v_{11})$ calculation is essential before constructing $L_{3}$.
-Then a question naturally arises: do we really require precomputing the boundary all-pair distance first?
-In other words, does there exist a chance to keep index correctness by dropping this time-consuming step?
+Then, a question arises naturally: do we really require precomputing the boundary all-pair distance first?
+In other words, is there a chance to keep index correctness by dropping this time-consuming step?
 Based on this brainstorm, we break the traditional misconception and propose a novel No-Boundary Strategy as follows, which significantly reduces the index construction/maintenance time by skipping the time-consuming pre-computation step.
@@
 It contains three steps as shown in Figure 5.
-With a graph partitioned into subgraphs $\{G_{i}\}$, we start by constructing the shortest path index $L_{i}$ for each subgraph $G_{i}$ parallelly in Step 1; Then Step 2 construct the overlay graph $\tilde{G}$ based on $\{L_{i}\}$.
+With a graph partitioned into subgraphs $\{G_{i}\}$, we start by constructing the partition index $L_{i}$ for each subgraph $G_{i}$ parallelly in Step 1; Then Step 2 construct the overlay graph $\tilde{G}$ based on $\{L_{i}\}$.
 For instance, $\tilde{G}$ is made up of boundary all-pair edges $(v_{3},v_{10})$, $(v_{8},v_{9})$, $(v_{5},v_{12})$, $(v_{6},v_{11})$ (with their weight calculated based on $L_{1},L_{2},L_{3}$, and $L_{4}$) and inter-edges.
-It should be noted that although the overlay graph $\tilde{G}$ has the same structure as that in the Pre-boundary Strategy, their edge weights differ as $\{L_{i}\}$ of them are constructed from different subgraphs; Step 3 constructs the shortest path index $\tilde{L}$ for $\tilde{G}$.
+It should be noted that although the overlay graph $\tilde{G}$ has the same structure as that in the Pre-boundary Strategy, their edge weights differ as $\{L_{i}\}$ of them are constructed from different subgraphs; Step 3 constructs the overlay index $\tilde{L}$ for $\tilde{G}$.
 Since the partition indexes $\{L_{i}\}$ are constructed in parallel first followed by the construction of $\tilde{L}$, the No-Boundary takes $max\{O_{c}(G_{i})\}+O_{c}(\tilde{G})$ time in index construction.
@@
 Now that $L_{i}$ cannot answer $G_{i}$’s query correctly as the global distance between boundary all-pair is not covered, then how can No-Boundary answer query correctly?
-Before revealing it, we first prove that the correctness of $\tilde{L}$ still holds even though it is built upon the incorrect $\{L_{i}\}$, that is the queries between boundary vertices can be correctly processed with $\tilde{L}$, using:
+Before revealing it, we first use the following theorem to prove that the correctness of $\tilde{L}$ still holds even though it is built upon the incorrect $\{L_{i}\}$, i.e., the queries between boundary vertices can be correctly processed with $\tilde{L}$.
 Theorem 1.
@@
 Proof.
-We divide all the scenarios into three cases as shown in Figure 5-(a): 1) $s,t\in G_{i}$ and $p_{G}(s,t)$ only passes through the interior of $G_{i}$, then it is obvious that $d_{G}(s,t)=d_{G_{i}}(s,t)$ holds.
-Since $\tilde{w}(s,t)=d_{G_{i}}(s,t)$ is leveraged for $\tilde{G}$’s construction, we can obtain that $d_{G}(s,t)=d_{\tilde{G}}(s,t)=d_{\tilde{L}}(s,t)$; 2) $s,t\in G_{i}$ with $p_{G}(s,t)$ going outside of $G_{i}$; 3) $s\in G_{i}$ and $t\in G_{j}(i\neq j)$.
+We divide all the scenarios into three cases as shown in Figure 6-(a): 1) $s$ and $t\in G_{i}$ and $p_{G}(s,t)$ only passes through the interior of $G_{i}$, then it is obvious that $d_{G}(s,t)=d_{G_{i}}(s,t)$ holds.
+Since $\tilde{w}(s,t)=d_{G_{i}}(s,t)$ is leveraged for $\tilde{G}$’s construction, we can obtain that $d_{G}(s,t)=d_{\tilde{G}}(s,t)=d_{\tilde{L}}(s,t)$; 2) $s$ and $t\in G_{i}$ with $p_{G}(s,t)$ going outside of $G_{i}$; 3) $s\in G_{i}$ and $t\in G_{j}(i\neq j)$.
 In the latter two cases, we take the concise form of $p_{G}(s,t)$ by extracting only the boundary vertices as $p_{c}=\left<s,b_{0},b_{1},\dots,b_{n},t\right>$ ($b_{i}\in B,0\leq i\leq n$).
@@
 Figure 6. Different Categories of OD Distribution
-Based on the above Theorem, we continue to answer the shortest distance queries by discussing all scenarios in two cases:
-Case 1: Same Partition $s,t\in G_{i}$, we report $Q(s,t)$ as follows:
+Based on the above Theorem, we answer the shortest distance queries by discussing all scenarios in two cases:
+Case 1: Same Partition $s,t\in G_{i}$, we report $Q(s,t)$ by distance concatenation as shown in the following theorem:
 Theorem 2.
@@
 Proof.
-We denote $d_{L_{i}}(s,t)$ as $d_{2}$, $\min\{d_{L_{i}}(s,b_{i1})+d_{\tilde{L}}(b_{i1},b_{i2})+d_{L_{i}}(b_{i2},t)\}$ as $d_{4}$, and divide all the scenarios into two subcases:
+We denote $d_{L_{i}}(s,t)$ as $d_{2}$, $\min\{d_{L_{i}}(s,b_{i1})+d_{\tilde{L}}(b_{i1},b_{i2})+d_{L_{i}}(b_{i2},t)\}$ as $d_{4}$, and consider two subcases:
 Subcase 1: $p_{G}(s,t)$ does not go outside of $G_{i}$, as shown in the first row of Figure 6.
-No matter $s$ and $t$ are borders or not, $d_{L_{i}}(s,t)$ (i.e., $d_{2}$) is enough to answer $d_{G}(s,t)$ as $L_{i}$ is built based on $G_{i}$ which contains all necessary information for finding the shortest path.
+No matter $s$ and $t$ are borders or not, $d_{L_{i}}(s,t)$ (i.e., $d_{2}$) is enough to answer $d_{G}(s,t)$ as $L_{i}$ is built based on $G_{i}$ which contains all necessary information for finding the shortest path;
 Subcase 2: $p_{G}(s,t)$ passes outside of $G_{i}$, as shown in the second row of Figure 6.
@@
 ∎
-Case 2: Cross-Partitions, i.e., $s\in G_{i},t\in G_{j},i\neq j$, we process queries the same manner as that of Pre-boundary Strategy.
+Case 2: Cross-Partitions, i.e., $s\in G_{i},t\in G_{j},i\neq j$, we process them in the same manner as Pre-Boundary Strategy.
 Lemma 3.
 The cross-partition queries can be correctly processed in the No-Boundary Strategy.
-Note that the proof of Lemma 3 and all the other Lemmas of this paper are available in the Appendix of our extended version (Zhang et al., 2023b).
-We can use $\tilde{L}$ to answer $d_{G}(s,t)$ when $s$ and $t$ are both boundary vertices; otherwise, we have to use $L_{i},L_{j}$ and $\tilde{L}$.
+Proof.
+We prove and discuss it with three subcases as shown in Figure 6-(c).
+Subcase 1: $s,t\in B$, then $d_{G}(s,t)$ can be correctly answered according to Theorem 1;
+Subcase 2: $s\in B$ or $t\in B$.
+As shown in the second illustration in Figure 6-(c), suppose $s$ is an inner vertex of $G_{i}$ and $t$ is a boundary vertex, we take the concise $p_{s,t}$ by extracting the boundary vertices as $p_{c}=\left<s,b_{0},\dots,b_{n},t\right>$.
+Then $b_{0}\in B_{i}$ and $p_{s,t}$ can be treated as concatenated by two sub-paths $p_{s,b_{0}}\oplus p_{b_{0},t}$.
+Specifically, $d_{G}(s,b_{0})=d_{L_{i}}(s,b_{0})$ by referring to the Subcase 1 of Case 1, and $d_{G}(b_{0},t)=d_{\tilde{L}}(b_{0},t)$ by referring to Theorem 1;
+Subcase 3: $s\notin B,t\notin B$, which is the extended case of Subcase 2, so $d_{G}(s,t)$ can also be correctly answered. ∎
+For cross-partition queries, we can use $\tilde{L}$ to answer $d_{G}(s,t)$ when $s$ and $t$ are both boundary vertices; otherwise, we have to use $L_{i},L_{j}$ and $\tilde{L}$.
 The intra-query complexity is $max\{O_{q}(L_{i}),B_{i}\times B_{j}\times O_{q}(\tilde{L}),B_{j}\times O_{q}(L_{j})\}$, while the inter-query takes $O_{q}(\tilde{L})$ time when $s$ and $t$ are both boundary vertices, and the complexity of other case is the worst of them: $max\{B_{i}\times O_{q}(L_{i}),B_{i}\times B_{j}\times O_{q}(\tilde{L}),B_{j}\times O_{q}(L_{j})\}$.
@@
 We repair this weakness by fixing the incorrect boundary all-pair distance in each partition $G_{i}$ and then transforming $\{L_{i}\}$ to their correct version.
-Specifically, we calculate the boundary all-pair distance $d(b_{i1},b_{i2})$ within each partition $G_{i}$ by leveraging $\tilde{L}$ and insert its corresponding edge $e(b_{i1},b_{i2})$ into $G_{i}$ to get $G^{\prime}_{i}$.
+Specifically, we calculate the boundary all-pair distance $d(b_{i1},b_{i2})$ within each partition $G_{i}$ by leveraging $\tilde{L}$ and insert its corresponding edge $e(b_{i1},b_{i2})$ into $G_{i}$ to get $G^{\prime}_{i}$ (Step 4).
 For instance, $d(v_{3},v_{10})$, $d(v_{8},v_{9})$, $d(v_{5},v_{12})$ and $d(v_{6},v_{11})$ are calculated based on $\tilde{L}$ and these edges (as colored red in Figure 5) are inserted into their corresponding partitions with new subgraphs $\{G^{\prime}_{i}\}$ formed.
-Followed by, we treat those newly inserted edges as graph updates and refresh $\{L_{i}\}$ to $\{L^{\prime}_{i}\}$ by invoking the index maintenance algorithms (Ouyang et al., 2020; Zhang et al., 2021c; Zhang and Yu, 2022; Farhan et al., 2022) in each partition parallelly.
+Followed by, we treat those newly inserted edges as graph updates and refresh $\{L_{i}\}$ to $\{L^{\prime}_{i}\}$ by invoking the index maintenance algorithms (Ouyang et al., 2020; Zhang et al., 2021c; Zhang and Yu, 2022; Farhan et al., 2022; Zhang et al., 2021b; Zhou et al., 2024b) in each partition parallelly (Step 5).
 As these procedures are implemented after the No-boundary, we call this strategy as Post-boundary.
@@
 5.3. Index Update for Different PSP Strategies
-The partitioned shortest path has been widely used, however, how to maintain the index to support its application in dynamic environments has never been discussed.
-In this section, we propose the partitioned index update algorithms to guarantee the correctness of the partitioned index in dynamic scenarios.
+The PSP algorithms have been widely used on static networks, however, how to maintain the PSP index to support its application in dynamic environments has never been discussed.
+In this section, we propose universal index update procedures for different PSP strategies to guarantee the correctness of the partitioned index in dynamic scenarios.
 Index Update for Pre-boundary Strategy.
@@
 Then we need to update the corresponding partition index $L_{i}$ and overlay index $\tilde{L}$.
+Step 1 takes $O(|B|\cdot(n\log n+m))$ time, while the partition and $\tilde{G}$ index takes $max\{O_{u}(G_{i}),O_{u}(\tilde{G})\}$ time to update in parallel, where $O_{u}$ is the update complexity for different indexes.
 Lemma 4.
-The indexes of Pre-Boundary Strategy can be correctly updated with the above strategy.
-Step 1 takes $O_{u}(nlogn+m)$ time, while the partition and $\tilde{G}$ index takes $max\{O_{u}(G_{i}),O_{u}(\tilde{G})\}$ time to update in parallel, where $O_{u}$ is the update complexity for different indexes.
+The update procedures of Pre-Boundary Strategy can be correctly deal with any edge weight updates.
+Proof.
+First of all, we need to recalculate Step 1 to identify the affected edges $e(b_{i1},b_{i2})$ between boundary vertex pairs in each partition.
+Even though this step is time-consuming, it cannot be skipped since it would be hard to identify the affected edges.
+For example, suppose the shortest path between the boundary pair $(b_{j1},b_{j2})$ in $G_{j}$ passes through an edge $e\in G_{i}$ with $d_{\tilde{G}}(b_{j1},b_{j2})=d_{0}$.
+When $e$ increases, we could update $L_{i}$ and then $\tilde{L}$.
+But $\tilde{L}$ cannot be correctly updated since it could be that $d_{\tilde{G}}(b_{j1},b_{j2})>d_{0}$, such that $d_{\tilde{G}}(b_{j1},b_{j2})$ cannot be refreshed to the correct value.
+It is because $d_{0}$ contains the old smaller edge weight while cannot be identified, since the shortest distance index always takes the smallest distance value.
+Then we could select those affected edges $e(b_{i1},b_{i2})$ by comparing their old and new weights.
+Lastly, we update their corresponding partition index $L_{i}$ and $\tilde{L}$ in parallel.
+∎
 Figure 7. Index Update of Different PSP Strategies
@@
 If there is an edge weight update, we need to further update $\tilde{L}$.
-The update complexity is $max\{O_{u}(G_{i})\}+O_{u}(\tilde{G})$.
+The update complexity is $max\{O_{u}(G_{i})\}+max\{|B_{i}|^{2}\cdot O_{q}(L_{i})\}+O_{u}(\tilde{G})$.
 Lemma 5.
-The indexes of No-Boundary can be correctly maintained with the above update strategy.
+The update procedures of No-Boundary Strategy can be correctly deal with any edge weight updates.
+Proof.
+In the inter-edge update case, since $e\in\tilde{G},e\notin G_{i},\forall e\in E_{inter}$, the weight change of $e$ could only affect the correctness of $\tilde{L}$.
+So only $\tilde{L}$ should be checked and updated.
+In the Intra-edge update case, since $e\in G_{i}$, its weight change will firstly affect $L_{i}$.
+Then it could affect $\tilde{G}$ as $e_{\tilde{G}}(b_{il_{1}},b_{il_{2}})=d_{L_{i}}(b_{il_{1}},b_{il_{2}})$.
+So $\tilde{L}$ also needs update if $e_{\tilde{G}}(b_{il_{1}},b_{il_{2}})$ changes after checked.
+∎
 Index Update for Post-Boundary Strategy.
-It is similar to No-Boundary, with an additional judgment and processing as shown in Figure 6-(b).
+It is similar to No-Boundary with an additional judgment and processing shown in Figure 7-(b).
 Scenario 1: Intra-edge weight change.
 Suppose $e\in E_{i}$ changes, we update $G_{i},L_{i}$ and then update $\tilde{G},\tilde{L}$ if any $d_{L_{j}}(b_{j1},b_{j2})$ changes.
-After that we may need to further update $\{G^{\prime}_{i}\},\{L^{\prime}_{i}\}$ if $d_{G_{i}^{\prime}}(d_{i1},d_{i2})$ and $d_{\tilde{L}}(d_{i1},d_{i2})$ are different; scenario 2: Inter-edge weight change.
+Then we update $\{G^{\prime}_{i}\},\{L^{\prime}_{i}\}$ if $d_{G_{i}^{\prime}}(d_{i1},d_{i2})$ and $d_{\tilde{L}}(d_{i1},d_{i2})$ are different; Scenario 2: Inter-edge weight change.
 Suppose $e\in E_{inter}$ changes, we update $\tilde{G},\tilde{L}$ and then update $\{G_{i}^{\prime}\},\{L_{i}^{\prime}\}$.
-Therefore, the update time complexity is the linear combination of the above procedures: $max\{O_{u}(G_{i})\}+O_{u}(\tilde{G})+O_{q}(\tilde{G})+max\{O_{u}(G_{i})\}$.
+Therefore, the update time complexity is $max\{O_{u}(G_{i})\}+max\{|B_{i}|^{2}\cdot O_{q}(L_{i})\}+O_{u}(\tilde{G})+max\{|B_{i}|^{2}\cdot O_{q}(\tilde{L})\}+max\{O_{u}(G_{i})\}$.
 Lemma 6.
-The indexes of Post-Boundary Strategy can be correctly maintained with the above update strategy.
-Remark 1.
-To construct the PSP index, Pre-Boundary precomputes the all-pair shortest distance among boundary vertices of each partition, and thus the newly added edges of $G_{i}$ and $\tilde{G}$ (red edges) are of correct edge weight.
-For example, the edge weight of $e(v_{5},v_{12})$ is 6 which is the path length of the shortest path $p_{G}(v_{5},v_{12})=\left<v_{5},v_{3},v_{10},v_{12}\right>$.
-While No-Boundary skips the heavy precomputation and only leverage $L_{i}$ to calculate boundary all-pair distance for overlay graph $\tilde{G}$ construction.
-As a result, the edge weight $w_{\tilde{G}}(v_{5},v_{12})=7$ is incorrect since it only preserves the local distance with corresponding lcoal shortest path $p(v_{5},v_{12})=\left<v_{5},v_{4},v_{12}\right>$ in $G_{3}$.
-This innovation largely saves the index construction time while drags down the query answering speed.
-Then Post-Boundary makes up the query efficiency loss by inserting the correct shortest distance value between boundary pairs into each partition (e.g., insert $w(v_{5},v_{12})=6$ to $G^{\prime}_{3}$) and refresh $\{L_{i}\}$ to its correct version $\{L^{\prime}_{i}\}$.
-In summary, there exists a balance of index performance among the three boundary strategies.
-In terms of the query processing, the No-Boundary is slower than Pre-Boundary and Post-Boundary, since the boundary vertices and path concatenation should be considered for OD within one partition.
-In terms of the index construction and update, the No-Boundary and Post-Boundary enjoy faster speed as the boundary all-pair distance computation can be neglected.
-In terms of the index size, the Post-Boundary needs twice the storage space as that of No-Boundary and Pre-Boundary, even though it has advantages in both query processing and index update.
-Therefore, no one strategy is better than the others in all aspects and could be selected based on the application scenario.
-Besides, it is worth noting that we illustrate PSP strategies in the form of edge-cut for clear presentation since the vertex-cut partitions can be easily converted to edge-cut partitions.
-Due to limited space, the conversion method is presented in extended version (Zhang et al., 2023b).
+The update procedures of Post-Boundary Strategy can be correctly deal with any edge weight updates.
+Proof.
+Firstly, we prove the necessity to keep both $\{G_{i}\},\{L_{i}\}$ and $\{G_{i}^{\prime}\},\{L_{i}^{\prime}\}$.
+Similarly to the Pre-Boundary Strategy, those boundary edges in $\{G_{i}^{\prime}\}$ would keep the old smaller value such that the index could not be correctly updated as explained in Lemma 4.
+So keeping $\{G_{i}\},\{L_{i}\}$ gives us a chance to update $\tilde{L},\{L_{i}\}$ correctly as proved in Theorem 2.
+Then, following the No-Boundary Strategy update, we recompute the shortest distance between the all-pair boundaries leveraging $\tilde{L}$ and compare their values on $L_{i}^{\prime}$, then update $G_{i}^{\prime},L_{i}^{\prime}$.
+∎
+Extension for Structural Updates.
+We next discuss how to support structural updates (edge/vertex insertion/deletion).
+For edge deletion, it is equivalent to increasing the edge weight to +$\infty$+ ∞.
+Vertex deletion can be regarded as deleting all its adjacent edges  (Ouyang et al., 2020; Zhang et al., 2021c, b).
+Regarding the edge insertion, it is equivalent to decreasing an edge weight from +$\infty$+ ∞ to a certain value (Ouyang et al., 2020; Zhang et al., 2021c, b) and thus we can utilize the edge weight decrease update.
+In particular, there are three cases for a newly inserted edge $e(u,v)$: 1) $u$ and $v\in G_{i}$ belong to the same partition; 2) $u,v$ belong to different partitions but they are both boundaries; 3) $u,v$ belong to different partitions and at least one of them is not boundary vertex.
+For case 1), we insert $e$ to $G_{i}$ and leverage edge decrease update to maintain $L_{i}$ (and $\tilde{L}$).
+For case 2), we update the overlay graph $\tilde{G}$ (and the partition index $L_{i}$ if the endpoints of the new edge belong to the same partition $G_{i}$) and overlay index $\tilde{L}$ by edge weight decrease update.
+For case 3), supposing $u\notin B$, this case results in a new boundary vertex $u$ and it affects the partition result, so index reconstruction is needed.
+The vertex insertion is equivalent to a set of edge insertions (Ouyang et al., 2020; Zhang et al., 2021c, b).
+Extension for Batch Updates.
+Our techniques also support batch updates.
+Since CH and TD initially supports the batch updates, we directly leverage the corresponding techniques (Ouyang et al., 2020; Zhang et al., 2021c) for the overlay and partition index maintenance.
+While for PLL that only supports streaming updates, a direct approach to process the batch weight update is to leverage the techniques designed for the streaming update.
+Nevertheless, the all-pair distances computation among boundary vertices could be shared among one batch update for faster index maintenance.
+Table 2. Time Complexity of Different PSP Strategies
+Opeartions Procedure Complexity
+Pre- Boundary Construction $\{B_{i}\times B_{i}\}\rightarrow\{L_{i}\},\tilde{L}$ $O(|B|\cdot(n\log n+m))+max\{O_{c}(G_{i})+O_{c}(\tilde{G})\}$
+Query Intra $L_{i}$ $O_{q}(L_{i})$
+Inter $L_{i}\oplus\tilde{L}\oplus L_{j}$ $max\{|B_{i}|\cdot O_{q}(L_{i}),|B_{i}|\cdot|B_{j}|\cdot O_{q}(\tilde{L}),|B_{j}|\cdot O_{q}(L_{j})\}$
+Update $\{B_{i}\times B_{i}\}\rightarrow\{L_{i}\},L$ $O(|B|\cdot(n\log n+m))+max\{O_{u}(G_{i})+O_{u}(\tilde{G})\}$
+No- Boundary Construction $\{L_{i}\}\rightarrow\tilde{G}\rightarrow\tilde{L}$ $max\{O_{c}(G_{i})\}+max\{|B_{i}|^{2}\cdot O_{q}(L_{i})\}+O_{c}(\tilde{G})$
+Query Intra $L_{i}\oplus\tilde{L}\oplus L_{i}$ $max\{|B_{i}|\cdot O_{q}(L_{i}),|B_{i}|^{2}\cdot O_{q}(\tilde{L}))\}$
+Inter $L_{i}\oplus\tilde{L}\oplus L_{j}$ $max\{|B_{i}|\cdot O_{q}(L_{i}),|B_{i}|\cdot|B_{j}|\cdot O_{q}(\tilde{L}),|B_{j}|\cdot O_{q}(L_{j})\}$
+Update $\{L_{i}\}\rightarrow\tilde{G}\rightarrow\tilde{L}$ $max\{O_{u}(G_{i})\}+max\{|B_{i}|^{2}\cdot O_{q}(L_{i})\}+O_{u}(\tilde{G})$
+Post- Boundary Construction $\{L_{i}\}\rightarrow\tilde{G}\rightarrow\tilde{L}$$\rightarrow\{B_{i}\times B_{i}\}\rightarrow\{L_{i}\}$ $max\{O_{c}(G_{i})\}+max\{|B_{i}|^{2}\cdot O_{q}(L_{i})\}+O_{c}(\tilde{G})$ $+max\{|B_{i}|^{2}\cdot O_{q}(\tilde{G})\}+max\{O_{u}(G_{i})\}$
+Query Intra $L_{i}$ $O_{q}(L_{i})$
+Inter $L_{i}\oplus\tilde{L}\oplus L_{j}$ $max\{|B_{i}|\cdot O_{q}(L_{i}),|B_{i}|\cdot|B_{j}|\cdot O_{q}(\tilde{L}),|B_{j}|\cdot O_{q}(L_{j})\}$
+Update $\{L_{i}\}\rightarrow\tilde{G}\rightarrow\tilde{L}$$\rightarrow\{B_{i}\times B_{i}\}\rightarrow\{L_{i}\}$ $max\{O_{u}(G_{i})\}+max\{|B_{i}|^{2}\cdot O_{q}(L_{i})\}+O_{u}(\tilde{G})$ $+max\{|B_{i}|^{2}\cdot O_{q}(\tilde{G})\}+max\{O_{u}(G_{i})\}$
+Because different SP indexes have different complexities but share the same logical procedure, we use the following notations to represent the logical complexities: $O_{c}$ is the index construction complexity, $O_{q}$ is the query complexity, and $O_{u}$ is the index update complexity.
+$max$ returns the maximum processing time among all partitions.
+We analyze and summarize the time complexity of PSP strategies in terms of index construction, query processing, and index update in Table 2.
+Notably, the boundary vertex numbers $|B|$ and $|B_{i}|$ are two crucial factors of the index performance.
+The advantage of the no-boundary and post-boundary strategies over pre-boundary strategy source from the replacement of the time-consuming Dijkstra’s searches ($O(|B|\cdot(m+n\log n)\})$) with index-based all-pair boundary query processing ($O(max\{|B_{i}|^{2}\cdot O_{q}(L_{i})\})$).
+Therefore, such an optimization generally results in much better performance if the underlying partition index is 2-hop labeling (in which case $O_{q}(L_{i})=w(G_{i})$, where $w(G_{i})$ is the treewidth of $G_{i}$).
+However, if the query processing is conducted in a search-based manner (e.g., direct search or CH) and the boundary number $|B_{i}|$ is very large, the no-boundary strategy’s performance may dramatically deteriorate.
+Such a deterioration is even worse for the post-boundary strategy.
+Nevertheless, no-boundary and post-boundary strategies generally have much better performance than the pre-boundary strategy in most cases.
 5.4. Pruning-based Overlay Graph Optimization
@@
 Then is it possible to improve index performance and decrease the density of $\tilde{G}$ by deleting some unnecessary edges in $\{B_{i}\times B_{i}\}$?
-As proved in Theorem 1, the shortest distance between boundary vertices is well-preserved in $\tilde{G}$, then could the distance still be preserved in $\tilde{G}$ if some edges are removed?
-We start exploring this question by categorizing the boundary vertices as followed:
+We explore this question by categorizing the boundary vertices as follows:
 Definition 0 (Half / Full-Connected Boundary Vertex).
@@
 Figure 8. Example of Overlay Graph Simplification
-As illustrated in Figure 7-(a), $b_{i3}$ is a half-connected boundary vertex since there exists $v_{i1}\in N(b_{i3}),v_{i1}\notin B_{i}$; and $b_{i1},b_{i2},b_{i4}$ are full-connected boundary vertices.
-Initially, $\tilde{G}$ can be viewed as composed of multiple complete graphs $\{B_{i}\times B_{i}\}$, which are connected by the inter-edges.
-Within this dense structure, the global distance between any two boundary vertices $s,t\in B$ is well-preserved.
-Let us consider a shortest path $p$ with starting vertex $b_{s}\in B_{i}^{F}$ and target vertex $b_{t}\in B$, then the second vertex $v_{0}$ along $p$ is a boundary vertex with $v_{0}\in B,v_{0}\in N(b_{s})$ according to the definition above.
-Therefore, the global distance information of $p$ will still be kept in $\tilde{G}$ if we delete edges $\{(b_{s},b_{i})\}$ with $b_{i}\in B_{i},b_{i}\notin N(b_{s})$, since it is impossible for $p$ to pass through these edges.
-The same principle holds for a shortest path ending with a fully connected boundary vertex $b_{t}\in B_{i}^{F}$ that its distance will also be preserved in $\tilde{G}$ after deleting edges $\{(b_{i},b_{t})\}$ with $b_{i}\in B_{i},b_{i}\notin N(b_{t})$.
-Then for the shortest path starting and ending with the half-connected boundary vertices $b_{s},b_{t}\in B_{i}^{H}$, it is necessary to insert an edge $(b_{s},b_{t})$ into $\tilde{G}$ to preserve its distance information, as the second vertex on this path could be a non-boundary vertex which is not contained in the overlay graph.
+As illustrated in Figure 8-(a), suppose partition $G_{i}$ has five boundary vertices, where $b_{1}$, $b_{2}$, and $b_{4}$ are full-connected boundary vertices since their neighbors are all boundary vertices while $b_{3}$ and $b_{5}$ are half-connected boundary vertex since they have neighbors that are non-boundary vertices ($v_{1}$ and $v_{2}$).
+We observe that the shortest paths from full-connected boundary vertices to other boundary vertices always pass through at least one of their boundary neighbors.
+Let us consider the shortest path $p$ from $b_{4}\in B_{i}^{F}$ to any other boundary vertex $b\in B$, then the second vertex $v$ along $p$ must be a boundary vertex with $v\in N(b_{4})$.
+Therefore, the global distance information of $p$ will still be kept in $\tilde{G}$ if we delete edges $\{(b_{4},b_{i})\}$ with $b_{i}\in B_{i},b_{i}\notin N(b_{4})$, since it is impossible for $p$ to pass through these edges.
+By contrast, for the shortest path starting and ending with the half-connected boundary vertices $b_{s},b_{t}\in B_{i}^{H}$, it is necessary to insert an edge $(b_{s},b_{t})$ into $\tilde{G}$ to preserve its distance information, as the second vertex on this path could be a non-boundary vertex which is not contained in the overlay graph.
 As analyzed above, we can shrink $\tilde{G}$ ($\{B_{i}\times B_{i}\}\cup E_{inter}$) to its slimmer version $\tilde{G}^{\prime}$ (composed of $\{B_{i}^{H}\times B_{i}^{H}\}$, $E_{inter}$ and $B_{i}^{F}$’s adjacent edges).
-Suppose $b_{i1},b_{i2}\in N(b_{i4})$ and let us only take $G_{i}$ as an example for the clear presentation, the original overlay graph as shown in Figure 7-(b) will be shrunk to a sparser one in Figure 7-(c) after the pruning technique.
-The following lemma demonstrates that the global distance can still be preserved in the overlay graph with the pruning technique.
+Let us only take $G_{i}$ as an example for the clear presentation, the original overlay graph as shown in Figure 8-(b) will be shrunk to a sparser one in Figure 8-(c) after the pruning technique, reducing 6 redundant shortcuts.
+In the following Lemma, we prove the global distance can still be preserved in the overlay graph after pruning.
 Lemma 8.
-$d_{G}(b_{s},b_{t})=d_{\tilde{G^{\prime}}}(b_{s},b_{t})$, $\forall(b_{s},b_{t})\in B\times B$.
-Therefore, we can enhance the index performance by pruning some edges in $\tilde{G}$ and utilizing its sparser version $\tilde{G}^{\prime}$, while preserving the global distance by referring to Lemma 8.
-4. SP-Oriented Partition Method Classification
-Having figured out how to achieve efficient query and update in partitioned graphs while guaranteeing the index correctness with the novel PSP strategies, we move on to the remaining two dimensions.
-As the SP index is well studied and analyzed in Section 2.4.
-Then, only the mist on the partition method remains: how to choose a partition method to achieve our preferred performance?
-This is not trivial because: 1) dozens of partition methods with different characteristics were proposed and applied in the past decades; 2) they are classified under different criteria (Buluç et al., 2016), for example, from the perspectives of partition manners (spectral, flow, graph glowing, contraction and multi-level), partition objectives (balance and minimal cut), computation manner (in-memory, distributed and streaming), and cut category (edge-cut and vertex-cut), but it is not clear which criterion is beneficial to SP index; 3) their relationship with (or the influence on) SP index is unknown and has never been studied.
-Based on these doubts, we propose a novel SP-oriented classification of partition methods as follows:
-1) Planar Partition treats partitions equally on one level.
-It decomposes a graph $G$ into multiple equally-important subgraphs $\{G_{i}|1\leq i\leq k\}(k\in[2,\infty])$ with 1) $\bigcup_{i\in[1,k]}V(G_{i})=V$, $V(G_{i})\cap V(G_{j})=\emptyset$, or 2) $\bigcup_{i\in[1,k]}E(G_{i})=E$, $E(G_{i})\cap E(G_{j})=\emptyset,\forall i,j\in[1,k],i\neq j$.
-Representative methods include spectral partitioning (Pothen et al., 1990; Barnard and Simon, 1994), graph growing (Diekmann et al., 2000), flow-based partitioning (Sanders and Schulz, 2011; Delling et al., 2011), node-swapping (Kernighan and Lin, 1970; Fiduccia and Mattheyses, 1982), multilevel graph partition (MGP) (Karypis and Kumar, 1998; Barat et al., 2018), etc..
-2) Core-Periphery Partition ((Borgatti and Everett, 2000; Rombach et al., 2014; Elliott et al., 2020; Maehara et al., 2014a)) treats the partitions discriminately by taking some important vertices as “Core” and the remaining ones as “Peripheries”.
-It decomposes a graph $G$ into two distinct parts: a core subgraph $G_{c}$ and $k-1$ peripheral subgraphs $\{G_{i}|1\leq i\leq k-1\}(k\in[2,\infty])$, satisfying $\bigcup_{i\in[1,k-1]}V(G_{i})\cup V(G_{c})=V,V(G_{c})\cap V(G_{i})=\emptyset(\forall 1\leq i\leq k-1)$ and $V(G_{i})\cap V(G_{j})=\emptyset(\forall i,j\in[1,k-1],i\neq j)$.
-There are two big streams depending on how the core is formed: Core-Tree Decomposition (Maehara et al., 2014b; Li et al., 2020a; Zheng et al., 2022) forms the core through tree decomposition (Robertson and Seymour, 1984), and the resulting periphery part is a set of small-width trees.
-Specifically, it leverages minimum degree elimination (Xu et al., 2005) to contract vertices of lower degree first such that the graph is contracted from the edge towards the center, generating a set of growing trees (“Peripheries”) around a shrinking while denser graph (“Core”) formed by those non-contracted vertices.
-The contraction terminates when the width of one tree reaches the previously set threshold; Sketch (Potamias et al., 2009; Das Sarma et al., 2010; Gubichev et al., 2010; Tretyakov et al., 2011; Qiao et al., 2012; Qi et al., 2013; Farhan et al., 2018; Wang et al., 2021; Farhan et al., 2022) selects a set of vertices as landmarks, which can be regarded as the core and treats the remaining parts as periphery.
-It generally works on extremely large graphs where the above tree decomposition is impossible.
-3) Hierarchical Partition organizes the network partitions hierarchically and each level is equivalent to planar partitioning.
-It organizes a graph $G$ hierarchically within $H$ levels where each level $h$ is a planar partitioning $\{G^{h}_{i}\}$ of $G$ and $\exists G^{h-1}_{j}\supset G^{h}_{i},\forall G^{h}_{i}(1\leq h<H)$.
-Classic methods include HiTi (Jung and Pramanik, 1996), SHARC (Bauer and Delling, 2010), G-Tree (Zhong et al., 2015), G*-Tree (Li et al., 2019a), SMOG (Holzer et al., 2009), and so on.A typical approach to create hierarchical partitions is by utilizing MGP, such as METIS (Karypis and Kumar, 1998) in G-Tree and G*-Tree.
-Figure 8. Different SP-Oriented Partitioning
-Table 1. SP-Oriented Partition Methods Classification.
-Partition Structure Partition Method Cut Category Partition Objectives Spatial-Aware
-Category Representative Methods Edge Cut Vertex Cut Balance Minimal Cut
-Planar Spectral Partitioning SP (Pothen et al., 1990), RSB (Barnard and Simon, 1994) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-Graph Growing Bubble (Diekmann et al., 2000) $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
-Flow-Based Partitioning PUNCH (Delling et al., 2011) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-Geometric Partitioning RCB (Simon, 1991; Huang et al., 1996) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-Node-Swapping Heuristic KL (Kernighan and Lin, 1970), FM (Fiduccia and Mattheyses, 1982) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-Minimum $k$-cut $k$-cut (Goldschmidt and Hochbaum, 1994) $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
-Multilevel Graph Partitioning SCOTCH (Pellegrini and Roman, 1996), METIS (Karypis and Kumar, 1998), KaHyPar (Akhremtsev et al., 2017) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-Streaming Edge Partitioning HDRF (Petroni et al., 2015; Mayer and Jacobsen, 2021), CLUGP (Kong et al., 2022) $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
-In-Memory Edge Partitioning NE (Zhang et al., 2017), HEP (Mayer and Jacobsen, 2021) $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
-Core-Periphery Core-Tree Decomposition Core-Tree (Maehara et al., 2014b; Li et al., 2020a; Zheng et al., 2022; Elliott et al., 2020) $\checkmark$ $\times$ $\times$ $\times$ $\times$
-Sketch Sketch (Potamias et al., 2009; Das Sarma et al., 2010; Gubichev et al., 2010; Tretyakov et al., 2011; Qiao et al., 2012; Qi et al., 2013; Farhan et al., 2018; Wang et al., 2021; Farhan et al., 2022) $\checkmark$ $\times$ $\times$ $\times$ $\times$
-Hierarchical Edge-Cut Based HP HiTi (Jung and Pramanik, 1996) $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$
-SHARC (Bauer and Delling, 2010), G-Tree (Zhong et al., 2015), G*-Tree (Li et al., 2019a) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-Vertex-Separator Based HP SMOG (Holzer et al., 2009), ROAD  (Lee et al., 2009, 2010) $\times$ $\checkmark$ $\times$ $\times$ $\times$
-Figure 8 shows the partition results of an example graph $G$ with different partitioning methods, where the red vertices represent the corresponding boundary vertices.
-Specifically, planar partitioning generates four partitions with equivalent vertex size while core-periphery partitioning produces both “Core” and “Periphery” (the periphery of Core-Tree decomposition is a set of small-width trees while Sketch treats the non-core part as periphery.
-The yellow vertex in each periphery is the “root vertex”).
-Hierarchical partitioning (HP) organizes partitions hierarchically, and the leaf nodes (the lowest level partition result) may have the same partition result as planar partitioning if the same partitioning method is used.
-In addition, we summarize and categorize existing partition methods under our SP-oriented classification, as shown in Table 1.
-Remark 2.
-As analyzed in Section 3, the boundary number of each partition will affect the partition method section as it is crucial to the computation complexity of PSP index.
-Specifically, the planar and hierarchical partitions have no limit to the boundary number, even though reducing the cut size is one of their optimization goals as the path index may suffer from large boundary numbers.
-Complementary, the core-tree decomposition limits the boundary number to be no larger than the pre-defined bandwidth such that their performance would not be deteriorated by the boundary if we set the bandwidth wisely.
-Therefore, the planar and hierarchical partitions are better used for small treewidth networks such that balanced partitions and fewer boundaries can be achieved together; while core-periphery is suitable for large treewidth networks as it deliberately limits the boundary vertex number by bandwidth.
+$\forall b_{s},b_{t}\in B,d_{G}(b_{s},b_{t})=d_{\tilde{G^{\prime}}}(b_{s},b_{t})$.
+Proof.
+We denote the concise form of $p_{G}(b_{s},b_{t})$ by only taking the borders as $\{b_{s}=b_{0},\dots,b_{t}=b_{h}\},(0<j<h)$.
+When $h=1$, if either $b_{s}\in B^{F}$ or $b_{t}\in B^{F}$, then $d_{G}(b_{s},b_{t})=w_{G}(b_{s},b_{t})$ holds with $b_{s}\in N(b_{t})$ by referring to definition 7.
+Since $(b_{s},b_{t})\in\tilde{G}^{\prime}$, $d_{G}(b_{s},b_{t})=d_{\tilde{G^{\prime}}}(b_{s},b_{t})$ holds.
+The same applies when $b_{s},b_{t}\in B^{H}$ with two endpoints in different subgraphs $b_{s}\in B_{i},b_{t}\in B_{j}(i\neq j)$.
+If $b_{s},b_{t}\in B^{H}$ with $b_{s},b_{t}\in B_{i}$, edge $(b_{s},b_{t})$ are inserted into $\tilde{G}^{\prime}$ with their global shortest distance in both Pre-Boundary and Post-Boundary strategies, so $d_{G}(b_{s},b_{t})=d_{\tilde{G}^{\prime}}(b_{s},b_{t})$ naturally holds.
+In the No-Boundary strategy, though only the local shortest distance is inserted in the overlay graph with $d_{G_{i}}(b_{s},b_{t})=d_{\tilde{G}^{\prime}}(b_{s},b_{t})$, it holds that $d_{G}(b_{s},b_{t})=d_{G_{i}}(b_{s},b_{t})$ since there exists no other boundary vertex besides two endpoints in $p_{G}(b_{s},b_{t})$ and it only pass through $G_{i}$, so $d_{G}(b_{s},b_{t})=d_{\tilde{G^{\prime}}}(b_{s},b_{t})$ holds.
+When $h>1$, the shortest path distance is accumulated as $d_{G}(b_{s},b_{t})=\sum_{j=0}^{h-1}d_{G}(b_{j},b_{j+1})$.
+Since $d_{G}(b_{j},b_{j+1})=d_{\tilde{G}^{\prime}}(b_{j},b_{j+1})$ holds for $0\leq j<h$ by referring to the case when $h$=1, we can get that $d_{G}(b_{s},b_{t})=d_{\tilde{G}^{\prime}}(b_{s},b_{t})$.
+So we prove that $d_{G}(b_{s},b_{t})=d_{\tilde{G^{\prime}}}(b_{s},b_{t})$ holds for all scenarios and boundary strategies.
+∎
+Therefore, we enhance the index performance by utilizing the sparser $\tilde{G}^{\prime}$.
+Note that Lemma 8 also holds with edge weight updates since this optimization only depends on the structure of $\tilde{G}$.
 6. Universal PSP Index Scheme and Newly Generated PSP Indexes
-In this section, we propose a universal PSP index scheme and generator and put forward five representative PSP indexes for query-oriented or update-oriented applications based on this scheme.
+In this section, we propose a universal PSP index scheme and then put forward five representative PSP indexes for query-oriented or update-oriented applications based on this scheme.
 Figure 9. PSP Index Generator with PSP Index Scheme and Representative Generated PSP Indexes
-5.1. PSP Index Scheme and Generator
-After elaborating on each component of the PSP index, we now have our PSP index scheme ready as shown in Figure 9.
+6.1. PSP Index Dimensions and Universal Scheme
+Based on the comprehensive literature review and analysis of the PSP index in previous Sections, we identify three critical dimensions of the PSP index as follows.
+1) Shortest Path Algorithm that decides the underlying methods for the overlay graph and partitions.
+The performance of different SP algorithms on dynamic networks has been thoroughly studied in our previous work (Zhang et al., 2021a);
+2) Graph Partition Method, which divides the graph into multiple subgraphs and is the key difference from the standalone SP indexes.
+The existing partition methods are not designed for path indexes, so how they affect the PSP index is unknown and deserves study;
+3) PSP Strategy, which provides a procedure to organize the index construction and update for different subgraphs and coordinate those indexes such that the query correctness is guaranteed.
+This dimension is crucial because it determines the complexity of both the query and update.
+However, only straightforward pre-boundary strategy exists currently.
+Thus we further propose two novel strategies to improve its index construction and maintenance performance.
+Given a dynamic network, these three dimensions provide a roadmap to construct a PSP index by selecting a specific graph partition method, SP algorithm and PSP strategy such that the application requirement (query / update efficiency) can be satisfied.
+Inspired by this, we propose a universal PSP Index Scheme as shown in Figure 9.
 The main part is the PSP index generator, which is formed by coupling those three components.
@@
 Meanwhile, planar and hierarchical partitions are more suitable for low-treewidth networks, while core-periphery can deal with high-treewidth graphs.
-Given the specific application requirements in terms of computation efficiency, space consumption, or partition type, our scheme could provide one PSP index with a specified PSP strategy, SP algorithm, and partition structure through the PSP index generator.
-For starters, the existing PSP indexes can find their positions in our scheme as will be discussed in Section 7, and we can compare them theoretically and fairly.
-Apart from this, we can further utilize it as a PSP Index Generator to create new PSP indexes that are suitable to user requirements, especially when all the existing ones are query-oriented, while there leaves a vast space of combinations for new ones.
+There are two benefits of our PSP index scheme.
+Firstly, the existing PSP indexes can find their positions in our scheme and we can compare them theoretically and fairly.
+For exmaple, 1) Pre-Boundary + Search/All-Pair + Hierarchy/ Planar: HiTi (Jung and Pramanik, 2002), CRP (Delling et al., 2017), PbS (Chondrogiannis and Gamper, 2014), ParDiSP (Chondrogiannis and Gamper, 2016) pre-compute shortest distances between boundaries to guide the search.
+In a broad sense, Arc-flag (Möhring et al., 2007) and SHARC (Bauer and Delling, 2010) also belong to this category.
+G-tree (Zhong et al., 2013; Li et al., 2019a) uses dynamic programming to compute the distance between layer’s all-pair information to replace searching and it is used in kNN (Zhong et al., 2015), ride-sharing (Ta et al., 2017), time-dependent (Wang et al., 2019), and distance embedding (Huang et al., 2021).
+They are slow to construct due to Pre-B and inefficient to query due to direct search; 2) Pre-Boundary+Search/PLL+Planar: COLA (Wang et al., 2016) builds labels for the skyline shortest path on the overlay graph for the constrained shortest path.
+Its construction suffers from Pre-B and query suffers from searching; 3) Pre-Boundary+TD/TD+Planar: FHL (Liu et al., 2021, 2022) builds the TD both within and between partitions for multi-dimensional skyline paths; 4) Pre-Boundary+PLL/PLL+Planar: T2Hop (Li et al., 2019c, 2020b) utilizes two layers of PLL to reduce the complexity of long-range time-dependent paths, and this structure’s performance is limited by PLL; 5) Pre-Boundary+Search/All-Pair/PLL+Sketch: This category works on huge graphs where index is impossible so only a small number of landmarks are selected to prune the search (Das Sarma et al., 2010; Wang et al., 2021; Farhan et al., 2022) or approximate the result (Gubichev et al., 2010; Qiao et al., 2012).
+Secondly, given the specific application requirements in terms of computation efficiency, space consumption, or partition type, our scheme could provide one PSP index with a specified PSP strategy, SP algorithm, and partition structure through the PSP index generator.
+Especially when all the existing ones are query-oriented, there is a vast space of combinations for new ones.
 Specifically, the generated PSP index needs to choose the components from each of the three dimensions, and its name also has three parts corresponding to the three dimensions.
-Although we can enumerate them all, we choose to introduce the representative ones from the perspectives of query- and update-oriented under different partition structures, as shown in Figure 9.
+Although we can enumerate them all, we choose to introduce five representative ones from the perspectives of query- and update-oriented under different partition structures, as shown in Figure 9.
+The details of these five PSP indexes will be elaborated in Section 6.2.
 6.2. Representative Generated New PSP Indexes
@@
 Query-Oriented Planar PSP Index: P-TD-P.
-Given a planar partition, the PSP strategies can apply directly to it.
-Next, we introduce how to construct the PSP indexes that are efficient in query processing under planar partition.
-Firstly in terms of PSP strategy, since both Pre-Boundary and Post-Boundary are efficient in query answering and our proposed Post-Boundary is faster than Pre-Boundary in index construction, we use it as the PSP strategy.
+We introduce how to construct a planar PSP index that is efficient in query processing.
+Firstly, in terms of PSP strategy, since both Pre-Boundary and Post-Boundary are efficient in query answering and our proposed Post-Boundary is generally faster than Pre-Boundary in index maintenance, we use it as the PSP strategy.
 Secondly, in terms of the SP algorithm, we could choose TD as the index for both partitions and overlay graph.
@@
 The core-periphery partition index comprises the core index $L_{c}$ and the periphery index $\{L_{i}\}$.
-It seems that $L_{c}$ and $L_{i}$ can be constructed by their corresponding subgraphs $G_{c}$ and $G_{i}$, respectively.
 Although the core does not belong to any partition, they are connected to the partitions, and we treat the core as the overlay graph.
@@
 As for index update, it takes $O(B_{max}\cdot V_{max}(logV_{max}+E_{max}))$ to update the partition all-pairs and $O(w^{2}_{\tilde{G}}\cdot\delta)$ to update the overlay graph.
-Remark 3.
-We note that these five PSP indexes are novel structures, each with unique technical challenges, especially in updating.
+In summary, these five PSP indexes are novel structures, each with unique technical challenges, especially in updating.
 Nevertheless, we believe the PSP scheme holds greater significance than these new indexes, as it can guide new PSP index designs.
-We omit their details due to limited space, but their source codes are provided (GitHub, [n.d.]).
-Besides, we take the N-CH-P index as an example and elaborate on its details in the Appendix (FullVersion).
-Table 2. Dataset Description
-[] Type Name Dataset $|V|$ $|E|$ Road Network NY 1 New York City 264,346 730,100 FL 1 Florida 1,070,376 2,687,902 W 1 Western USA 6,262,104 15,119,284 US 1 Full USA 23,947,347 57,708,624 Complex Network GO 2 Google 855,802 8,582,704 SK 3 Skitter 1,689,805 21,987,076 WI 3 Wiki-pedia 3,333,272 200,923,676 FR 2 com-Friendster 65,608,366 3,612,134,270 [1] http://www.dis.uniroma1.it/challenge9/download.shtml [2] http://snap.stanford.edu/data [3] http://konect.cc/
-Table 3. Average Vertex Degree of Overlay Graph
-Dataset NY FL GO SK
-Method PUNCH HEP PUNCH HEP KaHyPar HEP KaHyPar HEP
-w/o Opt 48.68 325.98 36.17 719.25 947.36 9754.52 1420.18 34889.5
-w/ Opt 48.68 320.61 36.17 714.30 912.24 7678.05 1409.98 27326.5
+We omit their details due to limited space, but their source codes are provided (GitHub, [n.
+d.]).
+Besides, we take the N-CH-P index as an example and elaborate on its details in the extended version (Zhang et al., [n.
+d.]).
 7. Experimental Evaluation
@@
 All the algorithms are implemented in C++ with full optimization on a server with 4 Xeon Gold 6248 2.6GHz CPUs (total 80 cores / 160 threads) and 1.5TB memory.
-The thread number is set to 150.
+The default thread number is set to 150.
+Table 3. Dataset Description
+[] Category Name Dataset $|V|$ $|E|$ Type Road Network NY 1 New York City 264,346 730,100 road FL 1 Florida 1,070,376 2,687,902 road W 1 Western USA 6,262,104 15,119,284 road US 1 Full USA 23,947,347 57,708,624 road Complex Network GO 2 Google 855,802 8,582,704 web SK 3 Skitter 1,689,805 21,987,076 social WI 3 Wiki-pedia 3,333,272 200,923,676 web FR 2 com-Friendster 65,608,366 3,612,134,270 social [1] http://www.dis.uniroma1.it/challenge9/download.shtml
+[2] http://snap.stanford.edu/data [3] http://konect.cc/
 7.1. Experimental Settings
-Datasets and Queries.
-We test on eight real-life datasets, including four weighted road networks (NY, FL, W, and US) and four complex networks (Web graph: GO and WI; Social network: SK and FR), as shown in Table 2.
-The edge weight of complex networks is generated by following the manner (Zhang et al., 2021b) that is inversely proportional to the highest degree of the endpoints.
-We randomly generate 10,000 queries and 1,000 update instances for each dataset to assess the query processing and index maintenance efficiency, respectively.
+Datasets.
+We test on eight real-life datasets (four road networks and four complex networks as shown in Table 3).
+Following (Zhang et al., 2021b; Zhou et al., 2024b; Zhang et al., 2021a; Zhou et al., 2024a), the edge weights of complex networks are randomly generated inversely proportional to the highest degree of the endpoints.
 Partition Methods and Their Performance Metrics.
-To select the suitable partition method for PSP index, we implement 8 representative partition methods: Bubble (Diekmann et al., 2000), RCB (Simon, 1991), KaHyPar (Akhremtsev et al., 2017), PUNCH (Delling et al., 2011), SCOTCH (Barat et al., 2018), METIS (Karypis and Kumar, 1998), HEP (Mayer and Jacobsen, 2021), CLUGP (Kong et al., 2022), all of which are planar partition methods with PUNCH and RCB only working on road networks.
-Since core-periphery partition (including Tree-Decomposition-based partition (Ouyang et al., 2018) and landmark selection in Sketch (Farhan et al., 2022)) and hierarchical partition (mostly using METIS (Karypis and Kumar, 1998)) methods are finite, they are selected by default in the corresponding PSP index.
+To select the suitable partition method for PSP index, we implement 8 representative methods: Bubble (Diekmann et al., 2000), RCB (Simon, 1991), KaHyPar (Akhremtsev et al., 2017), PUNCH (Delling et al., 2011b), SCOTCH (Barat et al., 2018), METIS (Karypis and Kumar, 1998), HEP (Mayer and Jacobsen, 2021), CLUGP (Kong et al., 2022), all of which are planar with PUNCH and RCB only working on road networks.
+Since core-periphery (core-tree (Zhou et al., 2024b) and Sketch (Farhan et al., 2022)) and hierarchical partition (uses METIS (Karypis and Kumar, 1998)) methods are finite, they are selected by default in the corresponding PSP index.
 We evaluate the performance of a partition method from three aspects: overall border number $|B|=\sum_{i\in[1,k]}|B_{i}|$, average partition border number $\overline{|B_{i}|}=\frac{\sum_{i\in[1,k]}|B_{i}|}{k}$, and partition connectivity $R_{C}=\frac{\sum_{i\in[1,k]}|C(G_{i})|}{k}$, where $C(G_{i})$ is the connected component number of subgraph $G_{i}$.
-PSP Indexes and Their Performance Metrics In terms of PSP indexes comparison, we evaluate 5 existing representative PSP indexes spreading over our proposed three partition structures: a) FHL (Liu et al., 2021, 2022a) and THop (Li et al., 2019c) belongs to planar partition; b) Core-Tree (CT) (Li et al., 2020a; Zheng et al., 2022) and Sketch (Farhan et al., 2022) belong to the core-periphery partition; c) G-Tree (Zhong et al., 2013) belongs to the hierarchical partition.
-Besides, we implement our 5 newly proposed PSP indexes (see Section 5): P-TD-P, N-CH-P, P-PT-CP, N-PC-CP, and N-TS-HP.
-For the state-of-the-art shortest path algorithms CH (Geisberger et al., 2008; Ouyang et al., 2020), TD (Ouyang et al., 2018; Zhang et al., 2021c), and PLL (Akiba et al., 2013; Zhang et al., 2021b), we implement their partitioned version (named PCH, PTD, and PPLL respectively) to test the performance of different PSP strategies and select the suitable partition number.
-We measure the performance of the PSP index from four aspects: index construction time $t_{c}$, query time $t_{q}$, update efficiency $t_{u}$, and index size $s$.
-Note that we do not report the result (or set it to $\infty$) if the running time of a partition method or PSP index exceeds 24 hours.
+PSP Indexes and Performance Metrics.
+We evaluate 5 SOTA PSP indexes combining with the three proposed partition structures: a) FHL (Liu et al., 2021, 2022) and THop (Li et al., 2019c) from planar partition; b) Core-Tree (CT) (Li et al., 2020a; Zheng et al., 2022; Zhou et al., 2024b) and Sketch (Farhan et al., 2022) from core-periphery; c) G-Tree (Zhong et al., 2013) from hierarchical partition.
+Besides, we implement our 5 newly proposed PSP indexes: P-TD-P, N-CH-P, P-PT-CP, N-PC-CP, and N-TS-HP.
+For the SOTA shortest path indexes CH (Geisberger et al., 2008; Ouyang et al., 2020), TD (Ouyang et al., 2018; Zhang et al., 2021c), and PLL (Akiba et al., 2013; Zhang et al., 2021b), we implement their partitioned version (named PCH, PTD, and PPLL).
+We measure the performance from four aspects: index construction time $t_{c}$, query time $t_{q}$, update time $t_{u}$, and index size $s$.
+We do not report the result ($\infty$) if index time exceeds 24 hours.
+Query and Update.
+For each dataset, we randomly generate a set of 10,000 queries and 1,000 edge update instances to evaluate $t_{q}$ and $t_{u}$.
+Following (Zhang et al., 2021c, a), for each selected edge $e$, we decrease or increase its weight to $\alpha\cdot|e|$, where $\alpha$ randomly ranges in $(0,2]$.
+Besides, we also randomly generate 1000 update instances for each type of structural update (edge/vertex insertion/deletion) for Exp 5.
 Parameter Setting.
-We set the default partition number as $32$ for all planar partitions and the bandwidth of core-tree indexes as $40$ as per experimental results.
-The landmark number of Sketch is set to $20$ by referring to (Farhan et al., 2022).
-As for the hierarchical partition, we follow G-tree (Zhong et al., 2013; Li et al., 2019a) and set the fan-out $f$ as $4$ and the maximum leaf node size $\tau$ as 128 in NY, 256 in FL, 512 in W and US, respectively.
-Figure 10. Partition Performance When Varying $k$.
+We set the default partition number as $32$ for all planar PSP and the bandwidth of CT-based PSP as $40$ as per our preliminary experimental results in (Zhang et al., [n.
+d.]).
+The landmark number of Sketch is $20$ (Farhan et al., 2022).
+For the hierarchical PSP, we follow G-tree (Zhong et al., 2013; Li et al., 2019a) and set the fan-out $f$ as $4$ and the leaf node size $\tau$ as 128 in NY, 256 in FL, 512 in W and US.
+We use No-boundary as the default strategy if the PSP strategy is not specified.
+Figure 10. Performance of Partition Methods When Varying Partition Number $k$.
 Figure 11. Effect of Partition Methods. Bar: Left, Ball: Right
-Figure 12. Effect of Partition Number. Bar: Left, Ball: Right
-Figure 13. Comparison with Non-partitioned SP Indexes. Bar: Left, Ball: Right
-Figure 14. Effectiveness of Partitioned Shortest Path Strategy. (Bar: Left Axis; Ball: Right Axis)
 7.2. Experiment Result Analysis
 Exp 1: Performance of Partition Methods When Varying $k$.
-We first report the performance of different partition methods with varying partition number $k$ (we only report results on FL, US, and GO due to limited space; the full results are in the full version (Zhang et al., 2023b))).
-As shown in Figure 10 (a)-(f), the overall border number $|B|$ generally increases, while the average border number $\overline{|B_{i}|}$ decreases as $k$ grows from 4 to 256.
-This indicates that having a larger partition number can potentially speed up the query processing of the PSP index due to a smaller average partition border number.
+As shown in Figure 10 (a)-(p), the overall border number $|B|$ generally increases while the average border number $\overline{|B_{i}|}$ decreases with the increase of $k$.
+This indicates that having a larger $k$ can potentially speed up the query processing of the partition index due to a smaller average partition border number.
 However, it also leads to a larger overlay graph, which can slow down the query processing and construction of the overlay index.
-Therefore, there should be a trade-off between them, and the best partition number typically varies for different PSP indexes.
-Furthermore, as per Figure 10 (g)-(i), only PUNCH and Bubble have a partition connectivity ratio $R_{C}$ of 1, which implies that at least one partition generated by other partition methods is unconnected (with multiple separate components).
-Since PTD requires the vertices within one partition to be connected, only PUNCH and Bubble suit it.
-Exp 2: Effect of Partition Method.
-Next, we further evaluate the effect of the partition method on the PSP indexes (PCH, PTD, and PPLL).
-In particular, we test PCH and PPLL with Bubble, PUNCH, KaHyPar, and METIS on road networks (NY and FL) while only testing PTD with PUNCH and Bubble by referring to Exp 1.
-Figure 11 shows that PUNCH and KaHyPar perform better than other partition methods in most cases due to smaller overall border number $|B|$ and average partition border number $\overline{B_{i}}$, which indicates that $|B|$ and $\overline{B_{i}}$ are important factors in the partition method selection.
-We select PUNCH as the default planar partition method due to its superior PSP index performance.
-It is worth noting that we also test PPLL with KaHyPar, METIS, and SCOTCH on complex networks.
-However, it fails to complete the index construction on the smallest graph GO within 48 hours due to the large and condensed overlay graph, which reveals that the PSP index with planar partition is generally unsuitable for complex networks.
-Exp 3: Effect of Partition Number.
-As shown in Figure 12 (a)-(b), the indexing time of PCH and PTD increased as $k$ increased from 4 to 256.
-This is because a larger $k$ could lead to a larger overlay graph that cannot leverage thread parallelization, thus resulting in higher indexing time.
-By contrast, PPLL’s indexing time first decreases and then increases in NY while tending to decrease in FL because a larger $k$ could also lead to a smaller $\overline{|B_{i}|}$, thus reducing the overlay graph construction time.
-In Figure 12 (c)-(d), we observe that a smaller $k$ generally improves the query time but leads to worse index update efficiency.
-Therefore, it is crucial to balance the query and update efficiency by selecting the appropriate partition number.
-We set the default value of $k$ as 32 since it provided satisfactory performance in most cases.
-Exp 4: Comparison with Non-partitioned SP Algorithms.
+Therefore, there is a trade-off between them.
+Furthermore, as per Figure 10 (q)-(x), among the partition methods used, only PUNCH and Bubble have a partition connectivity ratio $R_{C}$ of 1.
+This implies that the partitions generated by other partition methods generally contain multiple unconnected components.
+Since PTD requires the vertices within the same partition to be connected, only PUNCH and Bubble suit it.
+Overall, PUNCH and KaHyPar perform best in $|B|$ and $\overline{|B_{i}|}$ on road networks.
+When dealing with complex networks, it is advisable to use KaHyPar, METIS, or HEP.
+In the case of the largest graph FR, only HEP and CLUGP are capable of completing graph partitioning within 24 hours.
+Exp 2: Effect of Partition Method on PSP Index.
+Next, we choose Bubble, PUNCH, KaHyPar, and METIS as the partition methods according to Exp 1 and evaluate their effect on PSP indexes.
+We test PCH and PTD on road networks NY, FL, W, and US since they are infeasible on complex networks with large treewidth.
+We test PPLL on road networks and complex network GO and do not present the result on US since PPLL cannot finish index construction within 24 hours for any partition methods.
+Figure 11 shows that PUNCH outperforms other partition methods in most cases due to smaller border numbers $|B|$ and $\overline{B_{i}}$ as evidenced in Exp 1, which reveals that they are the most important factors in the partition method selection.
+For complex networks that PUNCH cannot be applied, KaHyPar has the best performance.
+We have tried to test PPLL on larger complex networks.
+However, it fails to complete the test within 24 hours due to the large and condensed overlay graph, which reveals that the planar PSP index is generally unsuitable for complex networks.
+Therefore, we select PUNCH and KaHyPar as the default planar partition methods for road networks and complex networks, respectively.
+Besides, we also evaluate the effect of partition number as presented in the appendix (Zhang et al., [n.
+d.]) due to limited space.
+In short, we set the default partition number as 32 for all datasets since it provided satisfactory performance in most cases.
+Table 4. Effectiveness of Overlay Optimization
+PSP Index PCH PTD PPLL
+Dataset Method $\tilde{D}$ $t_{c}$ $t_{q}$ $t_{u}$ $t_{c}$ $t_{q}$ $t_{u}$ $t_{c}$ $t_{q}$ $t_{u}$
+NY Bubble w/o opt 187.45 112.8 18.04 0.20 134.5 18.06 0.42 70.27 1743 11.77
+w/ opt 163.76 104.2 16.90 0.19 128.0 17.85 0.38 68.03 1561 10.00
+PUNCH w/o opt 53.53 4.67 3.29 0.02 5.33 0.32 0.016 25.94 9.70 0.19
+w/ opt 52.68 4.65 3.21 0.01 5.28 0.30 0.02 25.87 9.56 0.18
+FL Bubble w/o opt 235.80 143.6 44.56 0.06 195.9 33.75 0.12 748.7 3032 1.28
+w/ opt 211.87 138.0 38.82 0.05 187.8 31.37 0.09 742.4 2921 0.98
+PUNCH w/o opt 37.42 3.57 13.98 0.03 5.69 0.17 0.05 368.7 4.83 0.24
+w/ opt 37.26 3.43 13.81 0.03 5.61 0.16 0.04 365.7 4.34 0.24
+$\tilde{D}$: average vertex degree of $\tilde{G}$; $t_{c}$: indexing time (s); $t_{q}$: query time (ms); $t_{u}$: index update time (s).
+Figure 12. Effectiveness of PSP Strategy. Bar: Left, Ball: Right
+Exp 3: Effectiveness of Overlay Graph Optimization.
+We evaluate the effectiveness of overlay graph optimization by reporting $t_{c}$, $t_{q}$, $t_{u}$, and average degree $\tilde{D}$ of $\tilde{G}$ on NY and FL.
+We do not use large networks since they fail to report the results of PCH, PTD, and PPLL simultaneously.
+As shown in Table 4, it enhances the index performance in most cases with evident improvement in Bubble-based partitions (achieving $14.6\%$, $8.3\%$, $14.7\%$, and $17.8\%$ speed-up on $\tilde{D}$, $t_{c}$, $t_{q}$, and $t_{u}$).
+The improvement on PUNCH is marginal since it already has a great partition result.
+Nevertheless, this optimization improves the index performance by reducing the overlay graph density with a theoretical guarantee of the index correctness and we will use it by default in the following experiments.
+Figure 13. Performance of Different Index Update Types. Bar: Left, Ball: Right
+Figure 14. Edge Weight Batch Updates of PTD. Bar: Left, Polyline: Right
+Figure 15. Performance of Different Query Types
+Exp 4: Effectiveness of PSP Strategy.
+We evaluate the effectiveness of the proposed PSP strategies on PCH, PTD, PPLL, and CT in Figure 12.
+The no-boundary always the fastest in update, achieving up to 185.2×185.2$\times$185.2 ×, 91.3×91.3$\times$91.3 ×, 1759.4×1759.4$\times$1759.4 ×, and 5358.8×5358.8$\times$5358.8 × faster than the pre-boundary for PCH, PTD, PPLL, and CT.
+Moreover, it also outperforms the pre-boundary in index construction time in PCH and PTD with up to 2.4×2.4$\times$2.4 × and 1.7×1.7$\times$1.7 × faster.
+This is because no-boundary replaces the time-consuming Dijkstra’s searches among the boundary vertices with index query.
+Nevertheless, it is slow in query since the same-partition queries require concatenating the overlay and partition indexes.
+The post-boundary addresses this drawback by repairing the partition indexes, achieving equivalent query time as pre-boundary.
+Moreover, since it needs to repair all-boundary-pair distances, it is time-consuming when the query efficiency of $\tilde{L}$ is low.
+For example, because the CH query processing is relatively slow, the index construction and maintenance of PCH with post-boundary is even slower than the pre-boundary as it involves twice all-pair searches among the boundaries.
+Nevertheless, the post-boundary could achieve a significant update efficiency improvement (14.659.3×14.6-59.3$\times$14.6 - 59.3 × for PTD) for PTD, PPLL, and CT that use 2-hop labeling (TD and PLL) as the underlying SP index.
+Exp 5: Evaluation of Update Types.
+As shown in Figure 13, no-boundary is much faster (up to 185.2×185.2$\times$185.2 ×, 91.3×91.3$\times$91.3 ×, 1759.4×1759.4$\times$1759.4 × for PCH, PTD, and PPLL) than pre-boundary in all update types.
+The post-boundary works better than the pre-boundary on PTD and PPLL since it replaces all-pair searches with index query.
+Note that the performance of edge/vertex insertion (deletion) is almost equivalent to edge decrease (increase) since this structural update essentially shares a similar update mechanism with the edge weight update.
+Furthermore, the batch update for PTD is tested by varying batch size from 100 to 500.
+As shown in Figure 14, PTD-No and PTD-Post always outperform PTD-Pre.
+As the batch size increases, the improvement decreases since the all-pair boundary searches in the pre-boundary could be shared by different edge updates.
+Nonetheless, PTD-No and PTD-Post also share the computation for different edge updates during the all-pair boundary distance check and outperform PTD-Pre.
+Exp 6: Performance of Different Query Types.
+We classify the queries into four types based on $s$ and $t$ locations: Q1 (same partition), Q2 (both boundary), Q3 (one boundary one non-boundary), Q4 (different partitions).
+As shown in Figure 15, for PTD and PPLL, Q2 is the fastest since it only uses $\tilde{L}$.
+By contrast, Q3 and Q4 are much slower due to distance concatenation.
+For Q1, the post-boundary is the same with pre-boundary since they both only use the partition index, while the no-boundary has to concatenate $\tilde{L}$ and $\{L_{i}\}$.
+Lastly, PCH has similar query efficiency on different query types and strategies due to its search manner.
+Figure 16. Comparison with Non-partitioned SP Indexes. (Bar: Left axis, Ball: Right axis)
+Figure 17. PSP Indexes Comparison. Bar: Left, Ball: Right
+Exp 7: Comparison with Non-partitioned SP Algorithms.
 We conduct a comparison between PSP indexes (including PCH, PTD, PPLL, and CT) and their non-partitioned counterparts (namely CH (Geisberger et al., 2008), TD (Ouyang et al., 2018), PLL (Akiba et al., 2013)).
@@
 Moreover, CT demonstrates 72.3×72.3$\times$72.3 × and 3.58×3.58$\times$3.58 × improvement over PLL in terms of indexing time and index size, respectively.
-However, PSP indexes may compromise query processing efficiency and index update time.
-It is worth noting that it also shows that planar structure is more suited for road networks, while core-periphery structure is better suited for complex networks, as PPLL does not outperform PLL in terms of indexing time and scalability.
-Exp 5: Effectiveness of PSP Strategy.
-As shown in Figure 14 (a)-(d), the no-boundary strategy has the minimum index construction (about $14-35\%$ speed-up) and index update time (up to two orders of magnitude speed-up compared with the pre-boundary strategy for PTD and PCH), which indicates the superiority of our no-boundary strategy.
-Suffering from the highest index construction time, the post-boundary strategy has the same query time as the pre-boundary strategy but a significant update efficiency improvement (14.636.7×14.6-36.7$\times$14.6 - 36.7 × speed up for PTD).
-To further demonstrate the effectiveness of our post-boundary strategy, we test the same-partition query efficiency as presented in Figure 14 (e)-(f).
-The post-boundary archives about 3.76.3×3.7-6.3$\times$3.7 - 6.3 × and 4.830.5×4.8-30.5$\times$4.8 - 30.5 × speed-up compared with the no-boundary strategy for PTD and PPLL, respectively.
-Such a significant performance gain comes from the post-boundary’s correct local partition index.
-Finally, we evaluate the effectiveness of pruning-based overlay graph optimization by measuring the average vertex degree of the overlay graph, as shown in Table 3.
-We test PUNCH and HEP on road networks while the KaHyPar and HEP on complex networks.
-The degree decreases with the pruning-based optimization in most cases.
-It cannot improve PUNCH since PUNCH already has a great partition result.
-Figure 15. Comparison of All Shortest Path Indexes. Bar: Left, Ball: Right
-Exp 6: Comparison of All PSP Algorithms.
-We compare our proposed PSP indexes with the existing PSP indexes in Figure 15.
-We first analyze the performance of our proposed indexes against their counterparts: 1) P-TD-P has the same query performance and index size as FHL, but it is much faster to update with $1$ 202.7 × speed-up.
-2) N-CH-P has the fastest index construction and maintenance efficiency and the smallest index size among all planar PSP indexes, achieving up to 3.4×3.4$\times$3.4 ×, 1004×1004$\times$1004 ×, and 7.6×7.6$\times$7.6 × speed-up over FHL.
-3) N-PC-CP has faster index construction and update than CT since it utilizes faster CH for tree index.
-P-PT-CP outperforms CT in terms of query processing as it is faster in same-partition queries.
-4) N-TS-HP has faster index construction, smaller index size, faster query and update than G-Tree by leveraging TD for overlay index construction.
-5) Sketch can scale up to very large graphs (the only algorithm that can be applied on graph FR), but its query efficiency is much slower than N-PC-CP and P-PT-CP due to the light index.
+However, PSP indexes generally compromise query processing efficiency and even index update time (except for PTD).
+Besides, it can also be seen that planar structure is more suited for road networks, while core-periphery structure is better suited for complex networks, as only CT can finish tests on complex networks while PPLL cannot.
+Exp 8: Comparison of All PSP Algorithms.
+We compare our proposed PSP indexes with the existing ones in Figure 17.
+We first analyze the performance of our proposed indexes: 1) P-TD-P has the same query efficiency and index size as FHL but is much faster to update with $1$ 202.7 × speed-up; 2) N-CH-P has the fastest index construction and maintenance efficiency and the smallest index size among all planar PSP indexes, achieving up to 3.4×3.4$\times$3.4 ×, 1004×1004$\times$1004 ×, and 7.6×7.6$\times$7.6 × speed-up over FHL; 3) N-PC-CP has faster index construction and update than CT since it utilizes faster CH for tree index.
+P-PT-CP outperforms CT in terms of query processing as it is faster in same-partition queries; 4) N-TS-HP has faster index construction, smaller index size, faster query and update than G-Tree by leveraging TD for overlay index construction; 5) Sketch can scale up to very large graphs (the only algorithm that can be applied on graph FR), but its query efficiency is much slower than N-PC-CP and P-PT-CP due to the light index.
 We next discuss the advantages of different partition structures.
@@
 7.3. PSP Application Guidance
-Based on the experimental analysis in Section 6.2, we clarify the misconceptions mentioned in Section 1 and conclude the following observations: 1) all PSP indexes are essentially interrelated because they all consist of three dimensions in our PSP index scheme: graph partition method, PSP strategy, and shortest path algorithm.
+Based on the experimental analysis above, we clarify the misconceptions mentioned in Section 1 and conclude the following observations: 1) All PSP indexes are essentially interrelated because they all consist of three dimensions in our PSP index scheme: graph partition method, PSP strategy, and shortest path algorithm.
 Moreover, we could combine the elements of these three dimensions to create a new PSP index; 2) It is important to choose the right partition structure for different types of graphs.
@@
 On the other hand, core-periphery PSP indexes are more suitable for complex networks due to their excellent scalability; 3) Although PSP indexes typically have faster indexing efficiency, smaller index size, and better scalability than their non-partitioned counterparts, they sacrifice query and index update efficiency.
+Figure 18. Our Decision Tree for PSP Index Design
 According to our theoretical analysis and experimental results, we further provide our decision tree for PSP index design as illustrated in Figure 18.
@@
 If tree decomposition is infeasible, we recommend selecting PLL or even the direct search for SP algorithm selection.
-Figure 16. Our Decision Tree of PSP Index Design
-7. Related Work
-Recently, multiple surveys about the SP computation have emerged, focusing on either specific graph types (like transportation networks (Bast et al., 2016) or complex networks (Zhang et al., 2022)) or static graphs (Madkour et al., 2017; Li et al., 2017; Wu et al., 2012).
-(Zhang et al., 2021a) investigates the computation in dynamic graphs, but it does not consider the partitioned counterpart.
-The partitioned pathfinding framework is proposed in (Chondrogiannis and Gamper, 2014, 2016), however, the dynamic version is not discussed and they can be covered by the traditional PSP strategy.
-In this paper, we study the dynamic PSP problem systematically and explore its performance in both road networks and complex networks by involving the state-of-the-art discovery of SP algorithms and partition methods.
-Next, we discuss and categorize the existing PSP methods based on our scheme:
-1) Pre-Boundary + Search/All-Pair + Hierarchy/Planar: These methods (HiTi (Jung and Pramanik, 2002), Graph Separators, Customizable Route Planning (Delling et al., 2017) (Holzer et al., 2009), ParDiSP (Chondrogiannis and Gamper, 2014, 2016)) pre-compute the shortest distance between boundaries (some hierarchically) to guide the search.
-In a broad sense, Arc-flag (Möhring et al., 2007) and SHARC (Bauer and Delling, 2010) also belong to this category.
-G-tree (Zhong et al., 2013; Li et al., 2019a) uses dynamic programming to compute the distance between layer’s all-pair information to replace searching, and it is widely used in kNN (Zhong et al., 2015), ride-sharing (Ta et al., 2017), time-dependent routing (Wang et al., 2019), and machine learning-based path finding (Huang et al., 2021).
-They are slow to construct due to Pre-B and slower to query to direct search;
-2) Pre-Boundary+Search/PLL+Planar: COLA (Wang et al., 2016) builds the labels for the skyline shortest path on the overlay graph to answer the constrained shortest path query.
-This structure’s construction suffers from Pre-B and query suffers from searching;
-3) Pre-Boundary+TD/TD+Planar: FHL (Liu et al., 2021, 2022b, 2022a) partition builds the TD both within and between partitions for multi-dimensional skyline paths.
-As validated in Exp 4, our Post-B can speedup Pre-B’s construction dramatically;
-4) Pre-Boundary+PLL/ PLL+Planar: T2Hop (Li et al., 2019c, 2020b) utilizes two layers of PLL to reduce the complexity of long range time-dependent paths, and this structure’s performance is limited by PLL;
-5) No-Boundary+ PLL/TD+Core: Core-Tree (Li et al., 2020a; Zheng et al., 2022) is the baseline that we already discussed;
-6) Pre-Boundary+Search/All-Pair/PLL+Sketch: This category works on huge graphs where index is nearly impossible so only a small number of landmarks are selected to either help prune the search (Das Sarma et al., 2010; Wang et al., 2021; Farhan et al., 2022) or approximate result (Gubichev et al., 2010; Qiao et al., 2012).
-In sum, all the existing PSP methods have their limitation, warranting the comprehensiveinvestigation embarked on in this paper.
+8. Future Directions
+The insightful structure-based partition classification and universal PSP scheme proposed in this paper provide a systematical way to analyze and design the PSP index.
+Besides, the novel no-boundary and post-boundary strategies significantly enrich the design space for the PSP index.
+However, there are still some unresolved problems with the PSP index.
+Next, we will discuss some future directions and open challenges to inspire researchers in this field.
+1) I/O-efficient and Distributed PSP Algorithms.
+There are two general ways to scale up graph algorithms: scale up to external memory on a single machine (out-of-core PSP index), or scale out to a distributed cluster (distributed PSP index).
+Since the PSP indexes naturally decompose a graph into several subgraphs, we could regard each subgraph (or its partition index) as a whole for disk accessory to reduce the I/O numbers or distribute each subgraph to a server in the cluster to reduce communication.
+Nevertheless, how to partition the graph for fewer I/O numbers or communication workload according to the network structure and algorithm procedure is still an open challenge.
+Moreover, how to optimize the PSP strategy and underlying SP algorithm simultaneously according to the I/O accessory or communication pattern is also an open challenge.
+2) Automatic PSP Index Design System.
+Although we have propose a universal PSP index scheme and insightful guideline for the PSP index design, they are essentially experience-based.
+In the era of Artificial Intelligence (AI), how to leverage AI to automatically design the PSP index according to the system requirements is an open challenge.
+Moreover, how to make the elements of different dimensions in PSP index scheme portable such that the system could react promptly with the evolving system requirements is also an open challenge.
+3) High Throughput System Based on PSP Index.
+Another future direction is to leverage PSP index for high throughput shortest path query processing.
+For example, road networks generally incur high throughput queries while the edge weight of road networks (travel time) could dynamically changes due to evolving traffic conditions.
+Therefore, how to achieve fast index maintenance and query processing simultaneously is important to enhance the system throughput.
+The PSP index could leverage parallelization among partition indexes to speed up the index update.
+Besides, adopting 2-hop labeling as the underlying SP algorithm also contributes to a fast query processing, making the PSP index a promising option for building path-finding system with high throughput.
+Nevertheless, how to design such a PSP index for high throughput query processing on dynamic road networks is an open challenge.
+Besides, partial edge updates in a partition may not influence all queries on other partitions.
+Thus how to identify the affected queries such that the system could still leverage the unaffected index for fast query processing is another open challenge.
 9. Conclusions
-In this work, we systematically study the partitioned shortest path index and propose a universal PSP scheme by concomitantly considering three dimensions: PSP strategy, SP algorithm, and partition structure.
-For the PSP strategy, we propose two new strategies and pruned-boundary optimization for better index performance.
-We also provide index maintenance solutions for both traditional and novel PSP strategies.
-For partition structure, we propose a new path-oriented partition method classification.
+In this work, we systematically study the partitioned shortest path index and propose a universal PSP scheme by considering three dimensions: SP algorithm, partition method, and PSP strategy.
+We first briefly review the SP algorithms used in the PSP index and then propose a new structure-based classification to facilitate the selection of the partition method and conduct a comprehensive review of corresponding partition methods and PSP indexes.
+For the PSP strategy, we summarize the classic pre-boundary strategy and then propose two new strategies and pruning-based overlay graph optimization to improve the index performance.
+We also provide index maintenance methods for all PSP strategies.
 Furthermore, we construct our PSP index scheme by seamlessly coupling those three dimensions.
-We also propose a PSP index generator and five new PSP indexes that are either more efficient in query or update than the current state-of-the-art.
-Systematic and extensive experimental evaluations fully validate our findings and prove the effectiveness of our new indexes guided by the insightful PSP index scheme.
-Finally, PSP index design guidelines are provided to help future researchers and practitioners design their indexes.
-Our future work will address recommendation aspects for such designs.
+We also propose five new PSP indexes that are either more efficient in query or update than the current state-of-the-art.
+Systematic and extensive experimental evaluations demonstrate the effectiveness of our proposed techniques and new indexes guided by the insightful PSP index scheme.
+An insightful PSP index design guideline is provided to help future researchers and practitioners design their indexes.
+Lastly, the future directions and open challenges for the PSP index are also discussed.